### v1

# CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes

Abstract
We present CD-NGP, which is a fast and scalable representation for 3D reconstruction and novel view synthesis in dynamic scenes. Inspired by continual learning, our method first segments input videos into multiple chunks, followed by training the model chunk by chunk, and finally, fuses features of the first branch and subsequent branches. Experiments on the prevailing DyNeRF dataset demonstrate that our proposed novel representation reaches a great balance between memory consumption, model size, training speed, and rendering quality. Specifically, our method consumes $85\%$ less training memory ($<14$GB) than offline methods and requires significantly lower streaming bandwidth ($<0.4$MB/frame) than other online alternatives. Furthermore, we conduct experiments on a self-collected extra-long video sequences dataset, which demonstrates the superior scalability of our method over previous state-of-the-art approaches.
Keywords:
Continual learning Dynamic novel view synthesis Neural Radiance Field.
1 Introduction
Neural radiance Field (NeRF) [30] has made significant progress in 3D scene reconstruction and novel view synthesis. It takes multi-view RGB images with their camera poses as input and leverages multi-layer perceptrons (MLPs) to reconstruct both the geometry and color of the scene. Many variants have been proposed to improve NeRF’s efficiency [9, 38, 31, 56, 53], reconstruction quality [2, 3, 4], scalability [47, 50], and pose robustness [5, 26, 13, 52, 32, 51]. Recently, 3D Gaussian Splatting (3DGS) [22] proposes an explicit representation with differentiable rendering, which significantly speeds up the reconstruction and enables real-time novel view synthesis. Despite significant progress having been made in this field, the above-mentioned methods assume that the scene is static and may not work properly in dynamic scenes.
To address the dynamic reconstruction problem, amount of efforts have been made under the offline setting, requiring access to the entire dataset in the whole training stage. For example, [36, 28, 34, 33, 15, 29, 44] propose to use MLP-based representation and achieve high-quality results. [40, 17, 7, 16, 45] speed up the optimization by using tensor factorization and voxel representation, respectively. The more recent continual-learning-based methods [27, 48] can train models frame by frame, and streamable representations [43] enable the model to be loaded on the fly at inference. However, the offline methods require huge host memory e.g., taking more than $100GB$ when reconstructing multi-view dynamic videos of 300 frames, and the continual learning-based methods are weak in balancing the convergence speed and model compactness. As a result, these methods are limited in scalability, making them unsuitable for long videos.
In this paper, we propose Continual Dynamic neural graphics primitives (CD-NGP), a novel, fast, and scalable representation for reconstructing dynamic scenes, which enables dynamic reconstruction in long videos. As illustrated in Fig. 1, we evenly split the dynamic videos into multiple chunks based on the timestamp and conduct reconstruction with different model branches, where each branch includes MLP regressors, spatial encoding for $(x,y,z)$, and temporal encoding for $t$. We encode the different chunks by hash tables with different complexities to ensure model scalability and then continually train the model chunk by chunk to alleviate memory occupation. For each individual chunk, we locate the corresponding model branch, sample 3D points $(x,y,z)$ according to frame pixels and cached occupancy grids [31], and represent spatial features by spatial hash grids. The spatial features of the first branch and the current branch are fused by element-wise sum as default or concatenation if applicable. The temporal features are encoded by the hash table for time stamp $t$ in the current branch only. Spatial and temporal features are concatenated along with viewing direction $d$ and input to MLP regressors for rendering and optimization.
Extensive experiments demonstrate the great scalability of our method: compared with prevailing offline baselines, our method consumes $85\%$ less host memory at the training phase, $40\%$ less model size at the inference phase, and delivers comparable results. Compared with online baselines, our method achieves compact model size, high quality, and fast convergence simultaneously. The main contribution of this work is summarized as 3 folds:
Figure 1: The demonstration of our continual learning pipeline for dynamic scenes. Please note the dimensions of representations for different branches remain the same while the complexities of hash tables are flexible and could be configured to ensure scalability. The spatial hash features could be encoded along voxels [31], planes [17] or both [38].
  • We propose a novel method combining spatial and temporal hash encodings for dynamic scenes in short intervals with high scalability.
  • We propose a fast continual learning framework for dynamic scenes, producing comparable results to offline state-of-the-art methods while reducing memory and storage occupation significantly.
  • We provide a dataset of extra-long video sequences ($>3$ minutes) and further demonstrate the scalability of our method on long multi-view videos.
2 Related Work
2.1 Fast novel view synthesis in static scenes
The acceleration of NeRFs falls into two categories: training acceleration and rendering acceleration. Plenoxels [18] uses voxel grids to represent the scene and achieves magnitudes of acceleration over MLP backbones. TensoRF [9] decomposes the 3D scene as the product of tensors and achieves fast novel view synthesis. Plane-based methods [17, 21] project 3D points onto 3 orthogonal 2D planes to ensure sparsity and achieve fast convergence. Instant-NGP [31] leverages an occupancy grid and a multi-resolution hash encoding along with optimized tiny MLPs to enable fast novel view synthesis within a minute. The recent 3D Gaussian method [22] represents the scenes as a union of explicit Gaussian point clouds and achieves high fidelity, fast convergence, and real-time rendering on high-end hardware by eliminating the querying burden of spatial representations. KiloNeRF [37] distills the large MLP in NeRF [30] into tiny MLPs to accelerate rendering by thousands of times. Hedman et al. [20] propose the baking method to accelerate the rendering process. They alpha-compose the cached features along the ray and use MLPs only once per ray to overcome the computation load of the per-point MLP querying in NeRF. BakedSDF [56] stores the view-dependent features into Gaussian spheres along the reconstructed surfaces, and achieves real-time rendering on ordinary laptops. MobileNeRF[10] represents scenes into meshes and enables real-time rendering on mobile devices. Light field methods [42, 8, 19] directly represent the scene as a set of rays to eliminate volume rendering and enable real-time applications on mobile devices.
2.2 Representations of dynamic scenes
There are mainly 3 variants of NeRF for dynamic scenes: MLP, tensor factorization, and voxel representations. D-NeRF [36] leverages a large deformation MLP to model the transition and the trajectories of the objects and extends NeRF to simulated dynamic scenes. DyNeRF [28] uses long latent vectors to encode temporal information along with large MLPs to achieve high-fidelity novel view synthesis in real dynamic scenes. Park et al. [34] propose an interpolation-based MLP representation for dynamic scenes and achieve significant acceleration over DyNeRF. HyperReel [1] proposes a network to predict the sample locations and represent dynamic scenes by key-frame-based TensoRF [9] representation to ensure high fidelity and real-time rendering without any custom CUDA kernels. K-Planes [17] and HexPlane [7] factorize the whole 4D field to the product of 2D planes. Though differently implemented, they both achieve high-fidelity results with intermediate model size (200MB) and training time (100 minutes) on DyNeRF dataset. TineuVox [16] leverages the deformation MLP in D-NeRF and implements a 3D voxel grid with multi-scale interpolation. It achieves high fidelity and fast convergence on D-NeRF dataset. MixVoxels [45] leverages the standard deviation across the time of all pixels to estimate a variation field to decompose the dynamic scene into static and dynamic voxels. Equipped with different representations for static and dynamic objects, MixVoxels achieves high-fidelity results and super-fast convergence (15 minutes) on DyNeRF dataset. Song et al. [43] train Instant-NGP [31] and TensoRF [9] offline and stream the models along the feature channels for dynamic scenes. Concurrent works are leveraging 4D Gaussian point clouds [54, 49] or 3D Gaussians with deformation networks [55] achieving real-time rendering and high fidelity in dynamic scenes.
2.3 Continual learning of NeRF
Continual learning (online learning) aims to adapt to new tasks while defying catastrophic forgetting (resisting performance downgrade on the previous tasks) [12]. There are mainly 3 approaches to achieving continual learning: replay, regularization, and parameter isolation. The replay-based methods either store the previous data in a buffer and replay them [39], or use the model to generate the pseudo-ground truth of previous tasks while training on new tasks [41]. The regularization-based methods hold the parameters close to the solution of previous tasks [14]. The parameter-isolation methods use different parameters to cope with different tasks. Recent efforts based on generative replay [6, 35, 11] enable continual learning of NeRF in static scenes and achieve high fidelity. StreamRF [27] applies continual NeRF to dynamic scenes by parameter isolation. Their explicit voxel representation enables fast reconstruction (75 minutes for 300 frames) and could be streamed frame by frame. INV [48] uses a shared color MLP and trains different structure MLPs frame by frame to form a continual NeRF representation and achieves high-fidelity results in dynamic scenes.
3 Methodlogy
3.1 Preliminaries
Inspired by the previous NeRF variants [36, 28, 17, 7, 45, 31, 38] and continual learning methods [6, 35, 11, 27, 43, 48], we focus on the dynamic novel view synthesis problem in the multi-view setting and solve the problem in a continual learning manner to alleviate the huge memory consumption and bandwidth costs.
Following previous methods [36, 28, 17, 7], we consider the dynamic scene as a set of 3D points $\mathbf{p}=(x,y,z)$ with optical attributes $(\sigma,\mathbf{c})$, where $\sigma$ is the volume density of the point and $\mathbf{c}$ is the color of the point. We train our model along camera rays to reconstruct the scene. A camera ray is defined as:
$\mathbf{r}(u)=\mathbf{o}+u\mathbf{d},$ (1)
where $\mathbf{o}$ is the optical center of the camera, $\mathbf{d}$ is the direction off the camera. For each point along the camera ray with coordinates $(x,y,z)$ observed at direction $\mathbf{d}=(\phi,\psi)$, a field function $F$ is queried to obtain the optical attributes $(\sigma,\mathbf{c})$. Conditioned on time $t$, the entire model is defined as:
$(\sigma,\mathbf{c})=F_{\Theta}(x,y,z,t,\phi,\psi),$ (2)
where the function $F$ denotes the dynamic NeRF model and $\Theta$ denotes trainable model parameters.
Let $I_{i}$ denote the $i^{th}$ input image corresponding to observation direction $d_{i}$, $\mathbf{V}=\{\nu_{i}~{}|~{}\nu_{i}=(I_{i},d_{i}),1\leq i\leq N_{view}\}$ is the set of the training views, where $i,N_{view}\in\mathbb{Z^{+}}$ and $\mathbf{D_{a}}$ is the readily-accessible data at one step. Previous methods [6, 35] for static scenes define the continual learning problem as learning a unified representation with a limited subset of views $V_{a}$ available at every optimization step, i.e. $\mathbf{D_{a}}=V_{a}\subsetneq V$. However, a more natural definition of continual NeRF is in the dynamic scenes, where the observations at different time stamps arrive continually. Denoting all the frame times as $\mathbf{T}=\{\tau_{i}~{}|~{}1\leq i\leq N_{time}\},i,N_{time}\in\mathbb{Z^{+}}$, we define the continual learning problem as synthesizing a video sequence when the model could only have access to a limited number of frames $\mathbf{T_{a}}\subsetneq\mathbf{T}$ from all views $\mathbf{V}$ at each step, i.e. $\mathbf{D_{a}}=\{\mathbf{V}\times\mathbf{T_{a}}\}$ where $\times$ is the Cartesian product of sets.
To simplify the notations, let the whole sequence of frames be evenly divided into $k$ chunks where a chunk $\mathbf{T_{k}}=\{\tau_{i}~{}|~{}\lfloor\frac{N_{time}}{T_{chunk}}\rfloor=k\}$ is a sequential subset of length $T_{chunk}$. We define the continual learning problem as learning a representation where there are only the frames in the current chunk available at each optimization step, i.e. $\mathbf{T_{a}}=\mathbf{T_{k}}$. Accordingly, $T_{chunk}=1$ denotes training the model frame by frame, $T_{chunk}=10$ denotes training the model with a buffer of 10 frames, $T_{chunk}=300$ denotes training the entire 300-frame sequence offline. The definition is close to real-world applications since caching all training frames in memory or storing them on the disk is limited in scalability for long videos.
3.2 Representation of Short Dynamic Intervals
As pointed out by [46], naive 4D hash grids are not suitable for high-fidelity dynamic view synthesis. To achieve fast reconstruction and slim model size simultaneously, we represent time and spatial features separately and concatenate them to alleviate feature confusion. The spatial hash features may be encoded along 3D voxels [31], 2D planes [17, 21], or both [38]. The time feature is encoded per frame following [28]. Following previous state-of-the-art methods [30, 31, 16], we apply a two-fold field function for geometry and color individually. The architecture of the field function is thus defined as:
$(\sigma,\mathbf{H})=f_{\theta_{1}}(\Psi(\mathbf{x})\oplus\Gamma(t)),$ (3)
$\mathbf{c}=f_{\theta_{2}}(\mathbf{H},\phi,\psi).$ (4)
where $\Psi(\mathbf{x})$ denotes configurable spatial hash encodings, $f_{{\theta_{1}}}$ is a tiny MLP with trainable parameter ${\theta_{1}}$, $\Gamma(t)$ is the hash encoding for time axis alone, $\oplus$ is the concatenation of vectors. The MLP $f_{{\theta_{1}}}$ in Eq. 3 fuses features across the spatial domain and time domain and obtains the volume density $\sigma$. View-dependent colors $\mathbf{c}$ are regressed by another tiny MLP with parameter $\theta_{2}$ in Eq. 4. The spatial encoders $\Psi(x)$ follow a multi-resolution concatenated format, defined as:
$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}~{}P_{m}(x,y,z)$ (5)
where $P_{m}(x,y,z)\in\mathbf{R^{F}}$ denotes the tri-linearly interpolated value from the nearby 8 grid vertices w.r.t. 3D coordinates $(x,y,z)$ at the $l^{th}$ resolution in Instant-NGP [31], $L$ is the number of the resolutions. $\Psi(x)$ may also be configured as the fusion of 2D and 3D features in MERF [38], defined as:
$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}~{}P_{m}(x,y,z)+Q_{mxy}(x,y)+Q_{myz}(y,z)+Q_{mzx}(z,x)$ (6)
where $Q_{mij}(i,j)$ denotes the bi-linearly interpolated value from the 4 nearby vertices in the orthogonally projected planes w.r.t. axes $(i,j)$ at the $m^{th}$ resolution. Removing the $P_{m}(x,y,z)$ term leads to a pure plane representation like [17, 21].
3.3 CD-NGP
Unlike the existing replay-based methods for static scenarios [6, 35], we solve the continual learning problem of dynamic novel view synthesis via parameter isolation. The key insight is that the current state-of-the-art methods by tensor factorization [7, 17] or explicit voxels [45] are not accurate estimations of the object trajectory and inherent dynamics, but only simple regressions of the optical attributes in 3D space within a limited time. Therefore, the replay method is of limited use because the new transient processes are constantly arriving and the model of a limited size is not capable of reconstructing unlimited transient effects and thus suffers from catastrophic forgetting.
To address this issue, we propose our Continual Dynamic neural graphics primitives (CD-NGP) based on parameter isolation. We split the model into $N_{chunk}$ branches to represent different chunks of the dynamic scene accordingly. The $k^{th}$ branch contains spatial hash encoder $\Psi_{k}(\mathbf{x})$, temporal hash encoder $\Gamma_{k}(t)$, and MLPs $f_{\theta_{1k}},f_{\theta_{2k}}$. We train the first branch (base branch) with a large spatial hash table to better reconstruct the scenes with initial movements and static objects. Then we reuse the spatial features from the base branch and train the subsequent branches (auxiliary branches) to compensate for residuals. We thus substitute the $\Psi(\mathbf{x})$ in Eq. 3 with the fused feature to obtain the volume density $\sigma$ at the $k^{th}$ model branch under the continual learning setting as:
$(\sigma,\mathbf{H})=f_{\theta_{1k}}((\Psi_{0}(\mathbf{x})+\Psi_{k}(\mathbf{x}))\oplus\Gamma_{k}(t)),$ (7)
where $\Psi_{0}(\mathbf{x})$ denotes the spatial encoder of the base branch, $\Psi_{k}(\mathbf{x})$ denotes the spatial encoder of the current branch. The view-dependent colors are regressed as Eq. 4 with the color MLP $f_{\theta_{2k}}$.
The scalability of our method derives from the redundancy of dynamic scenes and the flexibility of hash table sizes. As pointed out by [45], most objects ($>90\%$) in the dynamic scenes are static. Therefore, the first chunk could be represented by a large hash table to reconstruct the scenes with initial movements and static objects, the subsequent chunks could be represented by small hash tables to compensate for errors, and the whole representation could be designed to leverage the fusion of features from the corresponding two branches to ensure scalability. To be specific, if we follow the default setting of Instant-NGP [31] and set the hash table in the first branch as $2^{P_{1}}=2^{19}$, the sizes of hash tables in subsequent branches could be configured as $2^{P_{2}}=2^{14}$, which consumes only $1/32$ of the parameters in the first branch. Since the spatial hash tables consume the most parameters in the model, and most chunks are encoded by the subsequent branches, the advantage of our asymmetric hash table structure is obvious: compared with the previous linear-streaming method [43], the total bandwidth consumption is reduced from $\mathcal{O}(\mathbf{K}\cdot 2^{19}LF)$ to $\mathcal{O}(N_{chunk}\cdot 2^{14}LF+2^{19}LF)$, where $\mathbf{K}$ is their bandwidth budget controlling the bitrates in streaming scenarios. If the bandwidth is extremely limited, the hash tables in our model could be arranged as Eq. 6 to reduce the bandwidth cost additionally.
3.4 Rendering and Optimization
Given the sampled optical attributes on the rays in Eq. 1, we adopt volume rendering to get the 2D pixels, defined as:
$\hat{C}(r)$ $=\sum_{i=1}^{N}T_{i}\alpha_{i}\mathbf{c_{i}}=\sum_{i=1}^{N}T_{i}(1-e^{\sigma_{i}\delta_{i}})\mathbf{c_{i}},$ (8)
where $T_{i}=\exp(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j})$ denotes the transmittance, $\delta_{i}$ is the interval between neighboring samples, $N$ is the number of samples along the ray. The model is optimized by L2 photometric loss with three regularization terms, defined as:
$L=\frac{1}{|\mathbf{R}|}\sum_{r\in\mathbf{R}}\|C(r)-\hat{C}(r)\|_{2}^{2}+\lambda_{d}L_{d}+\lambda_{o}L_{o}+\lambda_{r}L_{r}$ (9)
where $\mathbf{R}$ denotes the set of the camera rays. $L_{d}$ is the distortion loss in Mip-NeRF 360 [3], defined as:
$L_{d}$ $=\sum_{i,j}w_{i}w_{j}\left\lvert\frac{s_{i}+s_{i+1}}{2}-\frac{s_{j}+s_{j+1}}{2}\right\rvert$
$+\frac{1}{3}\sum_{i}w_{i}^{2}(s_{i+1}-s_{i})$ (10)
where $s_{i}$ denotes the normalized ray distance along the ray, $w_{i}=T_{i}\alpha_{i}$ is the weight of the $i^{th}$ sample. $L_{o}$ is the entropy on the opacity by [25], defined as:
$L_{o}=-o\log(o),\quad o=\sum_{j=1}^{N}T_{j}\alpha_{j}.$ (11)
$L_{r}$ is L1 regularization on the spatial features in the current branch as:
$L_{r}=||\Psi_{k}(\mathbf{x})||_{1}.$ (12)
We set $\lambda_{r}$ as 0.001 to encourage leverage of the base hash encoder. Based on [25], we set a larger $\lambda_{d},\lambda_{o}$ as 0.005 for alleviating foggy effects.
4 Experiments
4.1 Datasets
DyNeRF dataset [28] is the most prevailing dataset for dynamic view synthesis that provides high-quality multi-camera videos, on which most of the current state-of-the-art methods are evaluated. DyNeRF dataset contains 6 individual cooking scenes with varied illumination, food topological changes, and transient effects like flames. It contains five 300-frame videos and a 1200-frame video, with up to 21 synchronized camera views at the resolution of $2704\times 2028$ at 30 frames per second (FPS). To further prove the viability of our method on long video sequences, we provide a new dataset containing 5 dynamic scenes at $3840\times 2160$ resolution, the lengths of which are at least 200 seconds. Similar to the DyNeRF dataset [28], scenes in our dataset are recorded by 20 GoPro cameras at 30 FPS. Compared with those in the DyNeRF dataset, the dynamics are much harder, including large motions of human bodies and transient effects in the background. Please refer to the supplementary material for more details.
Figure 2: Example frames sampled from the proposed long multi-view video dataset.
4.2 Implementation details
Our implementation is based on HexPlane [7], Instant-NGP [31], and NGP-pl [25]. All the metrics are measured on a single RTX 3090 GPU if not otherwise specified. Similar to the previous work [7, 28], we select the view in the center with index 00 for evaluation on both datasets. Following [7], we exclude the unsynchronized coffee martini scene and conduct experiments on the others. Following most of the state-of-the-art methods [28, 27, 17, 45, 43] we train and evaluate the model at resolution $1352\times 1014$ on the DyNeRF dataset. We use $1280\times 720$ resolution for our dataset to fit into memory. We use the importance sampling scheme in [28] at a batch size of 1024 rays. The parameters are optimized by Adam [23], with $\epsilon=10^{-15},\beta=(0.9,0.96)$. The learning rate is set as 0.001 and scheduled by cosine function following [25]. We set the size of the hash table in the first branch as $2^{19}$ and the size of the hash table in the subsequent branches as $2^{14}$, and we set the structural parameters $(L,F)$ in the spatial hash encoders as $(12,2)$ if not otherwise specified. At the base branch, we initialize the models and train for 18k iterations, and for each of the subsequent branches we train for 3k iterations using the previous branch for initialization. We use LeakyReLU activation in the MLP regressors. For long video sequences, we assign 30 chunks as an episiode. The MLP regressors are re-trained from initialization in the first branch of each episode to adapt to recent changes. Strategies for updating the occupancy grids are identical with [31]. To save storage, we use only one shared occupancy grid for all the branches. We report the model sizes with the occupancy grid under bfloat16 precision and parameters in float32 precision. Please see the supplementary material for more details.
4.3 Results and Comparisons
Table 1: Comparisons between our CD-NGP and SOTA methods on the DyNeRF dataset. Memory denotes the maximum host memory required for training. - denotes the metrics unreported in the original publications. denotes the methods measured only on the most challenging flame salmon scene. Best, Second, and Third metrics are in highlight.
[…]
offline DyNeRF [28] 300 29.58 0.020 0.083 7 days 28MB >100GB
NeRF-T [28] 300 28.45 0.023 0.10 - - >100GB
Interp-NN [34] 300 29.88 - 0.096 2 days 20MB >100GB
HexPlane [7] 300 31.57 0.016 0.089 96 min 200MB 62GB
K-Planes [17] 300 31.63 - - 110 min 200MB >100GB
MixVoxels-T [45] 300 30.71 0.024 0.162 15 min 500MB >100GB
4DGS [54] 300 32.01 - 0.055 6 hours 4000MB <10GB
HyperReel [1] 50 31.1 - 0.096 9 hours 360MB 18GB
NeRFplayer [43] 300 30.29 - 0.152 5.5 hours - >100GB
online StreamRF [27] 1 28.26 0.029 0.226 75 min 5000MB <10GB
HexPlane-PR 10 24.03 0.081 0.244 100 min 318MB 14GB
HexPlane-GR 10 25.17 0.050 0.272 125 min 318MB 14GB
INV [48] 1 28.72 - 0.121 40 hours 336MB <10GB
CD-NGP (Ours) 10 30.23 0.029 0.198 75 min 113MB 14GB
4.3.1 Standard-length sequences.
Table 1 shows the quantitative comparisons between CD-NGP and other state-of-the-art methods on DyNeRF dataset. We use the officially released metrics. [28, 34] are trained on 8 V100 GPUs or equally 4 A100 GPUs. [28, 34, 43] have no official implementations available. We are unable to obtain the missing quality metrics in [17] because of their huge memory occupation in the default resolution. [7] uses resolution $(1024,768)$ in their implementation, so the memory occupation is much smaller than other offline methods. We use the released output of [27] for DSSIM and LPIPS metrics. [27, 48] are methods trained per-frame. [43] is trained on the whole video sequence but streamable upon inference. 4DGS [54] caches the frames onto disk for training to alleviate memory occupation with a trade-off on training speed. The reconstruction quality is measured by PSNR, DSSIM, and LPIPS [57] (by AlexNet [24]) following common practices [28, 7, 45]. To prove the viability of our parameter-isolation-based method, we apply the previous state-of-the-art method HexPlane [7] to continual learning setting with parameter-regularization (HexPlane-PR) and generative replay (HexPlane-GR) as additional baselines.
(a) HexPlane-offline
(b) MixVoxels-T
(c) StreamRF
(d) HexPlane-PR
(e) HexPlane-GR
(f) CD-NGP (ours)
Figure 3: Comparison of reconstruction quality on the most challenging flame salmon scene in DyNeRF dataset. PSNRs are computed under the resolution $1352\times 1014$.
We compare our method with both online and offline baselines. As demonstrated in Tab. 1, our CD-NGP is 32×32$\times$32 × faster than the MLP-based continual representation [48], $\times4$ smaller than the voxel-based continual representation [27] and significantly outperforms the two additional baselines in the continual learning setting. The undesirable results from HexPlane-PR and HexPlane-GR show the limitations of the regularization-based and replay-based methods in dynamic scenes: unlike the static scenes, the objects are not stationary across time and the models with limited complexity are not capable of reconstructing unlimited transient effects. Benefiting from the parameter isolation like previous methods [27, 48], our method achieves fast convergence and compact model size simultaneously: it leverages the hash tables to ensure fast convergence and reuses the features from differently-configured spatial hash tables to ensure high scalability. Compared with prevailing offline methods [28, 34, 7, 17, 45, 54], our method only requires caching $T_{chunk}=10$ frames into buffer by continual learning and thus reduces memory occupation by $85\%$. Our method also achieves $40\%$ less model size ($0.4$MB/frame) by leveraging the feature fusion of the base and the auxiliary branches with comparable quality.
Figure 3 provides qualitative comparisons. Compared with offline methods in Fig. 3(a), Fig. 3(b) and the online method in Fig. 3(c), our model is much smaller without large quality loss. Compared with regularization-based and replay-based continual learning methods in Fig. 3(d) and Fig. 3(e) respectively, our method achieves significantly higher quality. Results of other scenes including depth maps are shown in Fig. 4. As illustrated in Fig. 6, our method reaches a new balance between rendering quality, model size, and training time while remaining memory-friendly, in which we use metrics on the flame salmon scene for fair comparison. The perceptual consistency is measured by $1-$ LPIPS, and the speed is the reciprocal of the training time. The compactness and RAM-Efficiency are mapped by $-log$ function from the model size and memory occupation respectively.
Figure 4: The results of the proposed CD-NGP on DyNeRF dataset.
4.3.2 Long sequences.
We also provide quantitative comparisons on our proposed long multi-view video dataset in Tab. 3 and Fig. 6 to further show the advantage in scalability of our method. We split the sequence into 300-frame chunks for the compared methods since the offline baseline [7] is designed for 300-frame sequences in the DyNeRF dataset. We set the hash table size as $2^{12}$ in the auxiliary branches and use $T_{chunk}=5$ in our model. The offline baseline [7] produces unstable results on one of the 5 scenes in our dataset, so we only use the other 4 scenes to obtain the metrics of [7]. As shown in Tab. 3, our method reduces memory and storage consumption by at least 4×4$\times$4 × with comparable quality to the offline baseline. Compared with the results in Tab. 1, the performance gap in storage consumption is even larger in long video sequences, because the auxiliary branches consume the majority of the parameters and the complexity in storage is reduced to $\mathcal{O}(N_{chunk}\cdot 2^{12}LF)$. The PSNR curves in Fig. 6 further demonstrate the stability of our method. The offline baseline produces the best quality. The compared online methods suffer from catastrophic forgetting: the PSNRs at the beginning of the chunks are much lower than at the end. Our method resists the problem by parameter isolation and ensures the quality of the whole sequence. We also provide comparisons between rendered results in Fig. 8. The regularization method in Fig. 7(a) reconstructs the TV screen in the background erroneously. The replay method in Fig. 7(b) fails to recover the details. Ours in Fig. 7(d) produces comparable quality to the offline baseline in Fig. 7(c).
Figure 5: Comparison between different performance indexes of our method and recent representative baselines.
Figure 6: PSNR curves of the seminar scene. The solid line and dotted line denotes offline and online methods respectively.
4.4 Ablation studies
(a) HexPlane-PR
(b) HexPlane-GR
(c) HexPlane
(d) CD-NGP (ours)
Figure 7: Comparison of reconstruction quality on the 19th frame of the seminar scene in our long video dataset.
Figure 8: PSNR curves of asymmetric (solid line, ours) and symmetric (dotted line) hash table designs on the sear steak scene.
4.4.1 Lengths of hash tables.
Benefiting from the multi-branch hash table design, our model is highly efficient and configurable. We change the parameter numbers of the initial branch and the auxiliary branches and obtain the results on the DyNeRF dataset in Tab. 3. $(P_{1},P_{2})$ are the log2 values of the hash table lengths of the base branch and the other branches. Please note the whole model sizes include the sizes of the occupancy grid and are thus larger than those deduced from parameter numbers. Larger hash tables alleviate hash collisions and lead to better results. However, the default configuration $(19,14)$ in CD-NGP produces better results than the naive $(16,16)$ configuration even if the storage consumption is $40\%$ lower. Moreover, the $(19,12)$ configuration of CD-NGP significantly outperforms the naive $(14,14)$ configuration at a similar storage cost. To further demonstrate the scalability of our method, we provide the PSNR curves of different hash table designs on the sear steak scene in Fig. 8. Equipped with a larger hash table in the base branch, the asymmetric setting $(19,14)$ and $(19,12)$ outperforms others in the beginning and produces comparable metrics to the naive $(16,16)$, $(14,14)$ configuration respectively in the subsequent chunks. The advantage comes from our asymmetric hash table designs: the base branch with a large hash table recovers more environment patterns, and the auxiliary branches could thus use fewer parameters to cope with recent changes only.
Table 2: Comparisons between our CD-NGP and baselines on the long video sequences in our dataset. denotes the method trained offline. Metrics are measured on the first 1200 frames. Compared baselines are trained with a chunk size of 300 frames.
[…]
HexPlane-PR 20.90 0.286 1272MB 12GB
HexPlane-GR 24.82 0.282 1272MB 12GB
CD-NGP (ours) 26.50 0.201 201MB 12GB
HexPlane [7] 28.32 0.117 800MB 74GB
Table 3: Ablation study on hash table sizes. denotes the default setting.
[…]
$(18,18)$ 151M 30.55 638MB
$(16,16)$ 41M 30.08 191MB
$(14,14)$ 12M 29.32 75MB
$(19,14)^{{\dagger}}$ 21M 30.23 113MB
$(16,14)$ 13M 29.78 79MB
$(19,12)$ 14M 30.09 82MB
4.4.2 Spatial representations.
Table 4: Ablation study on spatial representations. Different representations lead to subtle changes in quality metrics.
[…]
CD-NGP 21M 30.23 0.198 75min 113MB
Voxel-C 21M 30.52 0.174 80min 114MB
Plane 14M 29.16 0.215 77min 83MB
MERF 7.5M 29.93 0.201 109min 57MB
MERF-C 7.5M 30.22 0.174 115min 57MB
Table 5: Ablation study on regularization weight $\lambda_{r}$. $10^{-3}$ is the default value.
[…]
$10^{-5}$ 30.20 0.0269 0.198
$10^{-4}$ 30.18 0.0268 0.196
$10^{-3}$ 30.23 0.0269 0.198
$10^{-2}$ 30.09 0.0279 0.197
1 29.66 0.0313 0.205
The proposed hash-based pipeline is not only practical with voxel hash grids but also compatible with other hash-based spatial backbones. We provide quantitative results on plane representation like [17, 21] and hybrid representation (MERF) [38] on the DyNeRF dataset in Tab. 4. Since the element-wise product for plane representation in [17] produces unsatisfactory results due to the hash collision, we remove the $P_{m}(x,y,z)$ in Eq. 6 and fuse the features $Q_{mij}(i,j)$ by concatenation in the implementation.
Table 4 also shows metrics of different feature fusion methods. We replace the element-wise sum in Eqs. 7 and 6 with vector concatenation and denote them as Voxel-C and MERF-C in Tab. 4. The hybrid representations (MERF and MERF-C) achieve comparable model sizes ($\sim$60MB) to methods for static scenarios [31, 21] without significant quality loss, reducing the average bandwidth cost to $0.2$MB/frame. Benefiting from larger feature dimensions, the vector concatenation leads to slightly better quality (especially when measured in LPIPS) by more training time, and the reasonable trade-off also shows the robustness of our method. Compared with the hash table designs in Tab. 3, the Voxel-C representation in Tab. 4 almost performs as well as the bulky $(18,18)$ configuration in Tab. 3 and reduces the model size by $80\%$.
We also provide ablation studies on the regularization weights of auxiliary spatial encodings. We apply regularizations on $\Psi_{k}(x)$ with different $\lambda_{r}$ and show the results in Tab. 5. The regularization is of slight improvement, and our method is robust within a wide range of $\lambda_{r}$. Please refer to the supplementary material for more ablation studies.
5 Conclusion and Limitation
We propose CD-NGP, a continual learning framework for multi-view dynamic scenes based on parameter isolation. Equipped with the multi-branch hash encoding structure, it shows high efficiency and scalability and is capable of reconstructing long video sequences with low host memory occupation. Compared with offline methods, it consumes $40\%$ less bandwidth and $85\%$ less memory than recent explicit baselines and achieves comparable quality while maintaining the speed advantage over the pure-MLP-based methods. Compared with online baselines, it significantly outperforms replay and regularization-based continual learning methods and reaches a new balance among reconstruction quality, training speed, bandwidth occupation, and memory cost. However, like most of the NeRF-based methods, it still requires querying the field function and is not appropriate for achieving real-time rendering. Further efforts could be made to apply baking methods or Gaussian point clouds to continual learning to achieve real-time rendering in long video sequences for wide applications.
Appendix 0.A Detailed Results on DyNeRF Dataset
We list the per-scene results of the proposed continual learning method with different spatial representations on the DyNeRF dataset in Tab. 6.
Table 6: Detailed results of the proposed CD-NGP and other hash-based methods on the DyNeRF dataset.
Model Flame Salmon Cook Spinach
[…]
CD-NGP 26.37 0.04586 0.2886 30.76 0.02440 0.1793
Voxel-C 26.65 0.04243 0.2680 31.03 0.02288 0.1593
Plane 25.48 0.06058 0.3173 29.41 0.03487 0.2010
MERF 26.15 0.04952 0.2921 30.35 0.02931 0.1858
MERF-C 26.56 0.04435 0.2546 30.69 0.02609 0.1583
Model Flame Steak Sear Steak
[…]
CD-NGP 31.27 0.02132 0.1736 31.71 0.01917 0.1701
Voxel-C 31.64 0.01823 0.1403 32.02 0.01698 0.1424
Plane 30.26 0.02949 0.1804 30.56 0.02723 0.1811
MERF 31.01 0.02508 0.1719 31.44 0.02181 0.1698
MERF-C 31.13 0.02424 0.1506 31.66 0.01989 0.1413
Model Cut Roasted Beef Average
[…]
CD-NGP 31.03 0.02366 0.1801 30.228 0.026882 0.19834
Voxel-C 31.26 0.02205 0.1578 30.520 0.024514 0.17356
Plane 30.11 0.03302 0.1951 29.164 0.037038 0.21498
MERF 30.71 0.02765 0.1856 29.932 0.030674 0.20104
MERF-C 31.04 0.02450 0.1632 30.216 0.027814 0.17360
Appendix 0.B More Implementation Details
0.B.1 Parameter numbers and streaming efficiency.
We set the dimensions of the latent feature $\mathbf{H}$ in Eq. 3 as 48. We use MLPs with two hidden layers and the dimensions of the hidden layers in Eqs. 3 and 4 are set as 128, and 64 respectively. We use MLPs with identical structures for different branches. We set the resolutions of the hash encoders of the proposed CD-NGP and MERF representations within $(16,2048)$ times the scale of the scene, and we use $(64,2048)$ for the plane representation.
Following MERF [38] for static scenes, we use a smaller hash table for the MERF and the MERF-C than pure voxel hash grid [31] representation in Tab. 4 to ensure the compactness and scalability of our method. Denoting the lengths of hash tables along pure voxels in a branch of CD-NGP as $2^{P}$, the lengths of hash tables in the MERF and the MERF-C are set as $2^{P-3}$ and $2^{P-4}$ for voxel and plane encoders respectively. The number of the parameters of the MERF representation is thus reduced to $1-2^{-3}-3\times 2^{-4}=31.25\%$ of the voxel hash grids. Since the hash tables take up the majority of the parameters, the model size is reduced from $21M$ to $7.5M$ as in Tab. 4. We use $2^{P-2}$-length hash tables for the Plane representation in Tab. 4 to balance the quality and the model size. The structural configurations $(L,F)$ of the plane representation in Tab. 4 are set as $(6,4)$ for better results.
The detailed parameter numbers and bandwidth consumption are thus listed in Tab. 7. 3D Enc. and 2D Enc. denote the spatial encodings along voxels and planes respectively. The tuples $(M_{1},M_{2})$ in the cells denote the parameter numbers of the base encoder and each of the auxiliary encoders. Following [27, 43, 1, 48] we also report the minimum per-frame bandwidth consumption $\mathbf{B_{min}}$, which only depends on the parameter numbers of the auxiliary branches and is thus significantly lower than the average bandwidth consumption $\mathbf{B_{avg}}$. Since the base encoder and the density grid could be shared among different branches and transmitted only once in the beginning, the minimum bandwidth consumption is reduced to $0.065\sim 0.16$ MB per frame, which is 100×100$\times$100 × less than the voxel-based StreamRF [27] and 7×7$\times$7 × less than the pure MLP representation INV [48].
Table 7: The parameter numbers of each module of the 30-branch representations in Tab. 4. $\mathbf{B_{min}}$ and $\mathbf{B_{avg}}$ are minimum and average bandwidth costs in MB/frame. ##\## denotes the number of parameters per branch (10 frames).
Model ##\##3D Enc. ##\##2D Enc. ##\##MLP ##\##Online ##\##Total $\mathbf{B_{min}}$ $\mathbf{B_{avg}}$
CD-NGP $(9.9M,0.37M)$ $0.042M$ $0.39M$ $21M$ 0.16 0.38
Voxel-C $(9.9M,0.37M)$ $0.046M$ $0.40M$ $21M$ 0.16 0.38
Plane - $(1.9M,0.082M)$ $0.048M$ $0.28M$ $14M$ 0.11 0.28
MERF $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.042M$ $0.16M$ $7.5M$ 0.061 0.19
MERF-C $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.052M$ $0.17M$ $7.5M$ 0.065 0.19
0.B.2 Continual learning pipeline.
We further illustrate the whole process of our method in Algorithm 1, where $\eta_{init}$ and $\eta_{aux}$ denote the numbers of iterations of the initial branch and the auxiliary branches, respectively. $\Psi_{i},\Gamma_{i},\theta_{1}^{(i)},\theta_{2}^{(i)}$ denote the model parameters in the branches. The whole model is trained chunk by chunk. The implementation will be available upon publication.
Algorithm 1 Continual learning pipeline
0:  $N_{chunk},T_{chunk},\eta_{init},\eta_{aux},\Psi_{i},\Gamma_{i},\theta_{1}^{(i)},\theta_{2}^{(i)},i\in\mathbb{N}$
0:  The whole representation for dynamic scenes.
[…]
2:  while $N_{chunk}$ do
[…]
5:     if $k~{}\mathbf{mod}~{}T_{episode}~{}=0$  then
[…]
7:     else
[…]
10:     end if
11:     Train the $k^{th}$ model branch $(\Psi_{k},\Gamma_{k},\theta_{1}^{(k)},\theta_{2}^{(k)})$ for $\eta$ steps.
[…]
13:  end while
Appendix 0.C More Details about the Long Video Dataset
We collect the dataset for evaluating the scalability of our method and further research. We have checked the videos to make sure they do not contain controversial or offensive content before using them for research purposes. The faces in the videos may contain identifiable information, but the information will neither be used for any other purposes nor affect the anonymity of the submission. All videos are collected with the explicit consent of all the participants. The view in the upper center is selected for validation. We synchronize the videos by environmental sound, and we put chessboard images in the scene to help calibrate the cameras and estimate the camera poses.
Appendix 0.D More Ablation Studies
0.D.1 Temporal representations.
Our continual learning pipeline is also compatible with a variety of representations for the time axis. Currently, there are 2 other prevailing representations for the time axis alone: pure frequency encoding [36] and MLP with frequency-encoding [16]. We replace the temporal hash encoding in our method with them on DyNeRF dataset and show the results in Tab. 9. Our time latent code slightly outperforms others on DyNeRF dataset. The robustness against changes in the temporal representations also shows the viability of the continual learning method.
Table 8: Ablation study on temporal representations.
[…]
CD-NGP 30.23 0.0269 0.198
Time-MLP 30.20 0.0273 0.202
Time-Freq 30.19 0.0272 0.203
Table 9: Ablation study on field composition methods.
[…]
CD-NGP 30.23 0.0269 0.198
Full composition 29.24 0.0367 0.228
Static-dynamic 28.27 0.0568 0.291
0.D.2 Field composition methods.
We also conduct ablation studies on the field composition methods and show the results in Tabs. 9 and 9. The field composition process is defined as:
$(\sigma,\mathbf{c})=(\sigma_{i}+\sigma_{a},\mathbf{c}_{i}+\mathbf{c}_{a}).$ (13)
Full composition in Tab. 9 denotes the naive composition of 2 fields from the base branch and the current branch respectively, i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ are the field from the base branch and $\sigma_{a}$ and $\mathbf{c}_{a}$ in Eq. 13 are the field from the current branch. Both branches are encoded with the same architecture as Eq. 3. Static-dynamic in Tab. 9 denotes a naive composition of the field from a static branch by the base encoder and a dynamic branch by the auxiliary encoders similar to [45], i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ in Eq. 13 are encoded without the $\Gamma(t)$ in Eq. 3. The online training strategies remain the same. Our CD-NGP outperforms the other methods in all metrics. Furthermore, as compared in Fig. 9, the compared composition methods in Figs. 9(c) and 9(d) suffer from salt and pepper noise because of the explicit composition, while our CD-NGP circumvents the problem by fusing features in the latent space and produces better results.
(a) Ground truth
(b) CD-NGP (Ours)
(c) Full composition
(d) Static-dynamic
Figure 9: Comparison of different field composition methods.
0.D.3 Chunk size.
We also conduct ablation studies on the chunk sizes and show the results in Tabs. 11 and 11. We only report the results on the cook spinach scene in the DyNeRF dataset for brevity. Our method is robust within a wide range of chunk sizes. As in Tab. 11, the smaller chunk size leads to better results when controlling the iteration numbers of each chunk. When the total iterations of the whole 300-frame sequence are held constant, the larger chunk size leads to better results as in Tab. 11. We do not cherry-pick the best chunk size and select $T_{chunk}=10$ as the default configuration in our method. The selection of the chunk size is a trade-off between the quality, robustness, training time, and host memory occupation and should depend on the data.
Table 10: Ablation study on chunk sizes. The iteration numbers of each chunk are held constant under different $T_{chunk}$ settings.
[…]
30 30.31 0.02654 0.1842 80MB
15 30.54 0.02557 0.1816 97MB
10 30.76 0.02440 0.1793 113MB
5 30.99 0.02268 0.1674 162MB
Table 11: Ablation study on chunk sizes. The total iterations of the whole 300-frame sequence are held constant under different $T_{chunk}$ settings.
[…]
30 30.84 0.02392 0.1505 80MB
15 30.76 0.02413 0.1684 97MB
10 30.76 0.02440 0.1793 113MB
5 30.39 0.02602 0.2075 162MB

### v3

@@
 Abstract
-We present CD-NGP, which is a fast and scalable representation for 3D reconstruction and novel view synthesis in dynamic scenes.
+We present CD-NGP, a fast and scalable representation for 3D reconstruction and novel view synthesis in dynamic scenes.
 Inspired by continual learning, our method first segments input videos into multiple chunks, followed by training the model chunk by chunk, and finally, fuses features of the first branch and subsequent branches.
@@
 Furthermore, we conduct experiments on a self-collected extra-long video sequences dataset, which demonstrates the superior scalability of our method over previous state-of-the-art approaches.
+A free copy of this paper and all supplemental materials are available at https://arxiv.org/abs/2409.05166.
 keywords:
-Continual learning Dynamic novel view synthesis Neural Radiance Field.
-1 Introduction
+Continual learning, Dynamic novel view synthesis, Neural Radiance Field.
+[…]
+The demonstration of our continual learning pipeline for dynamic scenes.
+Please note the dimensions of representations for different branches remain the same while the complexities of hash tables are flexible and could be configured to ensure scalability.
+The spatial hash features could be encoded along voxels [31], planes [17] or both [38].
+Introduction
 Neural radiance field (NeRF) [30] has made significant progress in 3D scene reconstruction and novel view synthesis.
@@
 Despite significant progress having been made in this field, the above-mentioned methods assume that the scene is static and may not work properly in dynamic scenes.
-To address the dynamic reconstruction problem, amount of efforts have been made under the offline setting, requiring access to the entire dataset in the whole training stage.
-For example, [36, 28, 34, 33, 15, 29, 44] propose to use MLP-based representation and achieve high-quality results.
-[40, 17, 7, 16, 45] speed up the optimization by using tensor factorization and voxel representation, respectively.
+To address the dynamic reconstruction problem, a number of efforts have been made in the offline setting, requiring access to the entire dataset during the training stage.
+For example, the methods in [36, 28, 34, 33, 15, 29, 44] propose to use MLP-based representation and achieve high-quality results.
+There are also methods speed up the optimization by using tensor factorization ([40, 17, 7]) and voxel representation [16, 45].
 The more recent continual-learning-based methods [27, 48] can train models frame by frame, and streamable representations [43] enable the model to be loaded on the fly at inference.
@@
 As a result, these methods are limited in scalability, making them unsuitable for long videos.
-In this paper, we propose Continual Dynamic neural graphics primitives (CD-NGP), a novel, fast, and scalable representation for reconstructing dynamic scenes, which enables dynamic reconstruction in long videos.
-As illustrated in Fig.
-1, we evenly split the dynamic videos into multiple chunks based on the timestamp and conduct reconstruction with different model branches, where each branch includes MLP regressors, spatial encoding for $(x,y,z)$, and temporal encoding for $t$.
-We encode the different chunks by hash tables with different complexities to ensure model scalability and then continually train the model chunk by chunk to alleviate memory occupation.
+In this paper, we propose continual dynamic neural graphics primitives (CD-NGP), a fast and scalable representation for reconstructing dynamic scenes, which enables dynamic reconstruction in long videos.
+As illustrated in CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes, we evenly split the dynamic videos into multiple chunks based on the timestamp and conduct reconstruction with different model branches, where each branch includes MLP regressors, spatial encoding for the spatial coordinates $(x,y,z)$, and temporal encoding for the time stamp $t$.
+We encode the different chunks by hash tables with different complexities to ensure the model scalability and then continually train the model chunk by chunk to alleviate memory occupation.
 For each individual chunk, we locate the corresponding model branch, sample 3D points $(x,y,z)$ according to frame pixels and cached occupancy grids [31], and represent spatial features by spatial hash grids.
@@
 Spatial and temporal features are concatenated along with viewing direction $d$ and input to MLP regressors for rendering and optimization.
-Extensive experiments demonstrate the great scalability of our method: compared with prevailing offline baselines, our method consumes $85\%$ less host memory at the training phase, $40\%$ less model size at the inference phase, and delivers comparable results.
+We validate the effectiveness of our method through comprehensive experimental results on both the widely used public datasets and the dataset proposed by us.
+As illustrated in Fig.
+1, our method reaches a new balance between rendering quality, model size, and training time while remaining memory-friendly, in which we use metrics on the flame salmon scene in the DyNeRF dataset for fair comparison.
+The perceptual consistency is measured by $1-$ LPIPS by AlexNet [57, 24], and the speed is the reciprocal of the training time.
+The compactness and RAM-Efficiency are mapped by $-log$ function from the model size and memory occupation respectively.
+Figure 1: Comparison between different performance indexes of our method and recent representative baselines.
+Compared with prevailing offline baselines, our method consumes $85\%$ less host memory at the training phase, $40\%$ less model size at the inference phase, and delivers comparable results.
 Compared with online baselines, our method achieves compact model size, high quality, and fast convergence simultaneously.
-The main contribution of this work is summarized as 3 folds:
-Figure 1: The demonstration of our continual learning pipeline for dynamic scenes.
-Please note the dimensions of representations for different branches remain the same while the complexities of hash tables are flexible and could be configured to ensure scalability.
-The spatial hash features could be encoded along voxels [31], planes [17] or both [38].
-  • We propose a novel method combining spatial and temporal hash encodings for dynamic scenes in short intervals with high scalability.
-  • We propose a fast continual learning framework for dynamic scenes, producing comparable results to offline state-of-the-art methods while reducing memory and storage occupation significantly.
-  • We provide a dataset of extra-long video sequences ($>3$ minutes) and further demonstrate the scalability of our method on long multi-view videos.
+Contributions of this paper are three-fold as follows:
+  • We propose a novel method combining spatial and temporal hash encodings for dynamic scenes and yield comparable results to both offline and online methods.
+  • We propose a fast continual learning framework for dynamic scenes and reduce memory and storage occupation significantly.
+  • We provide a dataset of multi-view long video sequences ($>3$ minutes) with large motion and further demonstrate the scalability of our method on the proposed dataset.
 1 Related Work
@@
 Hedman et al.
-[20] propose the baking method to accelerate the rendering process.
+[20] propose the baking method (converting the MLP representation into explicit meshes) to accelerate the rendering process.
 They alpha-compose the cached features along the ray and use MLPs only once per ray to overcome the computation load of the per-point MLP querying in NeRF.
@@
 1.2 Representations of dynamic scenes
-There are mainly 3 variants of NeRF for dynamic scenes: MLP, tensor factorization, and voxel representations.
+There are mainly 3 variants of NeRF for dynamic scenes.
+The first class is pure MLP-based representation.
 D-NeRF [36] leverages a large deformation MLP to model the transition and the trajectories of the objects and extends NeRF to simulated dynamic scenes.
@@
 [34] propose an interpolation-based MLP representation for dynamic scenes and achieve significant acceleration over DyNeRF.
+The second class is tensor factorization-based representation.
 HyperReel [1] proposes a network to predict the sample locations and represent dynamic scenes by key-frame-based TensoRF [9] representation to ensure high fidelity and real-time rendering without any custom CUDA kernels.
-K-Planes [17] and HexPlane [7] factorize the whole 4D field to the product of 2D planes.
-Though differently implemented, they both achieve high-fidelity results with intermediate model size (200MB) and training time (100 minutes) on DyNeRF dataset.
+K-Planes [17] and HexPlane [7] factorize the whole 4D field into the product of 2D planes.
+Though differently implemented, they both achieve high-fidelity results with intermediate model size (200MB) and training time (100 minutes) on the DyNeRF dataset.
+The third class is the explicit voxel-based representations.
 TineuVox [16] leverages the deformation MLP in D-NeRF and implements a 3D voxel grid with multi-scale interpolation.
-It achieves high fidelity and fast convergence on D-NeRF dataset.
+It achieves high fidelity and fast convergence on the D-NeRF dataset.
 MixVoxels [45] leverages the standard deviation across the time of all pixels to estimate a variation field to decompose the dynamic scene into static and dynamic voxels.
@@
 INV [48] uses a shared color MLP and trains different structure MLPs frame by frame to form a continual NeRF representation and achieves high-fidelity results in dynamic scenes.
-3 Methodlogy
+2 Method
 2.1 Preliminaries
@@
 where the function $F$ denotes the dynamic NeRF model and $\Theta$ denotes trainable model parameters.
-Let $I_{i}$ denote the $i^{th}$ input image corresponding to observation direction $d_{i}$, $\mathbf{V}=\{\nu_{i}~{}|~{}\nu_{i}=(I_{i},d_{i}),1\leq i\leq N_{view}\}$ is the set of the training views, where $i,N_{view}\in\mathbb{Z^{+}}$ and $\mathbf{D_{a}}$ is the readily-accessible data at one step.
-Previous methods [6, 35] for static scenes define the continual learning problem as learning a unified representation with a limited subset of views $V_{a}$ available at every optimization step, i.e.
-$\mathbf{D_{a}}=V_{a}\subsetneq V$.
-However, a more natural definition of continual NeRF is in the dynamic scenes, where the observations at different time stamps arrive continually.
-Denoting all the frame times as $\mathbf{T}=\{\tau_{i}~{}|~{}1\leq i\leq N_{time}\},i,N_{time}\in\mathbb{Z^{+}}$, we define the continual learning problem as synthesizing a video sequence when the model could only have access to a limited number of frames $\mathbf{T_{a}}\subsetneq\mathbf{T}$ from all views $\mathbf{V}$ at each step, i.e.
-$\mathbf{D_{a}}=\{\mathbf{V}\times\mathbf{T_{a}}\}$ where $\times$ is the Cartesian product of sets.
-To simplify the notations, let the whole sequence of frames be evenly divided into $k$ chunks where a chunk $\mathbf{T_{k}}=\{\tau_{i}~{}|~{}\lfloor\frac{N_{time}}{T_{chunk}}\rfloor=k\}$ is a sequential subset of length $T_{chunk}$.
-We define the continual learning problem as learning a representation where there are only the frames in the current chunk available at each optimization step, i.e.
+[…]
+To simplify the notations, let the whole sequence of frames be evenly divided into $k$ chunks where a chunk $\mathbf{T_{k}}=\{\tau_{i}\leavevmode\nobreak\ |\leavevmode\nobreak\ \lfloor\frac{N_{time}}{T_{chunk}}\rfloor=k\}$ is a sequential subset of length $T_{chunk}$.
+We define the continual learning problem as learning a representation where only the frames in the current chunk are available at each optimization step, i.e.
 $\mathbf{T_{a}}=\mathbf{T_{k}}$.
-Accordingly, $T_{chunk}=1$ denotes training the model frame by frame, $T_{chunk}=10$ denotes training the model with a buffer of 10 frames, $T_{chunk}=300$ denotes training the entire 300-frame sequence offline.
+Accordingly, $T_{chunk}=1$ denotes training the model online in a frame-by-frame manner, $T_{chunk}=10$ denotes training the model online with a buffer of 10 frames.
+If the video contains 300 frames in total, the $T_{chunk}=300$ denotes training the entire 300-frame sequence offline.
 The definition is close to real-world applications since caching all training frames in memory or storing them on the disk is limited in scalability for long videos.
 2.2 Representation of short dynamic intervals
-As pointed out by [46], naive 4D hash grids are not suitable for high-fidelity dynamic view synthesis.
+As shown in [46], naive 4D hash grids for dynamic scenes are short in reconstruction quality.
 To achieve fast reconstruction and slim model size simultaneously, we represent time and spatial features separately and concatenate them to alleviate feature confusion.
 The spatial hash features may be encoded along 3D voxels [31], 2D planes [17, 21], or both [38].
-The time feature is encoded per frame following [28].
+The time feature is encoded by short hash tables in a per-frame manner.
 Following previous state-of-the-art methods [30, 31, 16], we apply a two-fold field function for geometry and color individually.
@@
 $\mathbf{c}=f_{\theta_{2}}(\mathbf{H},\phi,\psi).$ (4)
-where $\Psi(\mathbf{x})$ denotes configurable spatial hash encodings, $f_{{\theta_{1}}}$ is a tiny MLP with trainable parameter ${\theta_{1}}$, $\Gamma(t)$ is the hash encoding for time axis alone, $\oplus$ is the concatenation of vectors.
+where $\mathbf{x}=(x,y,z)$ is the spatial coordinates, $\Psi(\mathbf{x})$ denotes configurable spatial hash encodings, $f_{{\theta_{1}}}$ is a tiny MLP with trainable parameter ${\theta_{1}}$, $\Gamma(t)$ is the hash encoding for time axis alone, and $\oplus$ is the concatenation of vectors.
 The MLP $f_{{\theta_{1}}}$ in Eq.
@@
 The spatial encoders $\Psi(x)$ follow a multi-resolution concatenated format, defined as:
-$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}~{}P_{m}(x,y,z)$ (5)
-where $P_{m}(x,y,z)\in\mathbf{R^{F}}$ denotes the tri-linearly interpolated value from the nearby 8 grid vertices w.r.t.
-3D coordinates $(x,y,z)$ at the $l^{th}$ resolution in Instant-NGP [31], $L$ is the number of the resolutions.
-$\Psi(x)$ may also be configured as the fusion of 2D and 3D features in MERF [38], defined as:
-$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}~{}P_{m}(x,y,z)+Q_{mxy}(x,y)+Q_{myz}(y,z)+Q_{mzx}(z,x)$ (6)
-where $Q_{mij}(i,j)$ denotes the bi-linearly interpolated value from the 4 nearby vertices in the orthogonally projected planes w.r.t.
-axes $(i,j)$ at the $m^{th}$ resolution.
-Removing the $P_{m}(x,y,z)$ term leads to a pure plane representation like [17, 21].
-3.3 CD-NGP
+$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}\leavevmode\nobreak\ P_{m}(x,y,z)$ (5)
+where the $P_{m}(x,y,z)\in\mathbf{R}^{F}$ denotes the tri-linearly interpolated value from the nearby 8 grid vertices w.r.t.
+3D coordinates $(x,y,z)$ at the $l$-th resolution, $L$ is the number of the resolutions, and $F$ is the dimensions of features $P_{m}(x,y,z)$, which are similar to those in Instant-NGP [31].
+2.3 Continual dynamic neural graphics primitives (CD-NGP)
 Unlike the existing replay-based methods for static scenarios [6, 35], we solve the continual learning problem of dynamic novel view synthesis via parameter isolation.
-The key insight is that the current state-of-the-art methods by tensor factorization [7, 17] or explicit voxels [45] are not accurate estimations of the object trajectory and inherent dynamics, but only simple regressions of the optical attributes in 3D space within a limited time.
-Therefore, the replay method is of limited use because the new transient processes are constantly arriving and the model of a limited size is not capable of reconstructing unlimited transient effects and thus suffers from catastrophic forgetting.
+The current state-of-the-art methods by tensor factorization [7, 17] or explicit voxels [45] are regressions of the optical attributes in 3D space within a limited time instead of the object trajectory and inherent dynamics.
+Therefore, the replay method is of limited use in our defined scenario: Since new transient processes are constantly arriving, the model of a limited size is not capable of reconstructing unlimited transient effects.
+Thus, the replay-based method suffers from catastrophic forgetting.
 To address this issue, we propose our continual dynamic neural graphics primitives (CD-NGP) based on parameter isolation.
 We split the model into $N_{chunk}$ branches to represent different chunks of the dynamic scene accordingly.
-The $k^{th}$ branch contains spatial hash encoder $\Psi_{k}(\mathbf{x})$, temporal hash encoder $\Gamma_{k}(t)$, and MLPs $f_{\theta_{1k}},f_{\theta_{2k}}$.
+The $k$-th branch contains spatial hash encoder $\Psi_{k}(\mathbf{x})$, temporal hash encoder $\Gamma_{k}(t)$, and MLPs $f_{\theta_{1k}},f_{\theta_{2k}}$.
 We train the first branch (base branch) with a large spatial hash table to better reconstruct the scenes with initial movements and static objects.
@@
 We thus substitute the $\Psi(\mathbf{x})$ in Eq.
-3 with the fused feature to obtain the volume density $\sigma$ at the $k^{th}$ model branch under the continual learning setting as:
+3 with the fused feature to obtain the volume density $\sigma$ at the $k$-th model branch under the continual learning setting as:
 $(\sigma,\mathbf{H})=f_{\theta_{1k}}((\Psi_{0}(\mathbf{x})+\Psi_{k}(\mathbf{x}))\oplus\Gamma_{k}(t)),$ (6)
@@
 The scalability of our method derives from the redundancy of dynamic scenes and the flexibility of hash table sizes.
-As pointed out by [45], most objects ($>90\%$) in the dynamic scenes are static.
-Therefore, the first chunk could be represented by a large hash table to reconstruct the scenes with initial movements and static objects, the subsequent chunks could be represented by small hash tables to compensate for errors, and the whole representation could be designed to leverage the fusion of features from the corresponding two branches to ensure scalability.
-To be specific, if we follow the default setting of Instant-NGP [31] and set the hash table in the first branch as $2^{P_{1}}=2^{19}$, the sizes of hash tables in subsequent branches could be configured as $2^{P_{2}}=2^{14}$, which consumes only $1/32$ of the parameters in the first branch.
-Since the spatial hash tables consume the most parameters in the model, and most chunks are encoded by the subsequent branches, the advantage of our asymmetric hash table structure is obvious: compared with the previous linear-streaming method [43], the total bandwidth consumption is reduced from $\mathcal{O}(\mathbf{K}\cdot 2^{19}LF)$ to $\mathcal{O}(N_{chunk}\cdot 2^{14}LF+2^{19}LF)$, where $\mathbf{K}$ is their bandwidth budget controlling the bitrates in streaming scenarios.
-If the bandwidth is extremely limited, the hash tables in our model could be arranged as Eq.
-6 to reduce the bandwidth cost additionally.
+As pointed out by [45], most points ($>90\%$) in the dynamic scenes are static.
+We thus leverage the shared static points across time and represent the first chunk with a large hash table and the subsequent chunks with small hash tables to ensure scalability.
+To be specific, if we follow the default setting of Instant-NGP [31] and set the hash table in the first branch as $2^{P_{1}}$, the sizes of hash tables in subsequent branches could be configured as $2^{P_{2}}$.
+Please note that $P_{1}$ is significantly larger than $P_{2}$ to ensure the scalability of the model.
+When ($P_{1},P_{2})=(19,14)$, the hash table in the second branch consumes only $1/32$ of the parameters in the first branch.
+Since the spatial hash tables occupy the most parameters in the model, and most chunks are encoded by the subsequent branches, our assymetric hash table design enables continual learning of dynamic scenes and high scalability in streaming scenarios simultaneously.
+The total bandwidth consumption of our method is $\mathcal{O}(N_{chunk}\cdot 2^{P_{2}}LF+2^{P_{1}}LF)$, while the previous method [43] requires offline training and consumes $\mathcal{O}(\mathbf{K}\cdot 2^{P_{1}}LF)$ bandwidth, where $\mathbf{K}$ ranging from $0.5$ to $16$ is their bandwidth budget controlling the bitrates in streaming scenarios.
+If the bandwidth is extremely limited, the hash tables in our model could be arranged as a combination of 3D representations and 2D representations like [38] to reduce the bandwidth cost additionally.
 2.4 Rendering and optimization
@@
 $+\frac{1}{3}\sum_{i}w_{i}^{2}(s_{i+1}-s_{i})$ (9)
-where $s_{i}$ denotes the normalized ray distance along the ray, $w_{i}=T_{i}\alpha_{i}$ is the weight of the $i^{th}$ sample. $L_{o}$ is the entropy on the opacity by [25], defined as:
+where $s_{i}$ denotes the normalized ray distance along the ray, $w_{i}=T_{i}\alpha_{i}$ is the weight of the $i$-th sample.
+$L_{o}$ in Eq. 8 is the entropy on the opacity by [25], defined as:
 $L_{o}=-o\log(o),\quad o=\sum_{j=1}^{N}T_{j}\alpha_{j}.$ (10)
-$L_{r}$ is L1 regularization on the spatial features in the current branch as:
+$L_{r}$ in Eq. 8 is the L1 regularization on the spatial features in the current branch as:
 $L_{r}=||\Psi_{k}(\mathbf{x})||_{1}.$ (11)
-We set $\lambda_{r}$ as 0.001 to encourage leverage of the base hash encoder. Based on [25], we set a larger $\lambda_{d},\lambda_{o}$ as 0.005 for alleviating foggy effects.
+We set $\lambda_{r}$ as 0.001 to encourage leverage of the base hash encoder. Based on [25], we set a larger $\lambda_{d},\lambda_{o}$ as 0.005 to alleviate foggy effects.
+2.5 Continual learning pipeline.
+We further illustrate the whole process of our method in Algorithm 1, where $\eta_{init}$ and $\eta_{aux}$ denote the numbers of iterations of the initial branch and the auxiliary branches, respectively.
+$\Psi_{i},\Gamma_{i},\theta_{1}^{(i)},\theta_{2}^{(i)}$ denote the model parameters in the branches.
+The whole model is trained chunk by chunk.
+The implementation will be available upon publication.
+Algorithm 1 Continual learning pipeline
+0:  $N_{chunk},T_{chunk},\eta_{init},\eta_{aux},\Psi_{i},\Gamma_{i},\theta_{1}^{(i)},\theta_{2}^{(i)},i\in\mathbb{N}$
+0:  The whole representation for dynamic scenes.
+[…]
+2:  while $N_{chunk}$ do
+[…]
+[…]
+5:     if $k\leavevmode\nobreak\ \mathbf{mod}\leavevmode\nobreak\ T_{episode}\leavevmode\nobreak\ =0$  then
+[…]
+7:     else
+[…]
+[…]
+10:     end if
+11:     Train the $k$-th model branch $(\Psi_{k},\Gamma_{k},\theta_{1}^{(k)},\theta_{2}^{(k)})$ for $\eta$ steps.
+[…]
+13:  end while
 3 Experiments
@@
 DyNeRF dataset contains 6 individual cooking scenes with varied illumination, food topological changes, and transient effects like flames.
-It contains five 300-frame videos and a 1200-frame video, with up to 21 synchronized camera views at the resolution of $2704\times 2028$ at 30 frames per second (FPS).
+It contains five 300-frame videos and one 1200-frame video, with up to 21 synchronized camera views at the resolution of $2704\times 2028$ at 30 frames per second (FPS).
+Meetroom dataset [27] is another dataset for dynamic view synthesis that contains 3 scenes captured by 13 synchronized cameras at $1280\times 720$ resolution.
+The lengths of the scenes are set as 300 frames and the refresh rates are set as 30 FPS.
 To further prove the viability of our method on long video sequences, we provide a new dataset containing 5 dynamic scenes at $3840\times 2160$ resolution, the lengths of which are at least 200 seconds.
 Similar to the DyNeRF dataset [28], scenes in our dataset are recorded by 20 GoPro cameras at 30 FPS.
+We synchronize the videos by environmental sound, and we put chessboard images in the scene to help calibrate the cameras and estimate the camera poses.
 Compared with those in the DyNeRF dataset, the dynamics are much harder, including large motions of human bodies and transient effects in the background.
-Please refer to the supplementary material for more details.
+We collect the dataset for evaluating the scalability of our method and further research.
+We have checked the videos to make sure they do not contain controversial or offensive content before using them for research purposes.
+The faces in the videos may contain identifiable information, but the information will neither be used for any other purposes except for research nor be disclosed to any third party.
+All videos are collected with the explicit consent of all the participants.
+The view in the upper center is selected for validation.
 Figure 2: Example frames sampled from the proposed long multi-view video dataset.
@@
 We report the model sizes with the occupancy grid under bfloat16 precision and parameters in float32 precision.
-Please see the supplementary material for more details.
+We set the dimensions of the latent feature $\mathbf{H}$ in Eq.
+3 as 48.
+We use MLPs with two hidden layers and the dimensions of the hidden layers in Eqs.
+3 and 4 are set as 128, and 64 respectively.
+We use MLPs with identical structures for different branches.
+We set the resolutions of the hash encoders of the CD-NGP within $(16,2048)$ times the scale of the scene, and we use $(64,2048)$ for the plane representation.
+Following MERF [38] for static scenes, we use a smaller hash table for the MERF and the MERF-C than pure voxel hash grid [31] representation in Table 5 to ensure the compactness and scalability of our method.
+Denoting the lengths of hash tables along pure voxels in a branch of CD-NGP as $2^{P}$, the lengths of hash tables in the MERF and the MERF-C are set as $2^{P-3}$ and $2^{P-4}$ for voxel and plane encoders respectively.
+The number of the parameters of the MERF representation is thus reduced to $1-2^{-3}-3\times 2^{-4}=31.25\%$ of the voxel hash grids.
+Since the hash tables take up the majority of the parameters, the model size is reduced from $21M$ to $7.5M$ as in Table 5.
+We use $2^{P-2}$-length hash tables for the Plane representation in Table 5 to balance the quality and the model size.
+The structural configurations $(L,F)$ of the plane representation in Table 5 are set as $(6,4)$ for better results.
 3.3 Results and comparisons
@@
 […]
-offline DyNeRF [28] 300 29.58 0.020 0.083 7 days 28MB >100GB
-NeRF-T [28] 300 28.45 0.023 0.10 - - >100GB
-Interp-NN [34] 300 29.88 - 0.096 2 days 20MB >100GB
-HexPlane [7] 300 31.57 0.016 0.089 96 min 200MB 62GB
-K-Planes [17] 300 31.63 - - 110 min 200MB >100GB
-MixVoxels-T [45] 300 30.71 0.024 0.162 15 min 500MB >100GB
-4DGS [54] 300 32.01 - 0.055 6 hours 4000MB <10GB
-HyperReel [1] 50 31.1 - 0.096 9 hours 360MB 18GB
-NeRFplayer [43] 300 30.29 - 0.152 5.5 hours - >100GB
-online StreamRF [27] 1 28.26 0.029 0.226 75 min 5000MB <10GB
-HexPlane-PR 10 24.03 0.081 0.244 100 min 318MB 14GB
-HexPlane-GR 10 25.17 0.050 0.272 125 min 318MB 14GB
-INV [48] 1 28.72 - 0.121 40 hours 336MB <10GB
-CD-NGP (Ours) 10 30.23 0.029 0.198 75 min 113MB 14GB
-4.3.1 Standard-length sequences.
+DyNeRF [28] $\checkmark$ 300 29.58 0.020 0.083 7 days 28MB >100GB
+NeRF-T [28] $\checkmark$ 300 28.45 0.023 0.10 - - >100GB
+Interp-NN [34] $\checkmark$ 300 29.88 - 0.096 2 days 20MB >100GB
+HexPlane [7] $\checkmark$ 300 31.57 0.016 0.089 96 min 200MB 62GB
+K-Planes [17] $\checkmark$ 300 31.63 - - 110 min 200MB >100GB
+MixVoxels-T [45] $\checkmark$ 300 30.71 0.024 0.162 15 min 500MB >100GB
+4DGS [55] $\checkmark$ 300 32.01 - 0.055 6 hours 4000MB <10GB
+HyperReel [1] $\checkmark$ 50 31.1 - 0.096 9 hours 360MB 18GB
+NeRFplayer [43] $\checkmark$ 300 30.29 - 0.152 5.5 hours - >100GB
+StreamRF [27] $\checkmark$ 1 28.26 0.029 0.226 75 min 5000MB <10GB
+HexPlane-PR $\checkmark$ 10 24.03 0.081 0.244 100 min 318MB 14GB
+HexPlane-GR $\checkmark$ 10 25.17 0.050 0.272 125 min 318MB 14GB
+INV [48] $\checkmark$ 1 28.72 - 0.121 40 hours 336MB <10GB
+CD-NGP (Ours) $\checkmark$ 10 30.23 0.029 0.198 75 min 113MB 14GB
+3.3.1 DyNeRF dataset.
 Table 1 shows the quantitative comparisons between CD-NGP and other state-of-the-art methods on DyNeRF dataset.
 We use the officially released metrics.
-[28, 34] are trained on 8 V100 GPUs or equally 4 A100 GPUs.
-[28, 34, 43] have no official implementations available.
+The methods in [28, 34] are trained on 8 V100 GPUs or equally 4 A100 GPUs.
+The methods in [28, 34, 43] have no official implementations available.
 We are unable to obtain the missing quality metrics in [17] because of their huge memory occupation in the default resolution.
-[7] uses resolution $(1024,768)$ in their implementation, so the memory occupation is much smaller than other offline methods.
+HexPlane [7] uses resolution $(1024,768)$ in their implementation, so the memory occupation is much smaller than other offline methods.
 We use the released output of [27] for DSSIM and LPIPS metrics.
-[27, 48] are methods trained per-frame.
-[43] is trained on the whole video sequence but streamable upon inference.
+StreamRF [27] and INV [48] are trained per-frame.
+NeRFplayer [43] is trained offline but streamable upon inference.
 4DGS [55] caches the frames onto disk for training to alleviate memory occupation with a trade-off on training speed.
@@
 We compare our method with both online and offline baselines.
-As demonstrated in Tab.
-1, our CD-NGP is 32×32$\times$32 × faster than the MLP-based continual representation [48], $\times4$ smaller than the voxel-based continual representation [27] and significantly outperforms the two additional baselines in the continual learning setting.
+As demonstrated in Table 1, our CD-NGP is 32×32$\times$32 × faster than the MLP-based continual representation INV [48], $\times4$ smaller than the voxel-based continual representation StreamRF [27] and significantly outperforms the two additional baselines in the continual learning setting.
 The undesirable results from HexPlane-PR and HexPlane-GR show the limitations of the regularization-based and replay-based methods in dynamic scenes: unlike the static scenes, the objects are not stationary across time and the models with limited complexity are not capable of reconstructing unlimited transient effects.
-Benefiting from the parameter isolation like previous methods [27, 48], our method achieves fast convergence and compact model size simultaneously: it leverages the hash tables to ensure fast convergence and reuses the features from differently-configured spatial hash tables to ensure high scalability.
+Benefiting from the parameter isolation like StreamRF [27] and INV [48], our method achieves fast convergence and compact model size simultaneously.
+It leverages the hash tables to ensure fast convergence and reuses the features from differently-configured spatial hash tables to ensure high scalability.
 Compared with prevailing offline methods [28, 34, 7, 17, 45, 55], our method only requires caching $T_{chunk}=10$ frames into buffer by continual learning and thus reduces memory occupation by $85\%$.
 Our method also achieves $40\%$ less model size ($0.4$MB/frame) by leveraging the feature fusion of the base and the auxiliary branches with comparable quality.
+Figure 4: The results of the proposed CD-NGP on DyNeRF dataset.
 Figure 3 provides qualitative comparisons.
@@
 4.
-As illustrated in Fig.
-6, our method reaches a new balance between rendering quality, model size, and training time while remaining memory-friendly, in which we use metrics on the flame salmon scene for fair comparison.
-The perceptual consistency is measured by $1-$ LPIPS, and the speed is the reciprocal of the training time.
-The compactness and RAM-Efficiency are mapped by $-log$ function from the model size and memory occupation respectively.
-Figure 4: The results of the proposed CD-NGP on DyNeRF dataset.
-4.3.2 Long sequences.
-We also provide quantitative comparisons on our proposed long multi-view video dataset in Tab.
-3 and Fig.
+(a) GT
+(b) HexPlane
+(c) MixVoxels
+(d) INV
+(e) StreamRF
+(f) Ours
+Figure 5: Comparison of reconstruction quality on the trimming scene in the Meetroom dataset.
+Our method produces comparable quality to both offline and online state-of-the-art methods but with much less memory and storage cost.
+3.3.2 Meetroom dataset.
+Table 2 and Fig.
+5 shows the quantitative and qualitative comparisons on the meetroom dataset.
+Fig.
+6 is the PSNR curve of the compared methods on the trimming scene.
+As shown in Table 2, HexPlane-GR and HexPlane-PR fail and produce unsatisfactory results on the meetroom dataset, we thus exclude them from Fig.
+6.
+Our method outperforms the implicit MLP-based online method [48] in terms of quality, speed, and model size significantly.
+Moreover, our method consumes only a fraction of storage than the explicit voxel-based online baseline [27] while maintaining comparable reconstruction quality as compared in Fig.
+5.
+When compared with offline baselines, our method produces comparable quality to MixVoxels [45] and outperforms HexPlane [7] in terms of quality and speed, and consumes much less memory as in Fig.
+5.
+We use the released configuration of [27] for the meetroom dataset.
+The full model of [27] results in much worse metrics than those reported in their paper, so we only report the metrics of their pilot model.
+Figure 6: PSNR curves of the compared methods on the trimming scene in the Meetroom dataset. Solid and dotted lines denote offline and online methods respectively.
+Table 2: Results on the Meetroom dataset.
+[…]
+Mixvoxels $\checkmark$ 26.96 0.027 0.1034 15min 500MB
+HexPlane $\checkmark$ 24.71 0.043 0.2206 100min 200MB
+HexPlane-GR $\checkmark$ 19.05 0.099 0.3999 100min 318MB
+HexPlane-PR $\checkmark$ 18.63 0.112 0.3589 100min 318MB
+StreamRF $\checkmark$ 26.48 0.029 0.1699 75min 1710MB
+INV $\checkmark$ 23.87 0.041 0.1195 40hours 336MB
+Ours $\checkmark$ 26.01 0.039 0.1906 77min 113MB
+3.3.3 Long sequence dataset and scalability analysis.
+We also provide quantitative comparisons on our proposed long multi-view video dataset in Table 3 and Fig.
 7 to further show the advantage in scalability of our method.
@@
 We set the hash table size as $2^{12}$ in the auxiliary branches and use $T_{chunk}=5$ in our model.
-The offline baseline [7] produces unstable results on one of the 5 scenes in our dataset, so we only use the other 4 scenes to obtain the metrics of [7].
-As shown in Tab.
-3, our method reduces memory and storage consumption by at least 4×4$\times$4 × with comparable quality to the offline baseline.
-Compared with the results in Tab.
-1, the performance gap in storage consumption is even larger in long video sequences, because the auxiliary branches consume the majority of the parameters and the complexity in storage is reduced to $\mathcal{O}(N_{chunk}\cdot 2^{12}LF)$.
+As shown in Table 3, our method reduces the memory and storage consumption by at least 4×4$\times$4 × with comparable quality to HexPlane [7].
+Compared with the results in Table 1, the performance gap in storage consumption is even larger in long video sequences.
+This is because the whole model shares one base branch and the following auxiliary branches occupy the majority of the parameters.
+We thus set the hash table size in the auxiliary branches to $2^{12}LF$ and leverage the re-use of the features from the base branch to reduce the model size to $\mathcal{O}(N_{chunk}\cdot 2^{12}LF)$.
+.
 The PSNR curves in Fig.
 7 further demonstrate the stability of our method.
-The offline baseline produces the best quality.
-The compared online methods suffer from catastrophic forgetting: the PSNRs at the beginning of the chunks are much lower than at the end.
-Our method resists the problem by parameter isolation and ensures the quality of the whole sequence.
-We also provide comparisons between rendered results in Fig.
+The compared online methods suffer from catastrophic forgetting: the PSNRs at the beginning of the chunks are much lower than that in the end.
+The problem is properly eliminated by our method through parameter isolation, and our method continuously produce high-quality results for the whole sequence.
+We also provide comparisons between the rendered results in Fig.
 8.
 The regularization method in Fig.
-7(a) reconstructs the TV screen in the background erroneously.
+8(a) reconstructs the TV screen in the background erroneously.
 The replay method in Fig.
-7(b) fails to recover the details.
+8(b) fails to recover the details.
 Ours in Fig.
-7(d) produces comparable quality to the offline baseline in Fig.
-7(c).
-Figure 5: Comparison between different performance indexes of our method and recent representative baselines.
-Figure 6: PSNR curves of the seminar scene. The solid line and dotted line denotes offline and online methods respectively.
+8(f) produces comparable quality to the offline baseline in Fig.
+8(c).
+Figure 7: PSNR curves of the seminar scene in our long video dataset. The solid lines and dotted lines denote offline and online methods respectively.
 3.4 Ablation studies
@@
 (c) HexPlane
-(d) CD-NGP (ours)
+(d) INV
+(e) StreamRF
+(f) CD-NGP (ours)
 Figure 8: Comparison of reconstruction quality on the 19th frame of the seminar scene in our long video dataset.
@@
 Benefiting from the multi-branch hash table design, our model is highly efficient and configurable.
-We change the parameter numbers of the initial branch and the auxiliary branches and obtain the results on the DyNeRF dataset in Tab.
-3.
+We change the parameter numbers of the initial branch and the auxiliary branches and obtain the results on the DyNeRF dataset in Table 4.
 $(P_{1},P_{2})$ are the log2 values of the hash table lengths of the base branch and the other branches.
-Please note the whole model sizes include the sizes of the occupancy grid and are thus larger than those deduced from parameter numbers.
-Larger hash tables alleviate hash collisions and lead to better results.
-However, the default configuration $(19,14)$ in CD-NGP produces better results than the naive $(16,16)$ configuration even if the storage consumption is $40\%$ lower.
+Since the whole model contains hash encoders, MLP regressors, and the occupancy grids.
+The sizes of the models are thus larger than the simple combinations of hash table and MLP regressors.
+We use the default configuration $(19,14)$ in the following experiments.
+Although larger hash tables alleviate hash collisions, the configuration $(19,14)$ in CD-NGP produces better results than the naive $(16,16)$ configuration even if the storage consumption is $40\%$ lower.
 Moreover, the $(19,12)$ configuration of CD-NGP significantly outperforms the naive $(14,14)$ configuration at a similar storage cost.
 To further demonstrate the scalability of our method, we provide the PSNR curves of different hash table designs on the sear steak scene in Fig.
-8.
+9.
 Equipped with a larger hash table in the base branch, the asymmetric setting $(19,14)$ and $(19,12)$ outperforms others in the beginning and produces comparable metrics to the naive $(16,16)$, $(14,14)$ configuration respectively in the subsequent chunks.
@@
 Metrics are measured on the first 1200 frames.
-Compared baselines are trained with a chunk size of 300 frames.
+The compared HexPlane baselines are trained with a chunk size of 300 frames.
 […]
-HexPlane-PR 20.90 0.286 1272MB 12GB
-HexPlane-GR 24.82 0.282 1272MB 12GB
-CD-NGP (ours) 26.50 0.201 201MB 12GB
-HexPlane [7] 28.32 0.117 800MB 74GB
+HexPlane-PR 20.90 0.286 1272MB 8 hours 12GB
+HexPlane-GR 24.82 0.282 1272MB 8 hours 12GB
+HexPlane 28.32 0.117 800MB 8 hours 74GB
+StreamRF 24.99 0.266 13GB 4 hours 12GB
+INV 27.07 0.086 1344MB 160 hours 12GB
+Ours 27.15 0.190 201MB 10 hours 12GB
 Table 4: Ablation study on hash table sizes. denotes the default setting.
@@
 $(19,12)$ 14M 30.09 82MB
-4.4.2 Spatial representations.
 Table 5: Ablation study on spatial representations. Different representations lead to subtle changes in quality metrics.
@@
 CD-NGP 21M 30.23 0.198 75min 113MB
+3D+4D 23M 29.34 0.243 82min 122MB
 Voxel-C 21M 30.52 0.174 80min 114MB
@@
 MERF-C 7.5M 30.22 0.174 115min 57MB
-Table 5: Ablation study on regularization weight $\lambda_{r}$. $10^{-3}$ is the default value.
-[…]
-$10^{-5}$ 30.20 0.0269 0.198
-$10^{-4}$ 30.18 0.0268 0.196
-$10^{-3}$ 30.23 0.0269 0.198
-$10^{-2}$ 30.09 0.0279 0.197
-1 29.66 0.0313 0.205
+3.4.2 Spatial representations and scalability analysis.
 The proposed hash-based pipeline is not only practical with voxel hash grids but also compatible with other hash-based spatial backbones.
-We provide quantitative results on plane representation like [17, 21] and hybrid representation (MERF) [38] on the DyNeRF dataset in Tab.
-4.
+To prove the flexibility and scalability of our method, we provide ablation studies on different spatial representations.
+We adopt the Eq.
+5 to the MERF representation [38], defined as:
+$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}\leavevmode\nobreak\ P_{m}(x,y,z)+Q_{mxy}(x,y)+Q_{myz}(y,z)+Q_{mzx}(z,x)$ (12)
+where $Q_{mij}(i,j)$ denotes the bi-linearly interpolated value from the 4 nearby vertices in the orthogonally projected planes w.r.t.
+axes $(i,j)$ at the $m$-th resolution.
+Removing the $P_{m}(x,y,z)$ term leads to a pure plane representation like [17, 21].
+We provide quantitative results on plane representation like [17, 21] and hybrid representation (MERF) [38] on the DyNeRF dataset in Table 5.
 Since the element-wise product for plane representation in [17] produces unsatisfactory results due to the hash collision, we remove the $P_{m}(x,y,z)$ in Eq.
 12 and fuse the features $Q_{mij}(i,j)$ by concatenation in the implementation.
+The 3D+4D representation similar to [34] is yielded by concatenating the spatial $(x,y,z)$ features and spatial-temporal $(x,y,z,t)$ features encoded by a 4D extension of Eq.
+5.
+By comparing our CD-NGP with the naive 3D+4D representation, we show the advantage of our asymmetric encoding of spatial-temporal $(x,y,z,t)$ features explained in Eqs.
+3 and 6: representing the features of spatial coordinates $(x,y,z)$ and temporal coordinate $(t)$ separately avoids feature collisions and lead to better results.
 Table 5 also shows metrics of different feature fusion methods.
 We replace the element-wise sum in Eqs.
-7 and 6 with vector concatenation and denote them as Voxel-C and MERF-C in Tab.
-4.
+6 and 12 with vector concatenation and denote them as Voxel-C and MERF-C in Table 5.
 The hybrid representations (MERF and MERF-C) achieve comparable model sizes ($\sim$60MB) to methods for static scenarios [31, 21] without significant quality loss, reducing the average bandwidth cost to $0.2$MB/frame.
 Benefiting from larger feature dimensions, the vector concatenation leads to slightly better quality (especially when measured in LPIPS) by more training time, and the reasonable trade-off also shows the robustness of our method.
-Compared with the hash table designs in Tab.
-3, the Voxel-C representation in Tab.
-4 almost performs as well as the bulky $(18,18)$ configuration in Tab.
-3 and reduces the model size by $80\%$.
-We also provide ablation studies on the regularization weights of auxiliary spatial encodings.
-We apply regularizations on $\Psi_{k}(x)$ with different $\lambda_{r}$ and show the results in Tab.
-5.
-The regularization is of slight improvement, and our method is robust within a wide range of $\lambda_{r}$.
-Please refer to the supplementary material for more ablation studies.
-5 Conclusion and Limitation
-We propose CD-NGP, a continual learning framework for multi-view dynamic scenes based on parameter isolation.
-Equipped with the multi-branch hash encoding structure, it shows high efficiency and scalability and is capable of reconstructing long video sequences with low host memory occupation.
-Compared with offline methods, it consumes $40\%$ less bandwidth and $85\%$ less memory than recent explicit baselines and achieves comparable quality while maintaining the speed advantage over the pure-MLP-based methods.
-Compared with online baselines, it significantly outperforms replay and regularization-based continual learning methods and reaches a new balance among reconstruction quality, training speed, bandwidth occupation, and memory cost.
-However, like most of the NeRF-based methods, it still requires querying the field function and is not appropriate for achieving real-time rendering.
-Further efforts could be made to apply baking methods or Gaussian point clouds to continual learning to achieve real-time rendering in long video sequences for wide applications.
-Appendix 0.A Detailed Results on DyNeRF Dataset
-We list the per-scene results of the proposed continual learning method with different spatial representations on the DyNeRF dataset in Tab. 6.
+Compared with the hash table designs in Table 4, the Voxel-C representation in Table 5 almost performs as well as the bulky $(18,18)$ configuration in Table 4 and reduces the model size by $80\%$.
+We list the per-scene results of the proposed continual learning method with different spatial representations on the DyNeRF dataset in Table 6.
 Table 6: Detailed results of the proposed CD-NGP and other hash-based methods on the DyNeRF dataset.
@@
 MERF-C 31.04 0.02450 0.1632 30.216 0.027814 0.17360
-Appendix 0.B More Implementation Details
-0.B.1 Parameter numbers and streaming efficiency.
-We set the dimensions of the latent feature $\mathbf{H}$ in Eq.
-3 as 48.
-We use MLPs with two hidden layers and the dimensions of the hidden layers in Eqs.
-3 and 4 are set as 128, and 64 respectively.
-We use MLPs with identical structures for different branches.
-We set the resolutions of the hash encoders of the proposed CD-NGP and MERF representations within $(16,2048)$ times the scale of the scene, and we use $(64,2048)$ for the plane representation.
-Following MERF [38] for static scenes, we use a smaller hash table for the MERF and the MERF-C than pure voxel hash grid [31] representation in Tab.
-4 to ensure the compactness and scalability of our method.
-Denoting the lengths of hash tables along pure voxels in a branch of CD-NGP as $2^{P}$, the lengths of hash tables in the MERF and the MERF-C are set as $2^{P-3}$ and $2^{P-4}$ for voxel and plane encoders respectively.
-The number of the parameters of the MERF representation is thus reduced to $1-2^{-3}-3\times 2^{-4}=31.25\%$ of the voxel hash grids.
-Since the hash tables take up the majority of the parameters, the model size is reduced from $21M$ to $7.5M$ as in Tab.
-4.
-We use $2^{P-2}$-length hash tables for the Plane representation in Tab.
-4 to balance the quality and the model size.
-The structural configurations $(L,F)$ of the plane representation in Tab.
-4 are set as $(6,4)$ for better results.
-The detailed parameter numbers and bandwidth consumption are thus listed in Tab.
-7.
+The detailed parameter numbers and bandwidth consumption are thus listed in Table 7.
 3D Enc.
@@
 Since the base encoder and the density grid could be shared among different branches and transmitted only once in the beginning, the minimum bandwidth consumption is reduced to $0.065\sim 0.16$ MB per frame, which is 100×100$\times$100 × less than the voxel-based StreamRF [27] and 7×7$\times$7 × less than the pure MLP representation INV [48].
-Table 7: The parameter numbers of each module of the 30-branch representations in Tab.
-4.
+Table 7: The parameter numbers of each module of the 30-branch representations in Table 5.
 $\mathbf{B_{min}}$ and $\mathbf{B_{avg}}$ are minimum and average bandwidth costs in MB/frame.
@@
 MERF-C $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.052M$ $0.17M$ $7.5M$ 0.065 0.19
-0.B.2 Continual learning pipeline.
-We further illustrate the whole process of our method in Algorithm 1, where $\eta_{init}$ and $\eta_{aux}$ denote the numbers of iterations of the initial branch and the auxiliary branches, respectively.
-$\Psi_{i},\Gamma_{i},\theta_{1}^{(i)},\theta_{2}^{(i)}$ denote the model parameters in the branches.
-The whole model is trained chunk by chunk.
-The implementation will be available upon publication.
-Algorithm 1 Continual learning pipeline
-0:  $N_{chunk},T_{chunk},\eta_{init},\eta_{aux},\Psi_{i},\Gamma_{i},\theta_{1}^{(i)},\theta_{2}^{(i)},i\in\mathbb{N}$
-0:  The whole representation for dynamic scenes.
-[…]
-2:  while $N_{chunk}$ do
-[…]
-[…]
-5:     if $k~{}\mathbf{mod}~{}T_{episode}~{}=0$  then
-[…]
-7:     else
-[…]
-[…]
-10:     end if
-11:     Train the $k^{th}$ model branch $(\Psi_{k},\Gamma_{k},\theta_{1}^{(k)},\theta_{2}^{(k)})$ for $\eta$ steps.
-[…]
-13:  end while
-Appendix 0.C More Details about the Long Video Dataset
-We collect the dataset for evaluating the scalability of our method and further research.
-We have checked the videos to make sure they do not contain controversial or offensive content before using them for research purposes.
-The faces in the videos may contain identifiable information, but the information will neither be used for any other purposes nor affect the anonymity of the submission.
-All videos are collected with the explicit consent of all the participants.
-The view in the upper center is selected for validation.
-We synchronize the videos by environmental sound, and we put chessboard images in the scene to help calibrate the cameras and estimate the camera poses.
-Appendix 0.D More Ablation Studies
-0.D.1 Temporal representations.
+3.4.3 Temporal representations.
 Our continual learning pipeline is also compatible with a variety of representations for the time axis.
 Currently, there are 2 other prevailing representations for the time axis alone: pure frequency encoding [36] and MLP with frequency-encoding [16].
-We replace the temporal hash encoding in our method with them on DyNeRF dataset and show the results in Tab.
-9.
+We replace the temporal hash encoding in our method with the ones mentioned above on DyNeRF dataset and show the results in Table 8.
 Our time latent code slightly outperforms others on DyNeRF dataset.
@@
 Time-Freq 30.19 0.0272 0.203
-Table 9: Ablation study on field composition methods.
+Table 9: Ablation study on the field composition methods.
 […]
@@
 Static-dynamic 28.27 0.0568 0.291
-0.D.2 Field composition methods.
-We also conduct ablation studies on the field composition methods and show the results in Tabs. 9 and 9. The field composition process is defined as:
+3.4.4 Field composition methods.
+We also conduct ablation studies on the field composition methods and show the results in Tables 9 and 10. The field composition process is defined as:
 $(\sigma,\mathbf{c})=(\sigma_{i}+\sigma_{a},\mathbf{c}_{i}+\mathbf{c}_{a}).$ (13)
-Full composition in Tab.
-9 denotes the naive composition of 2 fields from the base branch and the current branch respectively, i.e.
+Full composition in Table 9 denotes the naive composition of 2 fields from the base branch and the current branch respectively, i.e.
 $\sigma_{i}$ and $\mathbf{c}_{i}$ are the field from the base branch and $\sigma_{a}$ and $\mathbf{c}_{a}$ in Eq.
@@
 3.
-Static-dynamic in Tab.
-9 denotes a naive composition of the field from a static branch by the base encoder and a dynamic branch by the auxiliary encoders similar to [45], i.e.
+Static-dynamic in Table 9 denotes a naive composition of the field from a static branch by the base encoder and a dynamic branch by the auxiliary encoders similar to [45], i.e.
 $\sigma_{i}$ and $\mathbf{c}_{i}$ in Eq.
@@
 Furthermore, as compared in Fig.
-9, the compared composition methods in Figs.
-9(c) and 9(d) suffer from salt and pepper noise because of the explicit composition, while our CD-NGP circumvents the problem by fusing features in the latent space and produces better results.
+10, the compared composition methods in Figs.
+10(c) and 10(d) suffer from salt and pepper noise because of the explicit composition, while our CD-NGP circumvents the problem by fusing features in the latent space and produces better results.
 (a) Ground truth
@@
 Figure 10: Comparison of different field composition methods.
-0.D.3 Chunk size.
-We also conduct ablation studies on the chunk sizes and show the results in Tabs.
-11 and 11.
+3.4.5 Regularization weights.
+We also provide ablation studies on the regularization weights of auxiliary spatial encodings.
+We apply regularizations on $\Psi_{k}(x)$ with different $\lambda_{r}$ and show the results in Table 10.
+The regularization is of slight improvement, and our method is robust within a wide range of $\lambda_{r}$.
+Table 10: Ablation study on regularization weight $\lambda_{r}$.
+[…]
+$10^{-5}$ 30.20 0.0269 0.198
+$10^{-4}$ 30.18 0.0268 0.196
+$10^{-3}$ 30.23 0.0269 0.198
+$10^{-2}$ 30.09 0.0279 0.197
+1 29.66 0.0313 0.205
+3.4.6 Chunk size.
+We also conduct ablation studies on the chunk sizes and show the results in Tables 11 and 12.
 We only report the results on the cook spinach scene in the DyNeRF dataset for brevity.
 Our method is robust within a wide range of chunk sizes.
-As in Tab.
-11, the smaller chunk size leads to better results when controlling the iteration numbers of each chunk.
-When the total iterations of the whole 300-frame sequence are held constant, the larger chunk size leads to better results as in Tab.
-11.
-We do not cherry-pick the best chunk size and select $T_{chunk}=10$ as the default configuration in our method.
+As in Table 11, the smaller chunk size leads to better results when controlling the iteration numbers of each chunk.
+When the total iterations of the whole 300-frame sequence are held constant, the larger chunk size leads to better results as in Table 12.
+We select $T_{chunk}=10$ as the default configuration.
 The selection of the chunk size is a trade-off between the quality, robustness, training time, and host memory occupation and should depend on the data.
@@
 5 30.39 0.02602 0.2075 162MB
+4 Conclusion and Limitation
+We propose CD-NGP, a continual learning framework for multi-view dynamic scenes based on parameter isolation.
+Equipped with the multi-branch hash encoding structure, it shows high efficiency and scalability and is capable of reconstructing long video sequences with low host memory occupation.
+Compared with offline methods, it consumes $40\%$ less bandwidth and $85\%$ less memory than recent explicit baselines and achieves comparable quality while maintaining the speed advantage over the pure-MLP-based methods.
+Compared with online baselines, it significantly outperforms replay and regularization-based continual learning methods and reaches a new balance among reconstruction quality, training speed, bandwidth occupation, and memory cost.
+However, like most of the NeRF-based methods, it still requires querying the field function and is not appropriate for achieving real-time rendering.
+Further efforts could be made to apply baking methods or Gaussian point clouds to continual learning to achieve real-time rendering in long video sequences for wide applications.
+Appendix A Detailed results on the DyNeRF dataset

### v4

@@
 Abstract
-We present CD-NGP, a fast and scalable representation for 3D reconstruction and novel view synthesis in dynamic scenes.
-Inspired by continual learning, our method first segments input videos into multiple chunks, followed by training the model chunk by chunk, and finally, fuses features of the first branch and subsequent branches.
-Experiments on the prevailing DyNeRF dataset demonstrate that our proposed novel representation reaches a great balance between memory consumption, model size, training speed, and rendering quality.
-Specifically, our method consumes $85\%$ less training memory ($<14$GB) than offline methods and requires significantly lower streaming bandwidth ($<0.4$MB/frame) than other online alternatives.
-Furthermore, we conduct experiments on a self-collected extra-long video sequences dataset, which demonstrates the superior scalability of our method over previous state-of-the-art approaches.
-A free copy of this paper and all supplemental materials are available at https://arxiv.org/abs/2409.05166.
+Current methodologies for novel view synthesis (NVS) in dynamic scenes encounter significant challenges in harmonizing memory consumption, model complexity, training efficiency, and rendering fidelity.
+Existing offline techniques, while delivering high-quality results, are often characterized by substantial memory demands and limited scalability.
+In contrast, online methods grapple with the challenge of balancing rapid convergence with model compactness.
+To address these issues, we propose continual dynamic neural graphics primitives (CD-NGP).
+Our approach synergizes features from both temporal and spatial hash encodings to achieve high rendering quality, employs parameter reuse to enhance scalability, and leverages a continual learning framework to mitigate memory overhead.
+Furthermore, we introduce a novel dataset comprising multi-view, exceptionally long video sequences with substantial rigid and non-rigid motion, thereby substantiating the scalability of our method.
+We evaluate the quality, speed and scalability our method on both established public datasets and our exceptionally long video dataset.
+Notably, our method achieves an $85\%$ reduction in training memory consumption (less than 14GB) compared to offline techniques and significantly lowers streaming bandwidth requirements (less than 0.4MB/frame) relative to other online alternatives.
+The experimental results on our long video sequences dataset show the superior scalability and quality compared to existing state-of-the-art approaches.
 keywords:
@@
 The demonstration of our continual learning pipeline for dynamic scenes.
-Please note the dimensions of representations for different branches remain the same while the complexities of hash tables are flexible and could be configured to ensure scalability.
-The spatial hash features could be encoded along voxels [31], planes [17] or both [38].
+Our method segments input videos into multiple chunks, train the model chunk by chunk and leverage the re-use of parameters from the first branch to ensure high scalability.
+The spatial hash features could be encoded along voxels, orthogonally projected planes or both.
 Introduction
@@
 In this paper, we propose continual dynamic neural graphics primitives (CD-NGP), a fast and scalable representation for reconstructing dynamic scenes, which enables dynamic reconstruction in long videos.
-As illustrated in CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes, we evenly split the dynamic videos into multiple chunks based on the timestamp and conduct reconstruction with different model branches, where each branch includes MLP regressors, spatial encoding for the spatial coordinates $(x,y,z)$, and temporal encoding for the time stamp $t$.
-We encode the different chunks by hash tables with different complexities to ensure the model scalability and then continually train the model chunk by chunk to alleviate memory occupation.
-For each individual chunk, we locate the corresponding model branch, sample 3D points $(x,y,z)$ according to frame pixels and cached occupancy grids [31], and represent spatial features by spatial hash grids.
-The spatial features of the first branch and the current branch are fused by element-wise sum as default or concatenation if applicable.
-The temporal features are encoded by the hash table for time stamp $t$ in the current branch only.
-Spatial and temporal features are concatenated along with viewing direction $d$ and input to MLP regressors for rendering and optimization.
+The proposed methodology is illustrated in CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes: we partition dynamic videos into multiple segments based on timestamps and perform reconstruction using distinct model branches.
+Each branch includes MLP regressors, spatial hash encoding for spatial coordinates $(x,y,z)$, and temporal hash encoding for the timestamp $t$.
+For the first segment, we encode spatial features using a large hash table, and for subsequent segments, we use a smaller hash table.
+By reusing parameters from the initial branch, we enhance scalability and accelerate convergence.
+For each segment, we identify the corresponding model branch, sample 3D points $(x,y,z)$ according to frame pixels and cached occupancy grids [31], and represent spatial features using spatial hash grids.
+The spatial features from the initial branch and the current branch are fused through element-wise summation by default or concatenation if applicable.
+Temporal features are encoded by the hash table for the timestamp $t$ within the current branch only.
+These spatial and temporal features, along with the viewing direction $d$, are concatenated and input into MLP regressors for rendering and optimization.
 We validate the effectiveness of our method through comprehensive experimental results on both the widely used public datasets and the dataset proposed by us.
@@
 3D coordinates $(x,y,z)$ at the $l$-th resolution, $L$ is the number of the resolutions, and $F$ is the dimensions of features $P_{m}(x,y,z)$, which are similar to those in Instant-NGP [31].
+If necessary, Eq.
+5 could be adopted to the hybrid (MERF) representation [38], defined as:
+$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}\leavevmode\nobreak\ P_{m}(x,y,z)+Q_{mxy}(x,y)+Q_{myz}(y,z)+Q_{mzx}(z,x)$ (6)
+where $Q_{mij}(i,j)$ denotes the bi-linearly interpolated value from the 4 nearby vertices in the orthogonally projected planes w.r.t.
+axes $(i,j)$ at the $m$-th resolution.
+Removing the $P_{m}(x,y,z)$ term leads to a pure plane representation like [17, 21].
 2.3 Continual dynamic neural graphics primitives (CD-NGP)
@@
 4 with the color MLP $f_{\theta_{2k}}$.
+For simplicity, we denote the models adopting pure voxel encoders in Eq.
+5 as CD-NGP, the models adopting the hybrid encoders in Eq.
+6 as CD-MERF.
 The scalability of our method derives from the redundancy of dynamic scenes and the flexibility of hash table sizes.
@@
 3.1 Datasets
-DyNeRF dataset [28] is the most prevailing dataset for dynamic view synthesis that provides high-quality multi-camera videos, on which most of the current state-of-the-art methods are evaluated.
-DyNeRF dataset contains 6 individual cooking scenes with varied illumination, food topological changes, and transient effects like flames.
-It contains five 300-frame videos and one 1200-frame video, with up to 21 synchronized camera views at the resolution of $2704\times 2028$ at 30 frames per second (FPS).
+The DyNeRF dataset [28] is the most prevalent benchmark for dynamic view synthesis, providing high-quality multi-camera videos on which current state-of-the-art methods are primarily evaluated.
+This dataset includes six unique cooking scenes, each featuring varying illumination, food topological changes, and transient effects such as flames.
+It consists of five videos, each containing 300 frames, and one extended video with 1200 frames, captured at a resolution of $2704\times 2028$ with up to 21 camera views at 30 frames per second (FPS).
+Following the approach of HexPlane [7], we excluded the unsynchronized scene coffee martini and conducted experiments on the remaining scenes, using the first 300 frames of the 1200-frame video flame salmon for evaluation.
 Meetroom dataset [27] is another dataset for dynamic view synthesis that contains 3 scenes captured by 13 synchronized cameras at $1280\times 720$ resolution.
 The lengths of the scenes are set as 300 frames and the refresh rates are set as 30 FPS.
-To further prove the viability of our method on long video sequences, we provide a new dataset containing 5 dynamic scenes at $3840\times 2160$ resolution, the lengths of which are at least 200 seconds.
-Similar to the DyNeRF dataset [28], scenes in our dataset are recorded by 20 GoPro cameras at 30 FPS.
-We synchronize the videos by environmental sound, and we put chessboard images in the scene to help calibrate the cameras and estimate the camera poses.
-Compared with those in the DyNeRF dataset, the dynamics are much harder, including large motions of human bodies and transient effects in the background.
-We collect the dataset for evaluating the scalability of our method and further research.
-We have checked the videos to make sure they do not contain controversial or offensive content before using them for research purposes.
-The faces in the videos may contain identifiable information, but the information will neither be used for any other purposes except for research nor be disclosed to any third party.
-All videos are collected with the explicit consent of all the participants.
-The view in the upper center is selected for validation.
+To further validate the effectiveness of our method on long video sequences, we introduce a new dataset comprising 5 dynamic scenes at a resolution of $3840\times 2160$.
+Similar to the DyNeRF dataset [28], these scenes are captured by 20 GoPro cameras operating at 30 FPS.
+We synchronize the multi-view videos using environmental sound and incorporate chessboard images within the scenes to facilitate camera calibration and pose estimation.
+The dynamics in our dataset are notably more challenging than those in the DyNeRF dataset, featuring significant non-rigid motions of human bodies and transient effects such as slide effects on the screens in the backgrounds.
+To comprehensively assess the scalability of our method, we provide 5 scenes with a duration exceeding 200 seconds (6000 frames), which is over 10 times longer than the previous DyNeRF dataset.
+This dataset was meticulously collected to evaluate the scalability of our method and to support further research.
+We have carefully reviewed the videos to ensure they do not contain any controversial or offensive content before using them for research purposes.
+While the faces in the videos may contain identifiable information, this information will be used solely for research and will not be disclosed to any third party.
+All videos were collected with the explicit consent of all participants.
+The view in the upper center is designated for validation.
 Figure 2: Example frames sampled from the proposed long multi-view video dataset.
@@
 Similar to the previous work [7, 28], we select the view in the center with index 00 for evaluation on both datasets.
-Following [7], we exclude the unsynchronized coffee martini scene and conduct experiments on the others.
 Following most of the state-of-the-art methods [28, 27, 17, 45, 43] we train and evaluate the model at resolution $1352\times 1014$ on the DyNeRF dataset.
@@
 We set the size of the hash table in the first branch as $2^{19}$ and the size of the hash table in the subsequent branches as $2^{14}$, and we set the structural parameters $(L,F)$ in the spatial hash encoders as $(12,2)$ if not otherwise specified.
+Following MERF [38] for static scenes, we use a smaller hash table for the hybrid representation in Eq.
+6 than pure voxel hash grid [31] representation in Table 3 to ensure the compactness and scalability of our method.
+Denoting the lengths of hash tables along pure voxels in a branch of CD-NGP as $2^{P}$, the lengths of hash tables in the CD-MERF are set as $2^{P-3}$ and $2^{P-4}$ for voxel and plane encoders respectively.
+The number of the parameters of the MERF representation is thus reduced to $1-2^{-3}-3\times 2^{-4}=31.25\%$ of the voxel hash grids.
+Since the hash tables take up the majority of the parameters, the model size is reduced from $21M$ to $7.5M$ as in Table 3.
 At the base branch, we initialize the models and train for 18k iterations, and for each of the subsequent branches we train for 3k iterations using the previous branch for initialization.
@@
 We set the resolutions of the hash encoders of the CD-NGP within $(16,2048)$ times the scale of the scene, and we use $(64,2048)$ for the plane representation.
-Following MERF [38] for static scenes, we use a smaller hash table for the MERF and the MERF-C than pure voxel hash grid [31] representation in Table 5 to ensure the compactness and scalability of our method.
-Denoting the lengths of hash tables along pure voxels in a branch of CD-NGP as $2^{P}$, the lengths of hash tables in the MERF and the MERF-C are set as $2^{P-3}$ and $2^{P-4}$ for voxel and plane encoders respectively.
-The number of the parameters of the MERF representation is thus reduced to $1-2^{-3}-3\times 2^{-4}=31.25\%$ of the voxel hash grids.
-Since the hash tables take up the majority of the parameters, the model size is reduced from $21M$ to $7.5M$ as in Table 5.
-We use $2^{P-2}$-length hash tables for the Plane representation in Table 5 to balance the quality and the model size.
-The structural configurations $(L,F)$ of the plane representation in Table 5 are set as $(6,4)$ for better results.
 3.3 Results and comparisons
@@
 - denotes the metrics unreported in the original publications.
-denotes the methods measured only on the most challenging flame salmon scene.
 Best, Second, and Third metrics are in highlight.
+The full model of [27] results in much worse metrics than those reported in their paper, so we only report the metrics of their pilot model.
 […]
@@
 NeRFplayer [43] $\checkmark$ 300 30.29 - 0.152 5.5 hours - >100GB
-StreamRF [27] $\checkmark$ 1 28.26 0.029 0.226 75 min 5000MB <10GB
+StreamRF [27] $\checkmark$ 1 28.85 0.042 0.253 75 min 5000MB <10GB
 HexPlane-PR $\checkmark$ 10 24.03 0.081 0.244 100 min 318MB 14GB
 HexPlane-GR $\checkmark$ 10 25.17 0.050 0.272 125 min 318MB 14GB
-INV [48] $\checkmark$ 1 28.72 - 0.121 40 hours 336MB <10GB
+INV [48] $\checkmark$ 1 29.64 0.023 0.078 40 hours 336MB <10GB
 CD-NGP (Ours) $\checkmark$ 10 30.23 0.029 0.198 75 min 113MB 14GB
@@
 HexPlane [7] uses resolution $(1024,768)$ in their implementation, so the memory occupation is much smaller than other offline methods.
-We use the released output of [27] for DSSIM and LPIPS metrics.
+We use the official implementation of StreamRF [27] and INV [48] and report the average metrics on the 5 synchronized scenes of the dataset.
 StreamRF [27] and INV [48] are trained per-frame.
@@
 5.
-We use the released configuration of [27] for the meetroom dataset.
+We use the released configuration of StreamRF [27] for the meetroom dataset.
 The full model of [27] results in much worse metrics than those reported in their paper, so we only report the metrics of their pilot model.
@@
 INV $\checkmark$ 23.87 0.041 0.1195 40hours 336MB
-Ours $\checkmark$ 26.01 0.039 0.1906 77min 113MB
+CD-NGP (Ours) $\checkmark$ 26.01 0.039 0.1906 77min 113MB
+3.3.3 Spatial representations.
+Table 3: Ablation study on spatial representations. Different representations lead to subtle changes in quality metrics.
+[…]
+CD-NGP 21M 30.23 0.198 75min 113MB
+3D+4D 23M 29.34 0.243 82min 122MB
+CD-NGP-C 21M 30.52 0.174 80min 114MB
+CD-Plane 14M 29.16 0.215 77min 83MB
+CD-MERF 7.5M 29.93 0.201 109min 57MB
+CD-MERF-C 7.5M 30.22 0.174 115min 57MB
+Our proposed continual learning framework is compatible with multiple spatial backbones.
+We provide quantitative results on hybrid representation (denoted as CD-MERF) and plane representation like [17, 21] (denoted as CD-Plane) on the DyNeRF dataset in Table 3.
+Since the element-wise product for plane representation in [17] produces unsatisfactory results due to the hash collision, we remove the $P_{m}(x,y,z)$ in Eq.
+6 and fuse the features $Q_{mij}(i,j)$ by concatenation in the implementation.
+We use $2^{P-2}$-length hash tables for the Plane representation in Table 3 to balance the quality and the model size.
+The structural configurations $(L,F)$ of the plane representation in Table 3 are set as $(6,4)$ for better results.
+The 3D+4D representation similar to [34] is yielded by concatenating the spatial $(x,y,z)$ features and spatial-temporal $(x,y,z,t)$ features encoded by a 4D extension of Eq.
+5.
+By comparing our CD-NGP with the naive 3D+4D representation, we show the advantage of our asymmetric encoding of spatial-temporal $(x,y,z,t)$ features explained in Eqs.
+3 and 7: representing the features of spatial coordinates $(x,y,z)$ and temporal coordinate $(t)$ separately avoids feature collisions and lead to better results.
+Table 3 also shows metrics of different feature fusion methods.
+We replace the element-wise sum in Eqs.
+7 and 6 with vector concatenation and denote them as CD-NGP-C and CD-MERF-C in Table 3.
+The hybrid representations (CD-MERF and CD-MERF-C) achieve comparable model sizes ($\sim$60MB) to methods for static scenarios [31, 21] without significant quality loss, reducing the average bandwidth cost to $0.2$MB/frame.
+Benefiting from larger feature dimensions, the vector concatenation leads to slightly better quality (especially when measured in LPIPS) by more training time, and the reasonable trade-off also shows the robustness of our method.
+Compared with the hash table designs in Table 7, the CD-NGP-C representation in Table 3 almost performs as well as the bulky $(18,18)$ configuration in Table 7 and reduces the model size by $80\%$.
+We list the per-scene results of the proposed continual learning method with different spatial representations on the DyNeRF dataset in Table 4.
+Table 4: Detailed results of the proposed CD-NGP and other hash-based methods on the DyNeRF dataset.
+Model Flame Salmon Cook Spinach
+[…]
+CD-NGP 26.37 0.04586 0.2886 30.76 0.02440 0.1793
+CD-NGP-C 26.65 0.04243 0.2680 31.03 0.02288 0.1593
+CD-Plane 25.48 0.06058 0.3173 29.41 0.03487 0.2010
+CD-MERF 26.15 0.04952 0.2921 30.35 0.02931 0.1858
+CD-MERF-C 26.56 0.04435 0.2546 30.69 0.02609 0.1583
+Model Flame Steak Sear Steak
+[…]
+CD-NGP 31.27 0.02132 0.1736 31.71 0.01917 0.1701
+CD-NGP-C 31.64 0.01823 0.1403 32.02 0.01698 0.1424
+CD-Plane 30.26 0.02949 0.1804 30.56 0.02723 0.1811
+CD-MERF 31.01 0.02508 0.1719 31.44 0.02181 0.1698
+CD-MERF-C 31.13 0.02424 0.1506 31.66 0.01989 0.1413
+Model Cut Roasted Beef Average
+[…]
+CD-NGP 31.03 0.02366 0.1801 30.228 0.026882 0.19834
+CD-NGP-C 31.26 0.02205 0.1578 30.520 0.024514 0.17356
+CD-Plane 30.11 0.03302 0.1951 29.164 0.037038 0.21498
+CD-MERF 30.71 0.02765 0.1856 29.932 0.030674 0.20104
+CD-MERF-C 31.04 0.02450 0.1632 30.216 0.027814 0.17360
+The detailed parameter numbers and bandwidth consumption are listed in Table 5.
+3D Enc.
+and 2D Enc.
+denote the spatial encodings along voxels and planes respectively.
+The tuples $(M_{1},M_{2})$ in the cells denote the parameter numbers of the base encoder and each of the auxiliary encoders.
+Following [27, 43, 1, 48] we also report the minimum per-frame bandwidth consumption $\mathbf{B_{min}}$, which only depends on the parameter numbers of the auxiliary branches and is thus significantly lower than the average bandwidth consumption $\mathbf{B_{avg}}$.
+Since the base encoder and the density grid could be shared among different branches and transmitted only once in the beginning, the minimum bandwidth consumption is reduced to $0.065\sim 0.16$ MB per frame, which is 100×100$\times$100 × less than the voxel-based StreamRF [27] and 7×7$\times$7 × less than the pure MLP representation INV [48].
+Table 5: The parameter numbers of each module of the 30-branch representations in Table 3.
+$\mathbf{B_{min}}$ and $\mathbf{B_{avg}}$ are minimum and average bandwidth costs in MB/frame.
+##\## denotes the number of parameters per branch (10 frames).
+Model ##\##3D Enc. ##\##2D Enc. ##\##MLP ##\##Online ##\##Total $\mathbf{B_{min}}$ $\mathbf{B_{avg}}$
+CD-NGP $(9.9M,0.37M)$ $0.042M$ $0.39M$ $21M$ 0.16 0.38
+CD-NGP-C $(9.9M,0.37M)$ $0.046M$ $0.40M$ $21M$ 0.16 0.38
+CD-Plane - $(1.9M,0.082M)$ $0.048M$ $0.28M$ $14M$ 0.11 0.28
+CD-MERF $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.042M$ $0.16M$ $7.5M$ 0.061 0.19
+CD-MERF-C $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.052M$ $0.17M$ $7.5M$ 0.065 0.19
 3.3.4 Long sequence dataset and scalability analysis.
 We also provide quantitative comparisons on our proposed long multi-view video dataset in Table 6 and Fig.
-7 to further show the advantage in scalability of our method.
+7 to further show the superior scalability of our method.
 We split the sequence into 300-frame chunks for the compared methods since the offline baseline [7] is designed for 300-frame sequences in the DyNeRF dataset.
@@
 8.
-The regularization method in Fig.
+The regularization method displayed in Fig.
 8(a) reconstructs the TV screen in the background erroneously.
-The replay method in Fig.
+The replay method displayed in Fig.
 8(b) fails to recover the details.
@@
 8(c).
-Figure 7: PSNR curves of the seminar scene in our long video dataset. The solid lines and dotted lines denote offline and online methods respectively.
+To further demonstrate the scalability of our method, we conduct experiments on the first 2400 frames of the seminar scene in our long video dataset, show the PSNR curves in Fig.
+10 and compare the model sizes in Fig.
+11.
+For simplicity of the evaluation, we set the chunk size as 10 for the CD-NGP and CD-MERF method.
+Our method produces high-quality results, and consumes only a fraction of storage space compared with the other online baselines.
+Figure 7: PSNR curves of the first 1200 frames of seminar scene in our long video dataset. The solid lines and dotted lines denote offline and online methods respectively.
 3.4 Ablation studies
@@
 Figure 9: PSNR curves of asymmetric (solid line, ours) and symmetric (dotted line) hash table designs on the sear steak scene.
+Figure 10: PSNR curves of the first 2400 frames of seminar scene in our long video dataset.
+Figure 11: Comparison of the model sizes of our method and other online baselines. Our method show superior scalability over the compared baselines.
 3.4.1 Lengths of hash tables.
@@
 The advantage comes from our asymmetric hash table designs: the base branch with a large hash table recovers more environment patterns, and the auxiliary branches could thus use fewer parameters to cope with recent changes only.
-Table 3: Comparisons between our CD-NGP and baselines on the long video sequences in our dataset.
-denotes the method trained offline.
-Metrics are measured on the first 1200 frames.
+Table 6: Quantitative comparisons between our CD-NGP and baselines on the long video sequences in our dataset.
+Metrics are measured on the first 1200 frames and averaged among the different scenes.
 The compared HexPlane baselines are trained with a chunk size of 300 frames.
 […]
-HexPlane-PR 20.90 0.286 1272MB 8 hours 12GB
-HexPlane-GR 24.82 0.282 1272MB 8 hours 12GB
-HexPlane 28.32 0.117 800MB 8 hours 74GB
-StreamRF 24.99 0.266 13GB 4 hours 12GB
-INV 27.07 0.086 1344MB 160 hours 12GB
-Ours 27.15 0.190 201MB 10 hours 12GB
+HexPlane $\checkmark$ 28.32 0.117 800MB 8 hours 74GB
+HexPlane-PR $\checkmark$ 20.90 0.286 1272MB 8 hours 12GB
+HexPlane-GR $\checkmark$ 24.82 0.282 1272MB 8 hours 12GB
+StreamRF $\checkmark$ 24.99 0.266 13GB 4 hours 12GB
+INV $\checkmark$ 27.07 0.086 1344MB 160 hours 12GB
+CD-NGP (Ours) $\checkmark$ 27.15 0.190 201MB 10 hours 12GB
 Table 7: Ablation study on hash table sizes. denotes the default setting.
@@
 $(19,12)$ 14M 30.09 82MB
-Table 5: Ablation study on spatial representations. Different representations lead to subtle changes in quality metrics.
-[…]
-CD-NGP 21M 30.23 0.198 75min 113MB
-3D+4D 23M 29.34 0.243 82min 122MB
-Voxel-C 21M 30.52 0.174 80min 114MB
-Plane 14M 29.16 0.215 77min 83MB
-MERF 7.5M 29.93 0.201 109min 57MB
-MERF-C 7.5M 30.22 0.174 115min 57MB
-3.4.2 Spatial representations and scalability analysis.
-The proposed hash-based pipeline is not only practical with voxel hash grids but also compatible with other hash-based spatial backbones.
-To prove the flexibility and scalability of our method, we provide ablation studies on different spatial representations.
-We adopt the Eq.
-5 to the MERF representation [38], defined as:
-$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}\leavevmode\nobreak\ P_{m}(x,y,z)+Q_{mxy}(x,y)+Q_{myz}(y,z)+Q_{mzx}(z,x)$ (12)
-where $Q_{mij}(i,j)$ denotes the bi-linearly interpolated value from the 4 nearby vertices in the orthogonally projected planes w.r.t.
-axes $(i,j)$ at the $m$-th resolution.
-Removing the $P_{m}(x,y,z)$ term leads to a pure plane representation like [17, 21].
-We provide quantitative results on plane representation like [17, 21] and hybrid representation (MERF) [38] on the DyNeRF dataset in Table 5.
-Since the element-wise product for plane representation in [17] produces unsatisfactory results due to the hash collision, we remove the $P_{m}(x,y,z)$ in Eq.
-12 and fuse the features $Q_{mij}(i,j)$ by concatenation in the implementation.
-The 3D+4D representation similar to [34] is yielded by concatenating the spatial $(x,y,z)$ features and spatial-temporal $(x,y,z,t)$ features encoded by a 4D extension of Eq.
-5.
-By comparing our CD-NGP with the naive 3D+4D representation, we show the advantage of our asymmetric encoding of spatial-temporal $(x,y,z,t)$ features explained in Eqs.
-3 and 6: representing the features of spatial coordinates $(x,y,z)$ and temporal coordinate $(t)$ separately avoids feature collisions and lead to better results.
-Table 5 also shows metrics of different feature fusion methods.
-We replace the element-wise sum in Eqs.
-6 and 12 with vector concatenation and denote them as Voxel-C and MERF-C in Table 5.
-The hybrid representations (MERF and MERF-C) achieve comparable model sizes ($\sim$60MB) to methods for static scenarios [31, 21] without significant quality loss, reducing the average bandwidth cost to $0.2$MB/frame.
-Benefiting from larger feature dimensions, the vector concatenation leads to slightly better quality (especially when measured in LPIPS) by more training time, and the reasonable trade-off also shows the robustness of our method.
-Compared with the hash table designs in Table 4, the Voxel-C representation in Table 5 almost performs as well as the bulky $(18,18)$ configuration in Table 4 and reduces the model size by $80\%$.
-We list the per-scene results of the proposed continual learning method with different spatial representations on the DyNeRF dataset in Table 6.
-Table 6: Detailed results of the proposed CD-NGP and other hash-based methods on the DyNeRF dataset.
-Model Flame Salmon Cook Spinach
-[…]
-CD-NGP 26.37 0.04586 0.2886 30.76 0.02440 0.1793
-Voxel-C 26.65 0.04243 0.2680 31.03 0.02288 0.1593
-Plane 25.48 0.06058 0.3173 29.41 0.03487 0.2010
-MERF 26.15 0.04952 0.2921 30.35 0.02931 0.1858
-MERF-C 26.56 0.04435 0.2546 30.69 0.02609 0.1583
-Model Flame Steak Sear Steak
-[…]
-CD-NGP 31.27 0.02132 0.1736 31.71 0.01917 0.1701
-Voxel-C 31.64 0.01823 0.1403 32.02 0.01698 0.1424
-Plane 30.26 0.02949 0.1804 30.56 0.02723 0.1811
-MERF 31.01 0.02508 0.1719 31.44 0.02181 0.1698
-MERF-C 31.13 0.02424 0.1506 31.66 0.01989 0.1413
-Model Cut Roasted Beef Average
-[…]
-CD-NGP 31.03 0.02366 0.1801 30.228 0.026882 0.19834
-Voxel-C 31.26 0.02205 0.1578 30.520 0.024514 0.17356
-Plane 30.11 0.03302 0.1951 29.164 0.037038 0.21498
-MERF 30.71 0.02765 0.1856 29.932 0.030674 0.20104
-MERF-C 31.04 0.02450 0.1632 30.216 0.027814 0.17360
-The detailed parameter numbers and bandwidth consumption are thus listed in Table 7.
-3D Enc.
-and 2D Enc.
-denote the spatial encodings along voxels and planes respectively.
-The tuples $(M_{1},M_{2})$ in the cells denote the parameter numbers of the base encoder and each of the auxiliary encoders.
-Following [27, 43, 1, 48] we also report the minimum per-frame bandwidth consumption $\mathbf{B_{min}}$, which only depends on the parameter numbers of the auxiliary branches and is thus significantly lower than the average bandwidth consumption $\mathbf{B_{avg}}$.
-Since the base encoder and the density grid could be shared among different branches and transmitted only once in the beginning, the minimum bandwidth consumption is reduced to $0.065\sim 0.16$ MB per frame, which is 100×100$\times$100 × less than the voxel-based StreamRF [27] and 7×7$\times$7 × less than the pure MLP representation INV [48].
-Table 7: The parameter numbers of each module of the 30-branch representations in Table 5.
-$\mathbf{B_{min}}$ and $\mathbf{B_{avg}}$ are minimum and average bandwidth costs in MB/frame.
-##\## denotes the number of parameters per branch (10 frames).
-Model ##\##3D Enc. ##\##2D Enc. ##\##MLP ##\##Online ##\##Total $\mathbf{B_{min}}$ $\mathbf{B_{avg}}$
-CD-NGP $(9.9M,0.37M)$ $0.042M$ $0.39M$ $21M$ 0.16 0.38
-Voxel-C $(9.9M,0.37M)$ $0.046M$ $0.40M$ $21M$ 0.16 0.38
-Plane - $(1.9M,0.082M)$ $0.048M$ $0.28M$ $14M$ 0.11 0.28
-MERF $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.042M$ $0.16M$ $7.5M$ 0.061 0.19
-MERF-C $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.052M$ $0.17M$ $7.5M$ 0.065 0.19
 3.4.2 Temporal representations.
@@
 3.4.3 Field composition methods.
-We also conduct ablation studies on the field composition methods and show the results in Tables 9 and 10. The field composition process is defined as:
+We also conduct ablation studies on the field composition methods and show the results in Tables 9 and 12. The field composition process is defined as:
 $(\sigma,\mathbf{c})=(\sigma_{i}+\sigma_{a},\mathbf{c}_{i}+\mathbf{c}_{a}).$ (13)
@@
 Furthermore, as compared in Fig.
-10, the compared composition methods in Figs.
-10(c) and 10(d) suffer from salt and pepper noise because of the explicit composition, while our CD-NGP circumvents the problem by fusing features in the latent space and produces better results.
+12, the compared composition methods in Figs.
+12(c) and 12(d) suffer from salt and pepper noise because of the explicit composition, while our CD-NGP circumvents the problem by fusing features in the latent space and produces better results.
 (a) Ground truth
@@
 Further efforts could be made to apply baking methods or Gaussian point clouds to continual learning to achieve real-time rendering in long video sequences for wide applications.
-Appendix A Detailed results on the DyNeRF dataset

### v5

@@
 Abstract
-Current methodologies for novel view synthesis (NVS) in dynamic scenes encounter significant challenges in harmonizing memory consumption, model complexity, training efficiency, and rendering fidelity.
-Existing offline techniques, while delivering high-quality results, are often characterized by substantial memory demands and limited scalability.
-In contrast, online methods grapple with the challenge of balancing rapid convergence with model compactness.
+Current methods for novel view synthesis (NVS) in dynamic scenes encounter significant challenges in managing memory consumption, model complexity, training efficiency, and rendering fidelity.
+Existing offline techniques, while delivering high-quality results, face challenges from substantial memory demands and limited scalability.
+Conversely, online methods struggle to balance rapid convergence with model compactness.
 To address these issues, we propose continual dynamic neural graphics primitives (CD-NGP).
-Our approach synergizes features from both temporal and spatial hash encodings to achieve high rendering quality, employs parameter reuse to enhance scalability, and leverages a continual learning framework to mitigate memory overhead.
-Furthermore, we introduce a novel dataset comprising multi-view, exceptionally long video sequences with substantial rigid and non-rigid motion, thereby substantiating the scalability of our method.
-We evaluate the quality, speed and scalability our method on both established public datasets and our exceptionally long video dataset.
+Our approach leverages a continual learning framework to reduce memory overhead, and it also integrates features from distinct temporal and spatial hash encodings for high rendering quality.
+Meanwhile, our method employs parameter reuse to achieve high scalability.
+Additionally, we introduce a novel dataset featuring multi-view, exceptionally long video sequences with substantial rigid and non-rigid motion, which is seldom possessed by existing datasets.
+We evaluate the reconstruction quality, speed and scalability of our method on both the established public datasets and our exceptionally long video dataset.
 Notably, our method achieves an $85\%$ reduction in training memory consumption (less than 14GB) compared to offline techniques and significantly lowers streaming bandwidth requirements (less than 0.4MB/frame) relative to other online alternatives.
-The experimental results on our long video sequences dataset show the superior scalability and quality compared to existing state-of-the-art approaches.
+The experimental results on our long video sequences dataset show the superior scalability and reconstruction quality compared to existing state-of-the-art approaches.
 keywords:
@@
 […]
-The demonstration of our continual learning pipeline for dynamic scenes.
-Our method segments input videos into multiple chunks, train the model chunk by chunk and leverage the re-use of parameters from the first branch to ensure high scalability.
-The spatial hash features could be encoded along voxels, orthogonally projected planes or both.
+The outline of our continual learning pipeline for dynamic scenes.
+Our method segments input videos into multiple chunks and encodes the chunks by separate spatial and time hash encodings.
+Subsequently, we train the auxiliary branches incrementally on each chunk and leverage the re-use of parameters in the base branch to ensure high scalability.
+The spatial hash features could be encoded according to voxel grids, orthogonally projected plane grids or a combination of both.
 Introduction
-Neural radiance field (NeRF) [30] has made significant progress in 3D scene reconstruction and novel view synthesis.
+Neural radiance field (NeRF) introduced in [30] has made significant progress in 3D scene reconstruction and novel view synthesis.
 It takes multi-view RGB images with their camera poses as input and leverages multi-layer perceptrons (MLPs) to reconstruct both the geometry and color of the scene.
@@
 For example, the methods in [36, 28, 34, 33, 15, 29, 44] propose to use MLP-based representation and achieve high-quality results.
-There are also methods speed up the optimization by using tensor factorization ([40, 17, 7]) and voxel representation [16, 45].
-The more recent continual-learning-based methods [27, 48] can train models frame by frame, and streamable representations [43] enable the model to be loaded on the fly at inference.
-However, the offline methods require huge host memory e.g., taking more than $100GB$ when reconstructing multi-view dynamic videos of 300 frames, and the continual learning-based methods are weak in balancing the convergence speed and model compactness.
-As a result, these methods are limited in scalability, making them unsuitable for long videos.
+There are also methods that speed up the optimization by using the tensor factorization ([40, 17, 7]) and voxel representations [16, 45].
+The more recent continual-learning-based methods [27, 48] (also mentioned as online methods) can train models frame by frame, and streamable representations [43] enable the model to be loaded on the fly at inference.
+However, the offline methods require huge host memory e.g., taking more than $100GB$ when reconstructing multi-view dynamic videos of 300 frames, and the online methods are weak in balancing the convergence speed and model compactness.
+These limitations in scalability makes them impractical for the wide application on long videos.
 In this paper, we propose continual dynamic neural graphics primitives (CD-NGP), a fast and scalable representation for reconstructing dynamic scenes, which enables dynamic reconstruction in long videos.
-The proposed methodology is illustrated in CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes: we partition dynamic videos into multiple segments based on timestamps and perform reconstruction using distinct model branches.
-Each branch includes MLP regressors, spatial hash encoding for spatial coordinates $(x,y,z)$, and temporal hash encoding for the timestamp $t$.
-For the first segment, we encode spatial features using a large hash table, and for subsequent segments, we use a smaller hash table.
-By reusing parameters from the initial branch, we enhance scalability and accelerate convergence.
-For each segment, we identify the corresponding model branch, sample 3D points $(x,y,z)$ according to frame pixels and cached occupancy grids [31], and represent spatial features using spatial hash grids.
-The spatial features from the initial branch and the current branch are fused through element-wise summation by default or concatenation if applicable.
-Temporal features are encoded by the hash table for the timestamp $t$ within the current branch only.
-These spatial and temporal features, along with the viewing direction $d$, are concatenated and input into MLP regressors for rendering and optimization.
-We validate the effectiveness of our method through comprehensive experimental results on both the widely used public datasets and the dataset proposed by us.
-As illustrated in Fig.
-1, our method reaches a new balance between rendering quality, model size, and training time while remaining memory-friendly, in which we use metrics on the flame salmon scene in the DyNeRF dataset for fair comparison.
+As illustrated in CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes, CD-NGP derives from three key ideas: dividing the video into segments, encoding spatial and temporal features separately, and reusing features across different segments to ensure high scalability: Firstly, we partition dynamic videos into multiple segments according to time stamps $t$ and perform reconstruction using distinct model branches to facilitate continual learning.
+Each branch of our model includes MLP regressors, spatial hash encoding for spatial coordinates $(x,y,z)$, and temporal hash encoding for the timestamp $t$.
+Secondly, the spatial and temporal features are encoded separately to avoid feature confusion in different axes.
+For each model branch, we sample 3D points $(x,y,z)$ according to frame pixels and cached occupancy grids like [31], and encode the points into spatial features using spatial hash grids.
+We also encode temporal features using hash grids for the separated $t$ axis.
+These spatial and temporal features and the viewing direction $\mathbf{d}$ are concatenated and input into MLP regressors for rendering and optimization.
+Thirdly, we leverage the redundancy in the dynamic scenes and fuse the spatial features of different branches to yield high scalability.
+We set a larger hash table size ($2^{19}$) in the initial branch (base branch), and set a smaller size ($2^{14}$) for the other branches (auxiliary branches).
+This asymetric design facilitates reusing the features trained in the base branch, thereby reducing storage consumption and boosting rendering quality.
+Figure 1: Comparison between different performance indexes of our method and recent representative baselines.
 The perceptual consistency is measured by $1-$ LPIPS by AlexNet [57, 24], and the speed is the reciprocal of the training time.
 The compactness and RAM-Efficiency are mapped by $-log$ function from the model size and memory occupation respectively.
-Figure 1: Comparison between different performance indexes of our method and recent representative baselines.
+We validate the effectiveness of our method through comprehensive experimental results on both the widely used public datasets and a self-collected dataset.
+As illustrated in Fig.
+1, our method reaches a new balance between rendering quality, model size, and training time while remaining memory-friendly.
 Compared with prevailing offline baselines, our method consumes $85\%$ less host memory at the training phase, $40\%$ less model size at the inference phase, and delivers comparable results.
-Compared with online baselines, our method achieves compact model size, high quality, and fast convergence simultaneously.
-Contributions of this paper are three-fold as follows:
-  • We propose a novel method combining spatial and temporal hash encodings for dynamic scenes and yield comparable results to both offline and online methods.
-  • We propose a fast continual learning framework for dynamic scenes and reduce memory and storage occupation significantly.
-  • We provide a dataset of multi-view long video sequences ($>3$ minutes) with large motion and further demonstrate the scalability of our method on the proposed dataset.
+Compared with online baselines, our method achieves compact model size, high rendering quality, and fast convergence simultaneously.
+The contributions of this paper are three-fold as follows:
+  • We propose a novel method combining separate spatial and temporal hash encodings, achieving high scalability and rendering quality in dynamic scenes.
+  • We propose a fast and highly scalable continual learning framework for dynamic scenes.
+It enables reconstructing long videos with low memory usage and low storage consumption.
+Our method also outperforms online alternatives and achieves comparable results to offline methods.
+  • We provide a dataset of multi-view exceptionally long video sequences ($>3$ minutes) with large non-rigid motion, which is seldom possessed by existing datasets.
 1 Related Work
@@
 They alpha-compose the cached features along the ray and use MLPs only once per ray to overcome the computation load of the per-point MLP querying in NeRF.
-BakedSDF [56] stores the view-dependent features into Gaussian spheres along the reconstructed surfaces, and achieves real-time rendering on ordinary laptops.
+BakedSDF [56] stores the view-dependent features into Gaussian spheres along the reconstructed surfaces, and it achieves real-time rendering on ordinary laptops.
 MobileNeRF[10] represents scenes into meshes and enables real-time rendering on mobile devices.
@@
 The third class is the explicit voxel-based representations.
-TineuVox [16] leverages the deformation MLP in D-NeRF and implements a 3D voxel grid with multi-scale interpolation.
-It achieves high fidelity and fast convergence on the D-NeRF dataset.
+TineuVox [16] leverages the deformation MLP in D-NeRF and implements a 3D voxel grid with multi-scale interpolation, and it achieves high fidelity and fast convergence on the D-NeRF dataset.
 MixVoxels [45] leverages the standard deviation across the time of all pixels to estimate a variation field to decompose the dynamic scene into static and dynamic voxels.
@@
 Their explicit voxel representation enables fast reconstruction (75 minutes for 300 frames) and could be streamed frame by frame.
-INV [48] uses a shared color MLP and trains different structure MLPs frame by frame to form a continual NeRF representation and achieves high-fidelity results in dynamic scenes.
+INV [48] uses a shared color MLP and trains different structure MLPs frame by frame to form a continual NeRF representation.
+It achieves high reconstruction quality in dynamic scenes.
 2 Method
-2.1 Preliminaries
-Inspired by the previous NeRF variants [36, 28, 17, 7, 45, 31, 38] and continual learning methods [6, 35, 11, 27, 43, 48], we focus on the dynamic novel view synthesis problem in the multi-view setting and solve the problem in a continual learning manner to alleviate the huge memory consumption and bandwidth costs.
-Following previous methods [36, 28, 17, 7], we consider the dynamic scene as a set of 3D points $\mathbf{p}=(x,y,z)$ with optical attributes $(\sigma,\mathbf{c})$, where $\sigma$ is the volume density of the point and $\mathbf{c}$ is the color of the point.
-We train our model along camera rays to reconstruct the scene.
-A camera ray is defined as:
+2.1 Problem definition and preliminaries
+[…]
+Inspired by the previous NeRF variants [36, 28, 17, 7, 45, 31, 38] for dynamic novel view synthesis and continual learning methods [6, 35, 11, 27, 43, 48] in the static scenes, we focus on the dynamic novel view synthesis problem in a multi-view setting.
+We solve the problem using continual learning to alleviate the huge memory consumption and bandwidth costs.
+Unlike the existing replay-based continual learning methods for static scenarios [6, 35], we solve the continual learning problem of dynamic novel view synthesis via parameter isolation.
+This is because the current offline state-of-the-art methods by tensor factorization [7, 17] or explicit voxels [45] are regressions of the optical attributes in 3D space within a limited time instead of the object trajectory and inherent dynamics.
+Therefore, the replay method is of limited use in our defined scenario: Since new transient processes are constantly arriving, the model of a limited size is not capable of reconstructing unlimited transient effects.
+Thus, the replay-based method suffers from catastrophic forgetting.
+The design of our model follow the parameter isolation approach: Let the time stamps be divided into $N_{chunk}$ chunks, the model consists of $N_{chunk}$ branches to represent different chunks of the dynamic scene accordingly.
+Following previous methods [36, 28, 17, 7], we consider the dynamic scene as a set of 3D points $\mathbf{x}=(x,y,z)$ with optical attributes $(\sigma,\mathbf{c})$, where $\sigma$ is the volume density of the point and $\mathbf{c}$ is the color of the point.
+To reconstruct the scene, we train our model along camera rays which are defined as:
 $\mathbf{r}(u)=\mathbf{o}+u\mathbf{d},$ (1)
-where $\mathbf{o}$ is the optical center of the camera, $\mathbf{d}$ is the direction off the camera.
-For each point along the camera ray with coordinates $(x,y,z)$ observed at direction $\mathbf{d}=(\phi,\psi)$, a field function $F$ is queried to obtain the optical attributes $(\sigma,\mathbf{c})$.
+where $\mathbf{o}$ is the optical center of the camera, $\mathbf{d}$ is the direction from the camera to the scenes.
+For each point along the camera ray with coordinates $\mathbf{x}=(x,y,z)$ observed at direction $\mathbf{d}=(\phi,\psi)$, a field function $F$ is queried to obtain the optical attributes $(\sigma,\mathbf{c})$.
 Conditioned on time $t$, the entire model is defined as:
@@
 where the function $F$ denotes the dynamic NeRF model and $\Theta$ denotes trainable model parameters.
-[…]
-To simplify the notations, let the whole sequence of frames be evenly divided into $k$ chunks where a chunk $\mathbf{T_{k}}=\{\tau_{i}\leavevmode\nobreak\ |\leavevmode\nobreak\ \lfloor\frac{N_{time}}{T_{chunk}}\rfloor=k\}$ is a sequential subset of length $T_{chunk}$.
-We define the continual learning problem as learning a representation where only the frames in the current chunk are available at each optimization step, i.e.
-$\mathbf{T_{a}}=\mathbf{T_{k}}$.
-Accordingly, $T_{chunk}=1$ denotes training the model online in a frame-by-frame manner, $T_{chunk}=10$ denotes training the model online with a buffer of 10 frames.
-If the video contains 300 frames in total, the $T_{chunk}=300$ denotes training the entire 300-frame sequence offline.
-The definition is close to real-world applications since caching all training frames in memory or storing them on the disk is limited in scalability for long videos.
 2.2 Representation of short dynamic intervals
+To reconstruct the individual chunks of the dynamic scenes, we propose a representation of short dynamic intervals, which serves as the structures of the different branches of our model and handle the dynamic scnes in one chunk efficiently.
 As shown in [46], naive 4D hash grids for dynamic scenes are short in reconstruction quality.
-To achieve fast reconstruction and slim model size simultaneously, we represent time and spatial features separately and concatenate them to alleviate feature confusion.
-The spatial hash features may be encoded along 3D voxels [31], 2D planes [17, 21], or both [38].
+To achieve fast reconstruction and compact model size simultaneously, we represent time and spatial features separately and concatenate them to alleviate feature confusion.
+The spatial hash features can be encoded along 3D voxels [31], 2D planes [17, 21], or both [38].
 The time feature is encoded by short hash tables in a per-frame manner.
@@
 $(\sigma,\mathbf{H})=f_{\theta_{1}}(\Psi(\mathbf{x})\oplus\Gamma(t)),$ (3)
-$\mathbf{c}=f_{\theta_{2}}(\mathbf{H},\phi,\psi).$ (4)
+$\mathbf{c}=f_{\theta_{2}}(\mathbf{H},\phi,\psi),$ (4)
 where $\mathbf{x}=(x,y,z)$ is the spatial coordinates, $\Psi(\mathbf{x})$ denotes configurable spatial hash encodings, $f_{{\theta_{1}}}$ is a tiny MLP with trainable parameter ${\theta_{1}}$, $\Gamma(t)$ is the hash encoding for time axis alone, and $\oplus$ is the concatenation of vectors.
@@
 4.
-The spatial encoders $\Psi(x)$ follow a multi-resolution concatenated format, defined as:
-$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}\leavevmode\nobreak\ P_{m}(x,y,z)$ (5)
-where the $P_{m}(x,y,z)\in\mathbf{R}^{F}$ denotes the tri-linearly interpolated value from the nearby 8 grid vertices w.r.t.
-3D coordinates $(x,y,z)$ at the $l$-th resolution, $L$ is the number of the resolutions, and $F$ is the dimensions of features $P_{m}(x,y,z)$, which are similar to those in Instant-NGP [31].
-If necessary, Eq.
-5 could be adopted to the hybrid (MERF) representation [38], defined as:
-$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}\leavevmode\nobreak\ P_{m}(x,y,z)+Q_{mxy}(x,y)+Q_{myz}(y,z)+Q_{mzx}(z,x)$ (6)
-where $Q_{mij}(i,j)$ denotes the bi-linearly interpolated value from the 4 nearby vertices in the orthogonally projected planes w.r.t.
-axes $(i,j)$ at the $m$-th resolution.
-Removing the $P_{m}(x,y,z)$ term leads to a pure plane representation like [17, 21].
+The spatial hash encoders $\Psi(x)$ follow a multi-resolution concatenated format, defined as:
+$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}\leavevmode\nobreak\ P_{m}(x,y,z),$ (5)
+where the $P_{m}(x,y,z)\in\mathbf{R}^{F}$ denotes the tri-linearly interpolated feature from the nearby 8 grid vertices w.r.t.
+3D coordinates $(x,y,z)$ at the $m$-th resolution, $L$ is the number of the resolutions, and $F$ is the dimensions of features $P_{m}(x,y,z)$, which are similar to those in Instant-NGP [31].
+The voxel-based hash encoding structure is flexible and could be replaced by other spatial hash encodings, which will be discussed in the experiments.
 2.3 Continual dynamic neural graphics primitives (CD-NGP)
-Unlike the existing replay-based methods for static scenarios [6, 35], we solve the continual learning problem of dynamic novel view synthesis via parameter isolation.
-The current state-of-the-art methods by tensor factorization [7, 17] or explicit voxels [45] are regressions of the optical attributes in 3D space within a limited time instead of the object trajectory and inherent dynamics.
-Therefore, the replay method is of limited use in our defined scenario: Since new transient processes are constantly arriving, the model of a limited size is not capable of reconstructing unlimited transient effects.
-Thus, the replay-based method suffers from catastrophic forgetting.
-To address this issue, we propose our continual dynamic neural graphics primitives (CD-NGP) based on parameter isolation.
-We split the model into $N_{chunk}$ branches to represent different chunks of the dynamic scene accordingly.
-The $k$-th branch contains spatial hash encoder $\Psi_{k}(\mathbf{x})$, temporal hash encoder $\Gamma_{k}(t)$, and MLPs $f_{\theta_{1k}},f_{\theta_{2k}}$.
-We train the first branch (base branch) with a large spatial hash table to better reconstruct the scenes with initial movements and static objects.
-Then we reuse the spatial features from the base branch and train the subsequent branches (auxiliary branches) to compensate for residuals.
-We thus substitute the $\Psi(\mathbf{x})$ in Eq.
-3 with the fused feature to obtain the volume density $\sigma$ at the $k$-th model branch under the continual learning setting as:
+As defined in Section 2.1, we split the model into $N_{chunk}$ individual branches to represent the $N_{chunk}$ chunks of the dynamic scene.
+Consequently, the spatial features encoded by different model branches could be fused to reduce the number of the parameters and improve scalability.
+This approach, defined as continual dynamic neural graphics primitives (CD-NGP), is designed to handle the dynamic nature of scenes in the online setting.
+Equipped with the hash encoders and MLP regressors defined in Section 2.2, the $k$-th branch of the model contains spatial hash encoder $\Psi_{k}(\mathbf{x})$, temporal hash encoder $\Gamma_{k}(t)$, and MLPs $f_{\theta_{1k}},f_{\theta_{2k}}$.
+Please note the dimensions of the features is independent of the sizes of the hash tables: a larger hash table alleviates hash collision at the cost of more storage consumption, but the features encoded by different sizes of hash tables share the same size, i.e.
+$\Psi_{k}(x)\in\mathbf{R}^{LF}$ holds for all $k$.
+To alleviate hash collision and reduce model size simultaneously, we apply an asymmetric hash table design: the hash table in the first branch (base branch) is set as $2^{P_{1}}$, and the hash tables in the subsequent branches (auxiliary branches) are set as $2^{P_{2}}$, where $P_{1}$ is much larger than $P_{2}$.
+The base branch is designed to reconstruct the scenes with initial movements and static objects, and the auxiliary branches are designed to compensate for residuals.
+The re-use of the features in the base branch ensures high scalability and alleviates the catastrophic forgetting problem.
+When ($P_{1},P_{2})=(19,14)$, the hash table in the second branch consumes only $1/32$ of the parameters in the first branch.
+We substitute the $\Psi(\mathbf{x})$ in Eq.
+3 with the fused feature to obtain the volume density $\sigma$ at the $k$-th model branch under the continual learning setting, defined as:
 $(\sigma,\mathbf{H})=f_{\theta_{1k}}((\Psi_{0}(\mathbf{x})+\Psi_{k}(\mathbf{x}))\oplus\Gamma_{k}(t)),$ (6)
 where $\Psi_{0}(\mathbf{x})$ denotes the spatial encoder of the base branch, $\Psi_{k}(\mathbf{x})$ denotes the spatial encoder of the current branch.
-The view-dependent colors are regressed as Eq.
+The view-dependent colors are regressed as in Eq.
 4 with the color MLP $f_{\theta_{2k}}$.
-For simplicity, we denote the models adopting pure voxel encoders in Eq.
-5 as CD-NGP, the models adopting the hybrid encoders in Eq.
-6 as CD-MERF.
-The scalability of our method derives from the redundancy of dynamic scenes and the flexibility of hash table sizes.
-As pointed out by [45], most points ($>90\%$) in the dynamic scenes are static.
-We thus leverage the shared static points across time and represent the first chunk with a large hash table and the subsequent chunks with small hash tables to ensure scalability.
-To be specific, if we follow the default setting of Instant-NGP [31] and set the hash table in the first branch as $2^{P_{1}}$, the sizes of hash tables in subsequent branches could be configured as $2^{P_{2}}$.
-Please note that $P_{1}$ is significantly larger than $P_{2}$ to ensure the scalability of the model.
-When ($P_{1},P_{2})=(19,14)$, the hash table in the second branch consumes only $1/32$ of the parameters in the first branch.
-Since the spatial hash tables occupy the most parameters in the model, and most chunks are encoded by the subsequent branches, our assymetric hash table design enables continual learning of dynamic scenes and high scalability in streaming scenarios simultaneously.
-The total bandwidth consumption of our method is $\mathcal{O}(N_{chunk}\cdot 2^{P_{2}}LF+2^{P_{1}}LF)$, while the previous method [43] requires offline training and consumes $\mathcal{O}(\mathbf{K}\cdot 2^{P_{1}}LF)$ bandwidth, where $\mathbf{K}$ ranging from $0.5$ to $16$ is their bandwidth budget controlling the bitrates in streaming scenarios.
-If the bandwidth is extremely limited, the hash tables in our model could be arranged as a combination of 3D representations and 2D representations like [38] to reduce the bandwidth cost additionally.
-2.4 Rendering and optimization
-Given the sampled optical attributes on the rays in Eq. 1, we adopt volume rendering to get the 2D pixels, defined as:
-$\hat{C}(r)$ $=\sum_{i=1}^{N}T_{i}\alpha_{i}\mathbf{c_{i}}=\sum_{i=1}^{N}T_{i}(1-e^{\sigma_{i}\delta_{i}})\mathbf{c_{i}},$ (8)
-where $T_{i}=\exp(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j})$ denotes the transmittance, $\delta_{i}$ is the interval between neighboring samples, $N$ is the number of samples along the ray.
-The model is optimized by L2 photometric loss with three regularization terms, defined as:
-$L=\frac{1}{|\mathbf{R}|}\sum_{r\in\mathbf{R}}\|C(r)-\hat{C}(r)\|_{2}^{2}+\lambda_{d}L_{d}+\lambda_{o}L_{o}+\lambda_{r}L_{r}$ (9)
-Where $\mathbf{R}$ denotes the set of the camera rays. $L_{d}$ is the distortion loss in Mip-NeRF 360 [3], defined as:
-$L_{d}$ $=\sum_{i,j}w_{i}w_{j}\left\lvert\frac{s_{i}+s_{i+1}}{2}-\frac{s_{j}+s_{j+1}}{2}\right\rvert$
-$+\frac{1}{3}\sum_{i}w_{i}^{2}(s_{i+1}-s_{i})$ (10)
-where $s_{i}$ denotes the normalized ray distance along the ray, $w_{i}=T_{i}\alpha_{i}$ is the weight of the $i$-th sample.
-$L_{o}$ in Eq. 9 is the entropy on the opacity by [25], defined as:
-$L_{o}=-o\log(o),\quad o=\sum_{j=1}^{N}T_{j}\alpha_{j}.$ (11)
-$L_{r}$ in Eq. 9 is the L1 regularization on the spatial features in the current branch as:
-$L_{r}=||\Psi_{k}(\mathbf{x})||_{1}.$ (12)
-We set $\lambda_{r}$ as 0.001 to encourage leverage of the base hash encoder. Based on [25], we set a larger $\lambda_{d},\lambda_{o}$ as 0.005 to alleviate foggy effects.
-2.5 Continual learning pipeline.
+To cope with large motions and complex transient effects in the long time intervals, we assign $T_{episode}$ chunks as an episode.
+The MLP regressors are re-trained from initialization in the first branch of each episode to adapt to recent changes.
 We further illustrate the whole process of our method in Algorithm 1, where $\eta_{init}$ and $\eta_{aux}$ denote the numbers of iterations of the initial branch and the auxiliary branches, respectively.
 $\Psi_{i},\Gamma_{i},\theta_{1}^{(i)},\theta_{2}^{(i)}$ denote the model parameters in the branches.
-The whole model is trained chunk by chunk.
-The implementation will be available upon publication.
 Algorithm 1 Continual learning pipeline
-0:  $N_{chunk},T_{chunk},\eta_{init},\eta_{aux},\Psi_{i},\Gamma_{i},\theta_{1}^{(i)},\theta_{2}^{(i)},i\in\mathbb{N}$
+0:  $N_{chunk},T_{chunk},T_{episode},\eta_{init},\eta_{aux},\Psi_{i},\Gamma_{i},\theta_{1}^{(i)},\theta_{2}^{(i)},i\in\mathbb{N}$
 0:  The whole representation for dynamic scenes.
 […]
-2:  while $N_{chunk}$ do
+2:  while $k$<$N_{chunk}$ do
 […]
@@
 13:  end while
+The scalability of our method derives from the redundancy of dynamic scenes and the flexibility of hash table sizes.
+As pointed out by [45], most points ($>90\%$ in DyNeRF dataset by their observation) in the dynamic scenes are static.
+We thus leverage the shared static points across time and represent the first chunk with a large hash table and the subsequent chunks with small hash tables to ensure scalability.
+Since the spatial hash tables occupy the most parameters in the model, and most chunks are encoded by the subsequent branches, our asymetric hash table design enables continual learning of dynamic scenes and high scalability in streaming scenarios simultaneously.
+The total bandwidth consumption of our method is $\mathcal{O}(N_{chunk}\cdot 2^{P_{2}}LF+2^{P_{1}}LF)$, while the previous method [43] requires offline training and consumes $\mathcal{O}(\mathbf{K}\cdot 2^{P_{1}}LF)$ bandwidth, where $\mathbf{K}$ ranging from $0.5$ to $16$ is their bandwidth budget controlling the bitrates in streaming scenarios.
+If the bandwidth is extremely limited, the hash tables in our model could be arranged as a combination of 3D representations and 2D representations like [38] to reduce the bandwidth cost additionally.
+2.4 Rendering and optimization
+Given the sampled optical attributes on the rays in Eq. 1, we adopt volume rendering to get the 2D pixels, defined as:
+$\hat{C}(r)$ $=\sum_{i=1}^{N}T_{i}\alpha_{i}\mathbf{c_{i}}=\sum_{i=1}^{N}T_{i}(1-e^{-\sigma_{i}\delta_{i}})\mathbf{c_{i}},$ (7)
+where $T_{i}=\exp(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j})$ denotes the transmittance, $\delta_{i}$ is the interval between neighboring samples, $N$ is the number of samples along the ray, $\mathbf{c_{i}}$ is the output of Eq.
+4.
+The model is optimized by L2 photometric loss with three regularization terms, defined as:
+$L=\frac{1}{|\mathbf{R}|}\sum_{r\in\mathbf{R}}\|C(r)-\hat{C}(r)\|_{2}^{2}+\lambda_{d}L_{d}+\lambda_{o}L_{o}+\lambda_{r}L_{r},$ (8)
+where $\mathbf{R}$ denotes the set of the camera rays. $L_{d}$ is the distortion loss in Mip-NeRF 360 [3], defined as:
+$L_{d}$ $=\sum_{i,j}w_{i}w_{j}\left\lvert\frac{s_{i}+s_{i+1}}{2}-\frac{s_{j}+s_{j+1}}{2}\right\rvert$
+$+\frac{1}{3}\sum_{i}w_{i}^{2}(s_{i+1}-s_{i}),$ (9)
+where $s_{i}$ denotes the normalized ray distance along the ray, $w_{i}=T_{i}\alpha_{i}$ is the weight of the $i$-th sample.
+$L_{o}$ in Eq. 8 is the entropy on the opacity [25], defined as:
+$L_{o}=-o\log(o),\quad o=\sum_{j=1}^{N}T_{j}\alpha_{j},$ (10)
+$L_{r}$ in Eq. 8 is the L1 regularization on the spatial features in the current branch as:
+$L_{r}=||\Psi_{k}(\mathbf{x})||_{1}.$ (11)
+We set $\lambda_{r}$ as 0.001 to encourage using the features encoded by the base branch. Based on [25], we set a large value of $\lambda_{d},\lambda_{o}$ as 0.005 to alleviate foggy effects.
+Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 2: Example frames sampled from the proposed long multi-view video dataset.
 3 Experiments
@@
 Meetroom dataset [27] is another dataset for dynamic view synthesis that contains 3 scenes captured by 13 synchronized cameras at $1280\times 720$ resolution.
-The lengths of the scenes are set as 300 frames and the refresh rates are set as 30 FPS.
-To further validate the effectiveness of our method on long video sequences, we introduce a new dataset comprising 5 dynamic scenes at a resolution of $3840\times 2160$.
-Similar to the DyNeRF dataset [28], these scenes are captured by 20 GoPro cameras operating at 30 FPS.
-We synchronize the multi-view videos using environmental sound and incorporate chessboard images within the scenes to facilitate camera calibration and pose estimation.
-The dynamics in our dataset are notably more challenging than those in the DyNeRF dataset, featuring significant non-rigid motions of human bodies and transient effects such as slide effects on the screens in the backgrounds.
+The scenes are 300 frames long and captured at 30 FPS.
+To further validate the effectiveness of our method on long video sequences, we introduce a new dataset comprising 5 dynamic scenes at a resolution of $3840\times 2160$ as shown in Fig.
+2.
+Similar to those in the DyNeRF dataset [28], these scenes are captured by 20 GoPro cameras operating at 30 FPS.
+The cameras are arranged in two rows and placed in an arc facing the scenes.
+The camera in the center of the upper row is designated for validation.
+We synchronize the multi-view videos and incorporate chessboard images within the scenes to facilitate camera calibration and pose estimation.
+Our dataset features more challenging dynamics, including significant non-rigid human motions and transient effects like screen slides, compared to the DyNeRF dataset.
 To comprehensively assess the scalability of our method, we provide 5 scenes with a duration exceeding 200 seconds (6000 frames), which is over 10 times longer than the previous DyNeRF dataset.
-This dataset was meticulously collected to evaluate the scalability of our method and to support further research.
+Our dataset is collected to evaluate the scalability of our method and to support further research in the community.
 We have carefully reviewed the videos to ensure they do not contain any controversial or offensive content before using them for research purposes.
@@
 All videos were collected with the explicit consent of all participants.
-The view in the upper center is designated for validation.
-Figure 2: Example frames sampled from the proposed long multi-view video dataset.
 3.2 Implementation details
@@
 Similar to the previous work [7, 28], we select the view in the center with index 00 for evaluation on both datasets.
-Following most of the state-of-the-art methods [28, 27, 17, 45, 43] we train and evaluate the model at resolution $1352\times 1014$ on the DyNeRF dataset.
-We use $1280\times 720$ resolution for our dataset to fit into memory.
-We use the importance sampling scheme in [28] at a batch size of 1024 rays.
-The parameters are optimized by Adam [23], with $\epsilon=10^{-15},\beta=(0.9,0.96)$.
-The learning rate is set as 0.001 and scheduled by cosine function following [25].
-We set the size of the hash table in the first branch as $2^{19}$ and the size of the hash table in the subsequent branches as $2^{14}$, and we set the structural parameters $(L,F)$ in the spatial hash encoders as $(12,2)$ if not otherwise specified.
-Following MERF [38] for static scenes, we use a smaller hash table for the hybrid representation in Eq.
-6 than pure voxel hash grid [31] representation in Table 3 to ensure the compactness and scalability of our method.
-Denoting the lengths of hash tables along pure voxels in a branch of CD-NGP as $2^{P}$, the lengths of hash tables in the CD-MERF are set as $2^{P-3}$ and $2^{P-4}$ for voxel and plane encoders respectively.
-The number of the parameters of the MERF representation is thus reduced to $1-2^{-3}-3\times 2^{-4}=31.25\%$ of the voxel hash grids.
-Since the hash tables take up the majority of the parameters, the model size is reduced from $21M$ to $7.5M$ as in Table 3.
-At the base branch, we initialize the models and train for 18k iterations, and for each of the subsequent branches we train for 3k iterations using the previous branch for initialization.
+Following most of the state-of-the-art methods [28, 27, 17, 45, 43], we train and evaluate the model at resolution $1352\times 1014$ on the DyNeRF dataset.
+We follow [27] and use $1280\times 720$ resolution on the Meetroom dataset.
+We use $1280\times 720$ resolution for our dataset to fit into the host memory.
+We use the importance sampling approach described in [28] at a batch size of 1024 rays.
+The model parameters are optimized by Adam [23], with hyper parameters $\epsilon=10^{-15},\beta=(0.9,0.96)$ for stability.
+The learning rate is set as 0.001 and scheduled by cosine function.
+We set the hash table sizes as $2^{P_{1}}=2^{19}$ in the first branch and $2^{P_{2}}=2^{14}$ in the subsequent branches.
+The structural parameters $(L,F)$ in the spatial hash encoders are set as $(12,2)$ if not otherwise specified.
+At the base branch, we initialize the models and train for $\eta_{init}=18k$ iterations, and for each of the subsequent branches we train for $\eta_{aux}=3k$ iterations using the previous branch for initialization for the DyNeRF dataset and our dataset.
+We use $\eta_{init}=33k$ and $\eta_{aux}=3k$ for the Meetroom dataset.
 We use LeakyReLU activation in the MLP regressors.
-For long video sequences, we assign 30 chunks as an episiode.
-The MLP regressors are re-trained from initialization in the first branch of each episode to adapt to recent changes.
 Strategies for updating the occupancy grids are identical with [31].
@@
 We report the model sizes with the occupancy grid under bfloat16 precision and parameters in float32 precision.
+The episode length $T_{episode}$ is set as $30$ only for videos longer than 300 frames in our own dataset.
 We set the dimensions of the latent feature $\mathbf{H}$ in Eq.
@@
 We use MLPs with identical structures for different branches.
-We set the resolutions of the hash encoders of the CD-NGP within $(16,2048)$ times the scale of the scene, and we use $(64,2048)$ for the plane representation.
-3.3 Results and comparisons
-Table 1: Comparisons between our CD-NGP and SOTA methods on the DyNeRF dataset.
-Memory denotes the maximum host memory required for training.
-- denotes the metrics unreported in the original publications.
-Best, Second, and Third metrics are in highlight.
-The full model of [27] results in much worse metrics than those reported in their paper, so we only report the metrics of their pilot model.
+We set the resolutions of the hash encoders of the CD-NGP within $(16,2048)$ times the scale of the scene.
+3.3 Results and scalability analysis
+Table 1: Comparisons between our CD-NGP and state-of-the-art methods on the DyNeRF dataset.
+Memory denotes the maximum host memory required during training.
+$\dagger$ denotes the methods evaluated only on the flame salmon scene.
+- denotes the metrics not reported in the original publications.
 […]
@@
 K-Planes [17] $\checkmark$ 300 31.63 - - 110 min 200MB >100GB
-MixVoxels-T [45] $\checkmark$ 300 30.71 0.024 0.162 15 min 500MB >100GB
+MixVoxels [45] $\checkmark$ 300 30.71 0.024 0.162 15 min 500MB >100GB
 4DGS [55] $\checkmark$ 300 32.01 - 0.055 6 hours 4000MB <10GB
@@
 INV [48] $\checkmark$ 1 29.64 0.023 0.078 40 hours 336MB <10GB
-CD-NGP (Ours) $\checkmark$ 10 30.23 0.029 0.198 75 min 113MB 14GB
-3.3.1 DyNeRF dataset.
+CD-NGP (ours) $\checkmark$ 10 30.23 0.027 0.198 75 min 113MB 14GB
+3.3.1 Comparisons on DyNeRF dataset.
 Table 1 shows the quantitative comparisons between CD-NGP and other state-of-the-art methods on DyNeRF dataset.
-We use the officially released metrics.
-The methods in [28, 34] are trained on 8 V100 GPUs or equally 4 A100 GPUs.
+The methods in [28, 34] are trained on 8 V100 GPUs or equivalently 4 A100 GPUs.
 The methods in [28, 34, 43] have no official implementations available.
 We are unable to obtain the missing quality metrics in [17] because of their huge memory occupation in the default resolution.
-HexPlane [7] uses resolution $(1024,768)$ in their implementation, so the memory occupation is much smaller than other offline methods.
+We report the metrics in the publication if not otherwise specified.
+HexPlane [7] uses resolution $1024\times 768$ in their implementation, so the memory occupation is much smaller than other offline methods.
 We use the official implementation of StreamRF [27] and INV [48] and report the average metrics on the 5 synchronized scenes of the dataset.
-StreamRF [27] and INV [48] are trained per-frame.
+StreamRF [27] and INV [48] are trained per frame.
 NeRFplayer [43] is trained offline but streamable upon inference.
 4DGS [55] caches the frames onto disk for training to alleviate memory occupation with a trade-off on training speed.
-The reconstruction quality is measured by PSNR, DSSIM, and LPIPS [57] (by AlexNet [24]) following common practices [28, 7, 45].
+The reconstruction quality is measured by PSNR for pixel-wise similarities, DSSIM for structural similarities, and LPIPS [57] for perceptual similarities (by AlexNet [24]) following common practices [28, 7, 45].
 To prove the viability of our parameter-isolation-based method, we apply the previous state-of-the-art method HexPlane [7] to continual learning setting with parameter-regularization (HexPlane-PR) and generative replay (HexPlane-GR) as additional baselines.
+We compare our method with both online and offline baselines.
+As demonstrated in Table 1, our CD-NGP is 32×32$\times$32 × faster than the MLP-based continual representation INV [48], $\times4$ smaller than the voxel-based continual representation StreamRF [27] and significantly outperforms the two additional baselines in the continual learning setting.
+The undesirable results from HexPlane-PR and HexPlane-GR show the limitations of the regularization-based and replay-based methods in dynamic scenes: unlike the static scenes, the objects are not stationary across time and the models with limited complexity are not capable of reconstructing unlimited transient effects.
+Benefiting from the parameter isolation like StreamRF [27] and INV [48], our method achieves fast convergence and compact model size simultaneously.
+It leverages the hash tables to ensure fast convergence and reuses the features from differently configured spatial hash tables to ensure high scalability.
+Compared with prevailing offline methods [28, 34, 7, 17, 45, 55], our method only requires caching $T_{chunk}=10$ frames into the buffer by continual learning and thus reduces memory occupation by $85\%$.
+Our method also reduces $40\%$ less model size (to $0.4$MB/frame) by leveraging the feature fusion of the base and the auxiliary branches and delivers comparable quality to the offline methods.
 (a) HexPlane-offline
-(b) MixVoxels-T
+(b) MixVoxels
 (c) StreamRF
@@
 Figure 3: Comparison of reconstruction quality on the most challenging flame salmon scene in DyNeRF dataset. PSNRs are computed under the resolution $1352\times 1014$.
-We compare our method with both online and offline baselines.
-As demonstrated in Table 1, our CD-NGP is 32×32$\times$32 × faster than the MLP-based continual representation INV [48], $\times4$ smaller than the voxel-based continual representation StreamRF [27] and significantly outperforms the two additional baselines in the continual learning setting.
-The undesirable results from HexPlane-PR and HexPlane-GR show the limitations of the regularization-based and replay-based methods in dynamic scenes: unlike the static scenes, the objects are not stationary across time and the models with limited complexity are not capable of reconstructing unlimited transient effects.
-Benefiting from the parameter isolation like StreamRF [27] and INV [48], our method achieves fast convergence and compact model size simultaneously.
-It leverages the hash tables to ensure fast convergence and reuses the features from differently-configured spatial hash tables to ensure high scalability.
-Compared with prevailing offline methods [28, 34, 7, 17, 45, 55], our method only requires caching $T_{chunk}=10$ frames into buffer by continual learning and thus reduces memory occupation by $85\%$.
-Our method also achieves $40\%$ less model size ($0.4$MB/frame) by leveraging the feature fusion of the base and the auxiliary branches with comparable quality.
+Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption
 Figure 4: The results of the proposed CD-NGP on DyNeRF dataset.
@@
 3(c), our model is much smaller without large quality loss.
-Compared with regularization-based and replay-based continual learning methods in Fig.
+Compared with regularization-based and replay-based online methods in Fig.
 3(d) and Fig.
@@
 Our method produces comparable quality to both offline and online state-of-the-art methods but with much less memory and storage cost.
-3.3.2 Meetroom dataset.
+3.3.2 Comparisons on the Meetroom dataset.
 Table 2 and Fig.
-5 shows the quantitative and qualitative comparisons on the meetroom dataset.
+5 show the quantitative and qualitative comparisons on the meetroom dataset.
 Fig.
-6 is the PSNR curve of the compared methods on the trimming scene.
-As shown in Table 2, HexPlane-GR and HexPlane-PR fail and produce unsatisfactory results on the meetroom dataset, we thus exclude them from Fig.
-6.
-Our method outperforms the implicit MLP-based online method [48] in terms of quality, speed, and model size significantly.
+6(a) is the PSNR curve of the compared methods on the trimming scene.
+We use the released configuration of StreamRF [27] for the meetroom dataset to obtain the results.
+As shown in Table 2, HexPlane-GR and HexPlane-PR produce significantly lower quality than the other methods on the meetroom dataset, we thus exclude them from Fig.
+6(a) for better visualization.
+The metrics in Table 2 and the curves in Fig.
+6(a) show the advantage of our method over both online and offline baselines.
+It outperforms the implicit MLP-based online method [48] regarding reconstruction quality, speed, and model size.
 Moreover, our method consumes only a fraction of storage than the explicit voxel-based online baseline [27] while maintaining comparable reconstruction quality as compared in Fig.
 5.
-When compared with offline baselines, our method produces comparable quality to MixVoxels [45] and outperforms HexPlane [7] in terms of quality and speed, and consumes much less memory as in Fig.
-5.
-We use the released configuration of StreamRF [27] for the meetroom dataset.
-The full model of [27] results in much worse metrics than those reported in their paper, so we only report the metrics of their pilot model.
-Figure 6: PSNR curves of the compared methods on the trimming scene in the Meetroom dataset. Solid and dotted lines denote offline and online methods respectively.
-Table 2: Results on the Meetroom dataset.
+When compared with offline baselines, our method produces comparable quality to MixVoxels [45] and outperforms HexPlane [7] in terms of quality and speed by much less memory cost.
+Figure 6: Combined PSNR curves for various scenes and methods.
+(a) The trimming scene in the Meetroom dataset.
+Solid and dotted lines denote offline and online methods respectively.
+(b) The first 1200 frames of seminar scene in our dataset.
+The solid lines and dotted lines denote offline and online methods respectively.
+(c) The first 2400 frames of seminar scene in our long video dataset.
+(d) the sear steak scene in the DyNeRF dataset.
+Solid and dotted lines denote symmetric and asymmetric setting, respectively.
+Table 2: Comparisons of the metrics on the Meetroom dataset.
 […]
-Mixvoxels $\checkmark$ 26.96 0.027 0.1034 15min 500MB
-HexPlane $\checkmark$ 24.71 0.043 0.2206 100min 200MB
-HexPlane-GR $\checkmark$ 19.05 0.099 0.3999 100min 318MB
-HexPlane-PR $\checkmark$ 18.63 0.112 0.3589 100min 318MB
-StreamRF $\checkmark$ 26.48 0.029 0.1699 75min 1710MB
-INV $\checkmark$ 23.87 0.041 0.1195 40hours 336MB
-CD-NGP (Ours) $\checkmark$ 26.01 0.039 0.1906 77min 113MB
-3.3.3 Spatial representations.
-Table 3: Ablation study on spatial representations. Different representations lead to subtle changes in quality metrics.
-[…]
-CD-NGP 21M 30.23 0.198 75min 113MB
-3D+4D 23M 29.34 0.243 82min 122MB
-CD-NGP-C 21M 30.52 0.174 80min 114MB
-CD-Plane 14M 29.16 0.215 77min 83MB
-CD-MERF 7.5M 29.93 0.201 109min 57MB
-CD-MERF-C 7.5M 30.22 0.174 115min 57MB
-Our proposed continual learning framework is compatible with multiple spatial backbones.
-We provide quantitative results on hybrid representation (denoted as CD-MERF) and plane representation like [17, 21] (denoted as CD-Plane) on the DyNeRF dataset in Table 3.
-Since the element-wise product for plane representation in [17] produces unsatisfactory results due to the hash collision, we remove the $P_{m}(x,y,z)$ in Eq.
+Mixvoxels $\checkmark$ 26.96 0.027 0.103 15min 500MB
+HexPlane $\checkmark$ 24.71 0.043 0.221 100min 200MB
+HexPlane-PR $\checkmark$ 18.63 0.112 0.359 100min 318MB
+HexPlane-GR $\checkmark$ 19.49 0.099 0.400 100min 318MB
+StreamRF $\checkmark$ 26.48 0.029 0.170 75min 1710MB
+INV $\checkmark$ 23.87 0.041 0.120 40hours 336MB
+CD-NGP (ours) $\checkmark$ 26.01 0.039 0.191 70min 113MB
+3.3.3 Spatial representations and scalability analysis.
+Our proposed continual learning framework is robust against changes in spatial backbones, including spatial hash encoder configurations and feature fusion methods.
+The hash encoder in Eq.
+5 could be changed into the hybrid representation (MERF) [38] and the plane representation like those in [17, 21].
+The MERF representation is defined as:
+$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}\leavevmode\nobreak\ P_{m}(x,y,z)+Q_{mxy}(x,y)+Q_{myz}(y,z)+Q_{mzx}(z,x),$ (12)
+where $Q_{mij}(i,j)$ denotes the bi-linearly interpolated features from the 4 nearby vertices in the orthogonally projected planes w.r.t.
+axes $(i,j)$ at the $m$-th resolution.
+The plane representation is defined as:
+$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}Q_{mxy}(x,y)+Q_{myz}(y,z)+Q_{mzx}(z,x).$ (13)
+For the simplicity of the notations, we denote the models adopting Eq. 5, Eq. 12, and Eq. 13 as CD-NGP, CD-MERF, and CD-Plane respectively.
+We compare the results of different spatial representations on the DyNeRF dataset and list the parameter numbers in Table 3.
+The alternatives of spatial representations include CD-MERF and CD-Plane.
+The element-wise product for plane representation in [17] produces unsatisfactory results in the experiments, we thus remove the $P_{m}(x,y,z)$ in Eq.
 12 and fuse the features $Q_{mij}(i,j)$ by concatenation in the implementation.
-We use $2^{P-2}$-length hash tables for the Plane representation in Table 3 to balance the quality and the model size.
-The structural configurations $(L,F)$ of the plane representation in Table 3 are set as $(6,4)$ for better results.
-The 3D+4D representation similar to [34] is yielded by concatenating the spatial $(x,y,z)$ features and spatial-temporal $(x,y,z,t)$ features encoded by a 4D extension of Eq.
-5.
-By comparing our CD-NGP with the naive 3D+4D representation, we show the advantage of our asymmetric encoding of spatial-temporal $(x,y,z,t)$ features explained in Eqs.
-3 and 7: representing the features of spatial coordinates $(x,y,z)$ and temporal coordinate $(t)$ separately avoids feature collisions and lead to better results.
-Table 3 also shows metrics of different feature fusion methods.
+As shown in Table 3, the hybrid representation CD-MERF achieves high compression and comparable model sizes to methods for static scenarios [31, 21] ($\sim$60MB) without significant quality loss.
+We also adopt the 3D+4D representation similar to that in [34] for comparison.
+This representation is implemented by concatenating the spatial $(x,y,z)$ features and spatial-temporal $(x,y,z,t)$ features encoded by a 4D extension of Eq.
+5, i.e.
+$\Gamma(x,t)=\bigoplus_{m=1}^{L}P_{m}(x,y,z,t)$.
+However, the asymmetric hash sizes (using $(P_{1},P_{2})=(19,14)$) in the 3D+4D representation results in serious feature collisions and deliver bad results (PSNR < 16 dB on 3 scenes).
+We thus set $(P_{1},P_{2})=(14,14)$ for the 3D+4D representation in Table 3 to keep the size of the model close to our CD-NGP for fair comparison.
+The metrics show the advantage of our separated encoding of spatial-temporal $(x,y,z,t)$ features explained in Eqs.
+3 and 6: representing the features of spatial coordinates $(x,y,z)$ and temporal coordinate $(t)$ separately avoids feature collisions and lead to better results.
+The results in Table 3 also show the flexibility in the feature fusion methods.
 We replace the element-wise sum in Eqs.
-7 and 6 with vector concatenation and denote them as CD-NGP-C and CD-MERF-C in Table 3.
-The hybrid representations (CD-MERF and CD-MERF-C) achieve comparable model sizes ($\sim$60MB) to methods for static scenarios [31, 21] without significant quality loss, reducing the average bandwidth cost to $0.2$MB/frame.
-Benefiting from larger feature dimensions, the vector concatenation leads to slightly better quality (especially when measured in LPIPS) by more training time, and the reasonable trade-off also shows the robustness of our method.
-Compared with the hash table designs in Table 7, the CD-NGP-C representation in Table 3 almost performs as well as the bulky $(18,18)$ configuration in Table 7 and reduces the model size by $80\%$.
-We list the per-scene results of the proposed continual learning method with different spatial representations on the DyNeRF dataset in Table 4.
-Table 4: Detailed results of the proposed CD-NGP and other hash-based methods on the DyNeRF dataset.
-Model Flame Salmon Cook Spinach
-[…]
-CD-NGP 26.37 0.04586 0.2886 30.76 0.02440 0.1793
-CD-NGP-C 26.65 0.04243 0.2680 31.03 0.02288 0.1593
-CD-Plane 25.48 0.06058 0.3173 29.41 0.03487 0.2010
-CD-MERF 26.15 0.04952 0.2921 30.35 0.02931 0.1858
-CD-MERF-C 26.56 0.04435 0.2546 30.69 0.02609 0.1583
-Model Flame Steak Sear Steak
-[…]
-CD-NGP 31.27 0.02132 0.1736 31.71 0.01917 0.1701
-CD-NGP-C 31.64 0.01823 0.1403 32.02 0.01698 0.1424
-CD-Plane 30.26 0.02949 0.1804 30.56 0.02723 0.1811
-CD-MERF 31.01 0.02508 0.1719 31.44 0.02181 0.1698
-CD-MERF-C 31.13 0.02424 0.1506 31.66 0.01989 0.1413
-Model Cut Roasted Beef Average
-[…]
-CD-NGP 31.03 0.02366 0.1801 30.228 0.026882 0.19834
-CD-NGP-C 31.26 0.02205 0.1578 30.520 0.024514 0.17356
-CD-Plane 30.11 0.03302 0.1951 29.164 0.037038 0.21498
-CD-MERF 30.71 0.02765 0.1856 29.932 0.030674 0.20104
-CD-MERF-C 31.04 0.02450 0.1632 30.216 0.027814 0.17360
-The detailed parameter numbers and bandwidth consumption are listed in Table 5.
-3D Enc.
-and 2D Enc.
-denote the spatial encodings along voxels and planes respectively.
-The tuples $(M_{1},M_{2})$ in the cells denote the parameter numbers of the base encoder and each of the auxiliary encoders.
-Following [27, 43, 1, 48] we also report the minimum per-frame bandwidth consumption $\mathbf{B_{min}}$, which only depends on the parameter numbers of the auxiliary branches and is thus significantly lower than the average bandwidth consumption $\mathbf{B_{avg}}$.
-Since the base encoder and the density grid could be shared among different branches and transmitted only once in the beginning, the minimum bandwidth consumption is reduced to $0.065\sim 0.16$ MB per frame, which is 100×100$\times$100 × less than the voxel-based StreamRF [27] and 7×7$\times$7 × less than the pure MLP representation INV [48].
-Table 5: The parameter numbers of each module of the 30-branch representations in Table 3.
+6 and 12 with vector concatenation and denote them as CD-NGP-C and CD-MERF-C in Table 3.
+Benefiting from larger feature dimensions, these methods lead to slightly better quality (especially when measured in LPIPS) at the cost of more training time, and the reasonable trade-off also shows the robustness of our method.
+Table 3: Quality metrics, parameter numbers, and bandwidth costs of different spatial representations.
+The parameter numbers of each module of the 30-branch representations are also shown.
+#$\#3DEnc.$ and #$\#2DEnc.$ denote the parameter numbers of the 3D encoders and 2D encoders in the base branch and one auxiliary branch accordingly.
 $\mathbf{B_{min}}$ and $\mathbf{B_{avg}}$ are minimum and average bandwidth costs in MB/frame.
 ##\## denotes the number of parameters per branch (10 frames).
-Model ##\##3D Enc. ##\##2D Enc. ##\##MLP ##\##Online ##\##Total $\mathbf{B_{min}}$ $\mathbf{B_{avg}}$
-CD-NGP $(9.9M,0.37M)$ $0.042M$ $0.39M$ $21M$ 0.16 0.38
-CD-NGP-C $(9.9M,0.37M)$ $0.046M$ $0.40M$ $21M$ 0.16 0.38
-CD-Plane - $(1.9M,0.082M)$ $0.048M$ $0.28M$ $14M$ 0.11 0.28
-CD-MERF $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.042M$ $0.16M$ $7.5M$ 0.061 0.19
-CD-MERF-C $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.052M$ $0.17M$ $7.5M$ 0.065 0.19
+[…]
+CD-NGP 21M 30.23 0.198 75min 113MB $(9.9M,0.37M)$ $0.39M$ 0.16 0.38
+3D+4D 23M 29.31 0.242 82min 122MB $(0.73M,0.73M)$ $0.76M$ 0.31 0.41
+CD-NGP-C 21M 30.52 0.174 80min 114MB $(9.9M,0.37M)$ $0.40M$ 0.16 0.38
+CD-Plane 14M 29.16 0.215 77min 83MB - $(1.9M,0.082M)$ $0.28M$ 0.11 0.28
+CD-MERF 7.5M 29.93 0.201 109min 57MB $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.16M$ 0.061 0.19
+CD-MERF-C 7.5M 30.22 0.174 115min 57MB $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.17M$ 0.065 0.19
+The parameter numbers and bandwidth consumption in Table 3 also prove the superior scalability of our method, where 3D Enc.
+and 2D Enc.
+denote the spatial encodings according to voxel and plane indices respectively.
+Following MERF [38] for static scenes, we use a smaller hash table for the hybrid representation in Eq.
+12 than pure voxel hash grid [31] representation in Table 3 to ensure the compactness and scalability of our method.
+Denoting the lengths of hash tables along pure voxels in a branch of CD-NGP as $2^{P}$, the lengths of hash tables in the CD-MERF are set as $2^{P-3}$ and $2^{P-4}$ for voxel and plane encoders respectively.
+The number of the parameters of the MERF representation is reduced to $1-2^{-3}-3\times 2^{-4}=31.25\%$ of the voxel hash grids.
+Since the hash tables comprise most of the parameters, the model size is reduced from $21M$ to $7.5M$ as listed in Table 3.
+Likewise, we use $2^{P-2}$-length hash tables for the Plane representation in Table 3 to balance the quality and the model size.
+The structural configurations $(L,F)$ of the plane representation in Table 3 are set as $(6,4)$ , and the resolution of the hash encoders of the CD-Plane are set within $(64,2048)$ for better results.
+Following [27, 43, 1, 48] we also report the minimum per-frame bandwidth cost $\mathbf{B_{min}}$, which only depends on the parameter numbers of the auxiliary branches and is thus significantly lower than the average bandwidth consumption $\mathbf{B_{avg}}$.
+Since the base encoder and the density grid could be shared among different branches and transmitted only once in the beginning, the minimum bandwidth consumption is reduced to $0.065\sim 0.16$ MB per frame, which is 100×100$\times$100 × less than the voxel-based StreamRF [27] and 7×7$\times$7 × less than the pure MLP representation INV [48].
 3.3.4 Long sequence dataset and scalability analysis.
-We also provide quantitative comparisons on our proposed long multi-view video dataset in Table 6 and Fig.
-7 to further show the superior scalability of our method.
+As introduced in Section 2.3, our method shows superior scalability: the total size of the model is $\mathcal{O}(N_{chunk}\cdot 2^{P_{2}}LF+2^{P_{1}}LF)$, where $N_{chunk}$ is the number of chunks in the sequence.
+To prove the scalability of our method, we provide quantitative comparisons on our proposed long multi-view video dataset in Table 4 and Fig.
+6(b) to further show the superior scalability of our method.
 We split the sequence into 300-frame chunks for the compared methods since the offline baseline [7] is designed for 300-frame sequences in the DyNeRF dataset.
@@
 Compared with the results in Table 1, the performance gap in storage consumption is even larger in long video sequences.
-This is because the whole model shares one base branch and the following auxiliary branches occupy the majority of the parameters.
-We thus set the hash table size in the auxiliary branches to $2^{12}LF$ and leverage the re-use of the features from the base branch to reduce the model size to $\mathcal{O}(N_{chunk}\cdot 2^{12}LF)$.
-.
+This is because the majority of the parameters in our method are used for the auxiliary branches, and the base branch only takes up a small fraction of the parameters.
+Consequently, applying smaller hash tables in the auxiliary branches and leverage the re-use of the features in the base branch leads to a significant reduction in the model size.
 The PSNR curves in Fig.
-7 further demonstrate the stability of our method.
-The compared online methods suffer from catastrophic forgetting: the PSNRs at the beginning of the chunks are much lower than that in the end.
-The problem is properly eliminated by our method through parameter isolation, and our method continuously produce high-quality results for the whole sequence.
+6(b) further demonstrate the stability of our method.
+The compared regularization-based and replay-based methods suffer from catastrophic forgetting: the PSNRs at the beginning of the chunks are much lower than those in the end.
+The problem is properly eliminated by our method through parameter isolation, and our method continuously produces high-quality results for the whole sequence.
 We also provide comparisons between the rendered results in Fig.
+7.
+The regularization method displayed in Fig.
+7(a) reconstructs the TV screen in the background erroneously.
+The replay method displayed in Fig.
+7(b) fails to recover the details.
+Ours in Fig.
+7(f) produces comparable quality to the offline baseline in Fig.
+7(c).
+To further demonstrate the scalability of our method, we conduct experiments on the first 2400 frames of the seminar scene in our long video dataset.
+We further show the PSNR curves in Fig.
+6(c) and compare the model sizes in Fig.
 8.
-The regularization method displayed in Fig.
-8(a) reconstructs the TV screen in the background erroneously.
-The replay method displayed in Fig.
-8(b) fails to recover the details.
-Ours in Fig.
-8(f) produces comparable quality to the offline baseline in Fig.
-8(c).
-To further demonstrate the scalability of our method, we conduct experiments on the first 2400 frames of the seminar scene in our long video dataset, show the PSNR curves in Fig.
-10 and compare the model sizes in Fig.
-11.
-For simplicity of the evaluation, we set the chunk size as 10 for the CD-NGP and CD-MERF method.
-Our method produces high-quality results, and consumes only a fraction of storage space compared with the other online baselines.
-Figure 7: PSNR curves of the first 1200 frames of seminar scene in our long video dataset. The solid lines and dotted lines denote offline and online methods respectively.
+Our method produces high-quality results while consuming only a fraction of storage space compared with the other online baselines.
 3.4 Ablation studies
@@
 Figure 7: Comparison of reconstruction quality on the 19th frame of the seminar scene in our long video dataset.
-Figure 9: PSNR curves of asymmetric (solid line, ours) and symmetric (dotted line) hash table designs on the sear steak scene.
-Figure 10: PSNR curves of the first 2400 frames of seminar scene in our long video dataset.
-Figure 11: Comparison of the model sizes of our method and other online baselines. Our method show superior scalability over the compared baselines.
+Figure 8: Comparison of the model sizes of our method and other online baselines. Our method shows superior scalability over others.
 3.4.1 Lengths of hash tables.
 Benefiting from the multi-branch hash table design, our model is highly efficient and configurable.
-We change the parameter numbers of the initial branch and the auxiliary branches and obtain the results on the DyNeRF dataset in Table 7.
-$(P_{1},P_{2})$ are the log2 values of the hash table lengths of the base branch and the other branches.
-Since the whole model contains hash encoders, MLP regressors, and the occupancy grids.
-The sizes of the models are thus larger than the simple combinations of hash table and MLP regressors.
-We use the default configuration $(19,14)$ in the following experiments.
-Although larger hash tables alleviate hash collisions, the configuration $(19,14)$ in CD-NGP produces better results than the naive $(16,16)$ configuration even if the storage consumption is $40\%$ lower.
+We change the lengths of the hash tables of the initial branch and the auxiliary branches and obtain the results on the DyNeRF dataset in Table 5.
+$(P_{1},P_{2})$ are the log2 values of the hash table lengths of the base branch and the auxiliary branches.
+The whole model contains hash encoders, MLP regressors, and the occupancy gridd, the sizes of the models are thus larger than the value deduced from the combinations of the hash table and MLP regressors.
+Although larger hash tables lead to less hash collisions, the default configuration $(19,14)$ in CD-NGP produces better results than the naive $(16,16)$ configuration even if the storage consumption is $40\%$ lower.
 Moreover, the $(19,12)$ configuration of CD-NGP significantly outperforms the naive $(14,14)$ configuration at a similar storage cost.
-To further demonstrate the scalability of our method, we provide the PSNR curves of different hash table designs on the sear steak scene in Fig.
-9.
+To further demonstrate the superiority of the asymmetric hash table design and prove the scalability of our method, we provide the PSNR curves of different hash table designs on the sear steak scene in Fig.
+6(d).
 Equipped with a larger hash table in the base branch, the asymmetric setting $(19,14)$ and $(19,12)$ outperforms others in the beginning and produces comparable metrics to the naive $(16,16)$, $(14,14)$ configuration respectively in the subsequent chunks.
-The advantage comes from our asymmetric hash table designs: the base branch with a large hash table recovers more environment patterns, and the auxiliary branches could thus use fewer parameters to cope with recent changes only.
+The $(18,18)$ in Table 5 configuration delivers the best results (30.55 dB in PSNR) on the DyNeRF dataset, but at the cost of an exceptionally large model size (638MB).
+However, the CD-NGP-C configuration with $(P_{1},P_{2})=(19,14)$ listed in Table 3 achieves a comparable PSNR (30.52 dB) with a much smaller model size (114MB, $80\%$ compression).
+The advantage comes from our re-use of the spatial features encoded by different branches: the base branch with a large hash table recovers more environment details, and the auxiliary branches could thus use fewer parameters to cope with recent changes only.
+The combination of the base encode and the auxiliary encoders lead to a more efficient and scalable model.
 Table 4: Quantitative comparisons between our CD-NGP and baselines on the long video sequences in our dataset.
@@
 […]
-HexPlane $\checkmark$ 28.32 0.117 800MB 8 hours 74GB
-HexPlane-PR $\checkmark$ 20.90 0.286 1272MB 8 hours 12GB
-HexPlane-GR $\checkmark$ 24.82 0.282 1272MB 8 hours 12GB
-StreamRF $\checkmark$ 24.99 0.266 13GB 4 hours 12GB
-INV $\checkmark$ 27.07 0.086 1344MB 160 hours 12GB
-CD-NGP (Ours) $\checkmark$ 27.15 0.190 201MB 10 hours 12GB
+MixVoxels $\checkmark$ 26.94 0.022 0.118 2000MB 80 min 80GB
+HexPlane $\checkmark$ 27.74 0.022 0.117 800MB 8 hours 74GB
+HexPlane-PR $\checkmark$ 20.90 0.095 0.286 1272MB 8 hours 12GB
+HexPlane-GR $\checkmark$ 24.82 0.052 0.282 1272MB 8 hours 12GB
+StreamRF $\checkmark$ 24.72 0.036 0.273 13GB 4 hours 12GB
+INV $\checkmark$ 27.12 0.025 0.086 1344MB 160 hours 12GB
+CD-NGP (ours) $\checkmark$ 27.15 0.039 0.188 201MB 10 hours 12GB
 Table 5: Ablation study on hash table sizes. denotes the default setting.
 […]
-$(18,18)$ 151M 30.55 638MB
-$(16,16)$ 41M 30.08 191MB
-$(14,14)$ 12M 29.32 75MB
-$(19,14)^{{\dagger}}$ 21M 30.23 113MB
-$(16,14)$ 13M 29.78 79MB
-$(19,12)$ 14M 30.09 82MB
+$(18,18)$ 151M 30.55 0.024 0.161 638MB
+$(16,16)$ 41M 30.08 0.027 0.194 191MB
+$(14,14)$ 12M 29.32 0.035 0.248 75MB
+$(19,14)^{{\dagger}}$ 21M 30.23 0.027 0.198 113MB
+$(16,14)$ 13M 29.78 0.030 0.228 79MB
+$(19,12)$ 14M 30.09 0.028 0.209 82MB
 3.4.2 Temporal representations.
@@
 3.4.3 Field composition methods.
-We also conduct ablation studies on the field composition methods and show the results in Tables 9 and 12. The field composition process is defined as:
+We also conduct ablation studies on the field composition methods and show the results in Tables 7 and 9. The field composition process is defined as:
 $(\sigma,\mathbf{c})=(\sigma_{i}+\sigma_{a},\mathbf{c}_{i}+\mathbf{c}_{a}).$ (14)
@@
 Furthermore, as compared in Fig.
-12, the compared composition methods in Figs.
-12(c) and 12(d) suffer from salt and pepper noise because of the explicit composition, while our CD-NGP circumvents the problem by fusing features in the latent space and produces better results.
+9, the compared composition methods in Figs.
+9(c) and 9(d) suffer from salt and pepper noise because of the explicit composition, while our CD-NGP circumvents the problem by fusing features in the latent space and produces better results.
 (a) Ground truth
@@
 Figure 9: Comparison of different field composition methods.
-3.4.4 Regularization weights.
-We also provide ablation studies on the regularization weights of auxiliary spatial encodings.
-We apply regularizations on $\Psi_{k}(x)$ with different $\lambda_{r}$ and show the results in Table 10.
-The regularization is of slight improvement, and our method is robust within a wide range of $\lambda_{r}$.
-Table 10: Ablation study on regularization weight $\lambda_{r}$.
-[…]
-$10^{-5}$ 30.20 0.0269 0.198
-$10^{-4}$ 30.18 0.0268 0.196
-$10^{-3}$ 30.23 0.0269 0.198
-$10^{-2}$ 30.09 0.0279 0.197
-1 29.66 0.0313 0.205
-3.4.5 Chunk size.
-We also conduct ablation studies on the chunk sizes and show the results in Tables 11 and 12.
-We only report the results on the cook spinach scene in the DyNeRF dataset for brevity.
-Our method is robust within a wide range of chunk sizes.
-As in Table 11, the smaller chunk size leads to better results when controlling the iteration numbers of each chunk.
-When the total iterations of the whole 300-frame sequence are held constant, the larger chunk size leads to better results as in Table 12.
-We select $T_{chunk}=10$ as the default configuration.
-The selection of the chunk size is a trade-off between the quality, robustness, training time, and host memory occupation and should depend on the data.
-Table 11: Ablation study on chunk sizes. The iteration numbers of each chunk are held constant under different $T_{chunk}$ settings.
+Table 8: Ablation study on the chunk sizes. The iteration numbers of each chunk are held constant under different $T_{chunk}$ settings.
 […]
@@
 5 30.99 0.02268 0.1674 162MB
-Table 12: Ablation study on chunk sizes. The total iterations of the whole 300-frame sequence are held constant under different $T_{chunk}$ settings.
+Table 9: Ablation study on the chunk sizes. The total iterations of the whole 300-frame video are held constant under different $T_{chunk}$ settings.
 […]
@@
 5 30.39 0.02602 0.2075 162MB
+3.4.4 Chunk size.
+We also conduct ablation studies on the chunk sizes and show the results in Tables 8 and 9.
+We only report the results on the cook spinach scene in the DyNeRF dataset for brevity.
+As in Table 8, the smaller chunk size leads to better results when controlling the iteration numbers of each chunk.
+When the total iterations of the whole 300-frame sequence are held constant, the larger chunk size leads to better results as in Table 9.
+The selection of the chunk size is a trade-off between the quality, robustness, training time, and host memory occupation and should depend on the data, and our method is robust within a wide range of chunk sizes.
+We select $T_{chunk}=10$ as the default configuration.
 4 Conclusion and Limitation
-We propose CD-NGP, a continual learning framework for multi-view dynamic scenes based on parameter isolation.
-Equipped with the multi-branch hash encoding structure, it shows high efficiency and scalability and is capable of reconstructing long video sequences with low host memory occupation.
-Compared with offline methods, it consumes $40\%$ less bandwidth and $85\%$ less memory than recent explicit baselines and achieves comparable quality while maintaining the speed advantage over the pure-MLP-based methods.
-Compared with online baselines, it significantly outperforms replay and regularization-based continual learning methods and reaches a new balance among reconstruction quality, training speed, bandwidth occupation, and memory cost.
-However, like most of the NeRF-based methods, it still requires querying the field function and is not appropriate for achieving real-time rendering.
-Further efforts could be made to apply baking methods or Gaussian point clouds to continual learning to achieve real-time rendering in long video sequences for wide applications.
+We propose CD-NGP, a continual learning framework for multi-view dynamic scenes based on parameter isolation, along with a self-collected dataset for novel view synthesis under multi-view long-view scenarios.
+Our proposed method splits the input multi-view videos into different chunks, trains different model branches, and performs reconstruction accordingly.
+Each branch encodes spatial and temporal features separately using the multi-resolution hash encoding structure.
+It leverages the redundancy of the dynamic scenes and reuses the spatial features in the first branch to achieve fast convergence, high fidelity, and superior scalability.
+On the prevailing DyNeRF dataset, it consumes $40\%$ less bandwidth and $85\%$ less memory than recent explicit offline baselines, achieves comparable quality, and maintains the speed advantage over pure-MLP-based offline methods.
+Compared with online baselines, it delivers state-of-the-art quality and achieves a new balance among reconstruction quality, training speed, bandwidth occupation, and memory cost.
+On the long video dataset, our method delivers high quality and achieves fast convergence with only a fraction of the storage cost.
+However, like most NeRF-based methods, it still requires querying the field function and is not suitable for real-time rendering.
+Further efforts could be made to apply baking methods or Gaussian point clouds to continual learning, enabling real-time rendering in long video sequences for wide applications.
+Appendix A Per-scene results
+We list the per-scene results of the proposed continual learning method with different spatial representations on the DyNeRF dataset in Table 10, and the per-scene results of the methods in Table 4 on our self-collected long video dataset in Table 12.
+Table 10: Detailed results of the proposed CD-NGP and other hash-based methods on the DyNeRF dataset.
+Model Flame Salmon Cook Spinach Cut Roasted Beef
+[…]
+CD-NGP 26.37 0.046 0.289 30.76 0.024 0.179 31.03 0.024 0.180
+3D+4D 25.72 0.059 0.355 29.73 0.032 0.220 30.06 0.031 0.224
+CD-NGP-C 26.65 0.042 0.268 31.03 0.022 0.159 31.26 0.022 0.158
+CD-Plane 25.48 0.060 0.317 29.41 0.034 0.201 30.11 0.033 0.195
+CD-MERF 26.15 0.049 0.292 30.35 0.029 0.186 30.71 0.028 0.186
+CD-MERF-C 26.56 0.044 0.255 30.69 0.026 0.158 31.04 0.025 0.163
+Model Flame Steak Sear Steak Average
+[…]
+CD-NGP 31.27 0.021 0.174 31.71 0.019 0.170 30.23 0.027 0.198
+3D+4D 30.33 0.028 0.205 30.73 0.026 0.207 29.31 0.035 0.242
+CD-NGP-C 31.64 0.018 0.140 32.02 0.017 0.142 30.52 0.025 0.174
+CD-Plane 30.26 0.029 0.180 30.56 0.027 0.181 29.16 0.037 0.215
+CD-MERF 31.01 0.025 0.172 31.44 0.022 0.170 29.93 0.031 0.201
+CD-MERF-C 31.13 0.024 0.151 31.66 0.020 0.141 30.22 0.028 0.174
+Table 11: Per-scene results of the compared methods on the meetroom dataset in Table 2.
+Model Trimming Discussion
+[…]
+MixVoxels 27.21 0.026 0.093 27.42 0.026 0.110
+HexPlane 22.13 0.062 0.263 26.28 0.032 0.212
+HexPlane-PR 17.30 0.140 0.404 18.53 0.111 0.355
+HexPlane-GR 12.80 0.184 0.554 23.31 0.061 0.330
+StreamRF 26.72 0.029 0.170 26.65 0.028 0.169
+INV 27.08 0.026 0.089 24.57 0.036 0.110
+CD-NGP 27.58 0.040 0.164 26.01 0.040 0.214
+Model VRheadset Average
+[…]
+MixVoxels 26.25 0.029 0.107 26.96 0.027 0.103
+HexPlane 25.72 0.035 0.187 24.71 0.043 0.221
+HexPlane-PR 20.05 0.085 0.317 18.63 0.112 0.359
+HexPlane-GR 22.37 0.054 0.316 19.49 0.099 0.400
+StreamRF 26.06 0.029 0.171 26.48 0.029 0.170
+INV 19.95 0.061 0.160 23.87 0.041 0.120
+CD-NGP 24.43 0.037 0.193 26.01 0.039 0.191
+Table 12: Detailed results of the proposed CD-NGP and other hash-based methods on our long video dataset.
+Model Demo Discussion Kungfu
+[…]
+MixVoxels 27.50 0.021 0.103 28.01 0.016 0.097 24.43 0.033 0.140
+HexPlane 28.92 0.016 0.095 27.17 0.024 0.136 25.14 0.032 0.136
+HexPlane-PR 21.10 0.090 0.268 20.98 0.094 0.315 18.85 0.112 0.309
+HexPlane-GR 24.85 0.041 0.243 24.03 0.050 0.296 22.15 0.072 0.336
+StreamRF 24.86 0.032 0.233 24.41 0.029 0.258 23.75 0.043 0.301
+INV 26.55 0.025 0.093 28.51 0.017 0.054 24.57 0.037 0.120
+CD-NGP 26.66 0.030 0.161 29.66 0.057 0.137 23.66 0.055 0.257
+Model Reading Seminar Average
+[…]
+MixVoxels 25.99 0.021 0.144 28.78 0.018 0.108 26.94 0.022 0.118
+HexPlane 27.88 0.018 0.099 29.59 0.018 0.120 27.74 0.022 0.117
+HexPlane-PR 21.85 0.087 0.287 21.78 0.092 0.280 20.91 0.095 0.292
+HexPlane-GR 26.52 0.040 0.266 25.77 0.053 0.306 24.66 0.051 0.289
+StreamRF 24.06 0.040 0.325 26.53 0.037 0.250 24.72 0.036 0.273
+INV 26.39 0.026 0.106 29.60 0.018 0.054 27.12 0.025 0.085
+CD-NGP 26.70 0.031 0.212 28.97 0.024 0.169 27.13 0.039 0.201

### v6

@@
 Abstract
-Current methods for novel view synthesis (NVS) in dynamic scenes encounter significant challenges in managing memory consumption, model complexity, training efficiency, and rendering fidelity.
-Existing offline techniques, while delivering high-quality results, face challenges from substantial memory demands and limited scalability.
-Conversely, online methods struggle to balance rapid convergence with model compactness.
-To address these issues, we propose continual dynamic neural graphics primitives (CD-NGP).
-Our approach leverages a continual learning framework to reduce memory overhead, and it also integrates features from distinct temporal and spatial hash encodings for high rendering quality.
-Meanwhile, our method employs parameter reuse to achieve high scalability.
-Additionally, we introduce a novel dataset featuring multi-view, exceptionally long video sequences with substantial rigid and non-rigid motion, which is seldom possessed by existing datasets.
-We evaluate the reconstruction quality, speed and scalability of our method on both the established public datasets and our exceptionally long video dataset.
-Notably, our method achieves an $85\%$ reduction in training memory consumption (less than 14GB) compared to offline techniques and significantly lowers streaming bandwidth requirements (less than 0.4MB/frame) relative to other online alternatives.
-The experimental results on our long video sequences dataset show the superior scalability and reconstruction quality compared to existing state-of-the-art approaches.
+Novel view synthesis (NVS) in dynamic scenes faces persistent challenges in memory consumption, model complexity, training efficiency, and rendering quality.
+Offline methods offer high fidelity but suffer from high memory usage and limited scalability, while online approaches often trade quality for speed and compactness.
+We propose Continual Dynamic Neural Graphics Primitives (CD-NGP), a continual learning framework that reduces memory overhead and enhances scalability through parameter reuse.
+To avoid feature interference in dynamic scenes and improve rendering quality, our method combines spatial and temporal hash encodings, which compactly represent scene structures and motion patterns.
+We also introduce a new dataset comprising multi-view, long-duration ($>1200$ frames) videos with both rigid and non-rigid motion, which is not found in existing benchmarks.
+CD-NGP is evaluated on public datasets and our long video dataset, demonstrating superior scalability and reconstruction quality.
+It significantly reduces training memory usage (to $<$14GB) and requires only 0.4MB/frame in streaming bandwidth on DyNeRF—substantially lower than most online baselines.
 keywords:
@@
 […]
-The outline of our continual learning pipeline for dynamic scenes.
-Our method segments input videos into multiple chunks and encodes the chunks by separate spatial and time hash encodings.
-Subsequently, we train the auxiliary branches incrementally on each chunk and leverage the re-use of parameters in the base branch to ensure high scalability.
-The spatial hash features could be encoded according to voxel grids, orthogonally projected plane grids or a combination of both.
+Overview of our continual learning pipeline for dynamic scenes.
+Our method segments multi-view input videos into multiple chunks and subsequently reconstructs them using individual model branches.
+Each branch contains individual spatial hash encodings for $(x,y,z)$, temporal hash encodings for $t$, and MLP regressors for downstream rendering and optimization.
+Benefiting from the reuse of the larger hash table in the first branch (base branch) and the fusion of features from the subsequent branches (auxiliary branches), our method achieves high scalability and fast convergence.
+Leveraging the flexible hash encoding structure, it also demonstrates strong adaptability to different encoding types, including 3D voxel grids, 2D plane grids, or hybrid variants.
 Introduction
@@
 For example, the methods in [39, 28, 37, 36, 15, 29, 49] propose to use MLP-based representation and achieve high-quality results.
-There are also methods that speed up the optimization by using the tensor factorization ([40, 17, 7]) and voxel representations [16, 45].
+Some methods speed up the optimization by using the tensor factorization ([45, 17, 7]) and voxel representations [16, 50].
 The more recent continual-learning-based methods [27, 53] (also mentioned as online methods) can train models frame by frame, and streamable representations [48] enable the model to be loaded on the fly at inference.
-However, the offline methods require huge host memory e.g., taking more than $100GB$ when reconstructing multi-view dynamic videos of 300 frames, and the online methods are weak in balancing the convergence speed and model compactness.
-These limitations in scalability makes them impractical for the wide application on long videos.
+On the one hand, most offline approaches require loading all video frames of the training set into memory, imposing substantial demands on hardware resources (e.g., approximately 100 GB for the DyNeRF dataset [28] with 20 views $\times$ 300 frames).
+Lazy loading reduces peak memory usage [61], but still requires prohibitive disk space for full video frame caching.
+Moreover, the overall resource consumption scales linearly with the number of frames ($O(N_{frames})$), which makes it unsuitable for lengthy video sequences.
+On the other hand, recent online methods [27, 53] fail to simultaneously achieve fast convergence and compact model size, which constrains their applicability in long-term dynamic scenarios.
+These limitations motivate us to design a more scalable and memory-efficient representation for modeling dynamic scenes over extended durations.
 In this paper, we propose continual dynamic neural graphics primitives (CD-NGP), a fast and scalable representation for reconstructing dynamic scenes, which enables dynamic reconstruction in long videos.
@@
 Each branch of our model includes MLP regressors, spatial hash encoding for spatial coordinates $(x,y,z)$, and temporal hash encoding for the timestamp $t$.
-Secondly, the spatial and temporal features are encoded separately to avoid feature confusion in different axes.
+Secondly, the spatial and temporal features are encoded separately to avoid feature interference in different axes.
 For each model branch, we sample 3D points $(x,y,z)$ according to frame pixels and cached occupancy grids like [34], and encode the points into spatial features using spatial hash grids.
@@
 Thirdly, we leverage the redundancy in the dynamic scenes and fuse the spatial features of different branches to yield high scalability.
-We set a larger hash table size ($2^{19}$) in the initial branch (base branch), and set a smaller size ($2^{14}$) for the other branches (auxiliary branches).
-This asymetric design facilitates reusing the features trained in the base branch, thereby reducing storage consumption and boosting rendering quality.
-Figure 1: Comparison between different performance indexes of our method and recent representative baselines.
-The perceptual consistency is measured by $1-$ LPIPS by AlexNet [57, 24], and the speed is the reciprocal of the training time.
-The compactness and RAM-Efficiency are mapped by $-log$ function from the model size and memory occupation respectively.
-We validate the effectiveness of our method through comprehensive experimental results on both the widely used public datasets and a self-collected dataset.
+We set a larger hash table size ($2^{19}$) in the initial branch (the base branch), and set a smaller size ($2^{14}$) for the other branches (the auxiliary branches).
+This asymmetric design facilitates reusing the features trained in the base branch, thereby reducing storage consumption and boosting rendering quality.
+Figure 1: Comparison of various performance indices between our method and recent representative baselines.
+Perceptual consistency is measured by $1-$ LPIPS using AlexNet [63, 24], and convergence speed is defined as the reciprocal of training time.
+Compactness and RAM efficiency are derived by applying a $-\log$ transformation to model size and memory usage, respectively.
+We validate the effectiveness of our method through comprehensive experimental results on both widely used public datasets and a self-collected dataset.
 As illustrated in Fig.
 1, our method reaches a new balance between rendering quality, model size, and training time while remaining memory-friendly.
-Compared with prevailing offline baselines, our method consumes $85\%$ less host memory at the training phase, $40\%$ less model size at the inference phase, and delivers comparable results.
-Compared with online baselines, our method achieves compact model size, high rendering quality, and fast convergence simultaneously.
+Compared with most prevailing offline baselines, our method requires significantly less host memory during training, maintains a smaller model size at inference, and achieves comparable rendering quality by quantitative metrics.
+In contrast to online baselines, it simultaneously attains compact model size, high rendering fidelity, and fast convergence.
 The contributions of this paper are three-fold as follows:
-  • We propose a novel method combining separate spatial and temporal hash encodings, achieving high scalability and rendering quality in dynamic scenes.
+  • We propose a novel method that combines separate spatial and temporal hash encodings to efficiently encode both scene geometry and motion patterns.
+This design improves scalability and rendering quality in dynamic scenes.
   • We propose a fast and highly scalable continual learning framework for dynamic scenes.
@@
 Our method also outperforms online alternatives and achieves comparable results to offline methods.
-  • We provide a dataset of multi-view exceptionally long video sequences ($>3$ minutes) with large non-rigid motion, which is seldom possessed by existing datasets.
+  • We provide a dataset of multi-view exceptionally long video sequences ($>3$ minutes) with large non-rigid motion, which is rarely found in existing datasets.
 1 Related Work
@@
 MixVoxels [50] leverages the standard deviation across the time of all pixels to estimate a variation field to decompose the dynamic scene into static and dynamic voxels.
-Equipped with different representations for static and dynamic objects, MixVoxels achieves high-fidelity results and super-fast convergence (15 minutes) on DyNeRF dataset.
+Equipped with different representations for static and dynamic objects, MixVoxels achieves high-fidelity results and super-fast convergence (15 minutes) on the DyNeRF dataset.
 Song et al.
 [48] train Instant-NGP [34] and TensoRF [9] offline and stream the models along the feature channels for dynamic scenes.
-Concurrent works are leveraging 4D Gaussian point clouds [55, 49] or 3D Gaussians with deformation networks [54] achieving real-time rendering and high fidelity in dynamic scenes.
+4K4D [58] uses foreground masks to separate the dynamic objects from the static background, and leverages a point cloud-based method to achieve real-time rendering.
+Concurrent works are leveraging 4D Gaussian point clouds [61, 54] or 3D Gaussians with deformation networks [60] achieving real-time rendering and high fidelity simultaneously.
 1.3 Continual learning of NeRF
-Continual learning (online learning) aims to adapt to new tasks while defying catastrophic forgetting (resisting performance downgrade on the previous tasks) [12].
+Continual learning (online learning) aims to adapt to new tasks while mitigating catastrophic forgetting (resisting performance downgrade on the previous tasks) [12].
 There are mainly 3 approaches to achieving continual learning: replay, regularization, and parameter isolation.
@@
 The regularization-based methods hold the parameters close to the solution of previous tasks [14].
-The parameter-isolation methods use different parameters to cope with different tasks.
-Recent efforts based on generative replay [6, 35, 11] enable continual learning of NeRF in static scenes and achieve high fidelity.
-StreamRF [27] applies continual NeRF to dynamic scenes by parameter isolation.
-Their explicit voxel representation enables fast reconstruction (75 minutes for 300 frames) and could be streamed frame by frame.
+The parameter-isolation methods use disjoint parameters to cope with different tasks.
+Depending on the application scenarios, the continual learning methods can be divided into two categories: static and dynamic.
+On the one hand, several recent works leverage MLP [11] or hash grid representations [6, 38] and replay to enable continual learning of NeRF in static scenes.
+On the other hand, for the dynamic scenes, the continual learning is mainly based on parameter isolation: StreamRF [27] uses parameter isolation of explicit voxel grids to extend continual NeRF to dynamic scenes.
+Their representation enables fast reconstruction (75 minutes for 300 frames) and can be streamed frame by frame.
 INV [53] uses a shared color MLP and trains different structure MLPs frame by frame to form a continual NeRF representation.
-It achieves high reconstruction quality in dynamic scenes.
+It delivers high rendering quality in dynamic scenes.
 2 Method
+Previous online methods for dynamic scenes [27, 53] do not consider the scalability of the model: their model size scales almost linearly with the number of frames, which is of limited use in long video sequences.
+Our method combines hash representation and parameter isolation for continual learning and achieves high scalability: it not only supports continually reconstructing the dynamic scene but also scales with a significantly lower model size complexity.
 2.1 Problem definition and preliminaries
@@
 Thus, the replay-based method suffers from catastrophic forgetting.
-The design of our model follow the parameter isolation approach: Let the time stamps be divided into $N_{chunk}$ chunks, the model consists of $N_{chunk}$ branches to represent different chunks of the dynamic scene accordingly.
+The design of our model follows the parameter isolation approach: Let the time stamps be divided into $N_{chunk}$ chunks, and the model consists of $N_{chunk}$ branches to represent different chunks of the dynamic scene accordingly.
 Following previous methods [39, 28, 17, 7], we consider the dynamic scene as a set of 3D points $\mathbf{x}=(x,y,z)$ with optical attributes $(\sigma,\mathbf{c})$, where $\sigma$ is the volume density of the point and $\mathbf{c}$ is the color of the point.
@@
 2.2 Representation of short dynamic intervals
-To reconstruct the individual chunks of the dynamic scenes, we propose a representation of short dynamic intervals, which serves as the structures of the different branches of our model and handle the dynamic scnes in one chunk efficiently.
+To reconstruct the individual chunks of the dynamic scenes, we propose a representation of short dynamic intervals, which serve as the structures of the different branches of our model and handle the dynamic scenes in one chunk efficiently.
 As shown in [51], naive 4D hash grids for dynamic scenes are short in reconstruction quality.
-To achieve fast reconstruction and compact model size simultaneously, we represent time and spatial features separately and concatenate them to alleviate feature confusion.
-The spatial hash features can be encoded along 3D voxels [31], 2D planes [17, 21], or both [38].
+To simultaneously achieve fast reconstruction and compact model size, we use hash encodings to represent temporal and spatial features separately, then we concatenate them to alleviate feature interference.
+The spatial hash encodings can be implemented as hash tables of 3D voxels [34], orthogonal 2D planes [17, 21], or both [41].
 The time feature is encoded by short hash tables in a per-frame manner.
@@
 As defined in Section 2.1, we split the model into $N_{chunk}$ individual branches to represent the $N_{chunk}$ chunks of the dynamic scene.
-Consequently, the spatial features encoded by different model branches could be fused to reduce the number of the parameters and improve scalability.
+Consequently, the spatial features encoded by different model branches could be fused to reduce the number of parameters and improve scalability.
 This approach, defined as continual dynamic neural graphics primitives (CD-NGP), is designed to handle the dynamic nature of scenes in the online setting.
 Equipped with the hash encoders and MLP regressors defined in Section 2.2, the $k$-th branch of the model contains spatial hash encoder $\Psi_{k}(\mathbf{x})$, temporal hash encoder $\Gamma_{k}(t)$, and MLPs $f_{\theta_{1k}},f_{\theta_{2k}}$.
-Please note the dimensions of the features is independent of the sizes of the hash tables: a larger hash table alleviates hash collision at the cost of more storage consumption, but the features encoded by different sizes of hash tables share the same size, i.e.
+Please note the dimensions of the features are independent of the sizes of the hash tables: a larger hash table alleviates hash collision at the cost of more storage consumption, but the features encoded by different sizes of hash tables share the same size, i.e.
 $\Psi_{k}(x)\in\mathbf{R}^{LF}$ holds for all $k$.
@@
 4 with the color MLP $f_{\theta_{2k}}$.
-To cope with large motions and complex transient effects in the long time intervals, we assign $T_{episode}$ chunks as an episode.
-The MLP regressors are re-trained from initialization in the first branch of each episode to adapt to recent changes.
-We further illustrate the whole process of our method in Algorithm 1, where $\eta_{init}$ and $\eta_{aux}$ denote the numbers of iterations of the initial branch and the auxiliary branches, respectively.
+We further illustrate the entire process of our method in Algorithm 1, where $\eta_{init}$ and $\eta_{aux}$ denote the numbers of iterations of the first branch and the auxiliary branches, respectively.
 $\Psi_{i},\Gamma_{i},\theta_{1}^{(i)},\theta_{2}^{(i)}$ denote the model parameters in the branches.
+The model is trained in a continual learning manner: the first branch is trained from scratch to reconstruct the initial chunk of the dynamic scene, while the MLP decoders of the auxiliary branches are initialized from the previous branch to reconstruct subsequent chunks, facilitating faster convergence.
+To cope with large motions and complex transient effects in long time intervals, we assign $T_{episode}$ chunks as an episode.
+The MLP regressors are re-trained from initialization in the starting branch of each episode to adapt to recent changes.
 Algorithm 1 Continual learning pipeline
@@
 The scalability of our method derives from the redundancy of dynamic scenes and the flexibility of hash table sizes.
-As pointed out by [45], most points ($>90\%$ in DyNeRF dataset by their observation) in the dynamic scenes are static.
+As pointed out by [50], most points ($>90\%$ in the DyNeRF dataset by their observation) in the dynamic scenes are static.
 We thus leverage the shared static points across time and represent the first chunk with a large hash table and the subsequent chunks with small hash tables to ensure scalability.
-Since the spatial hash tables occupy the most parameters in the model, and most chunks are encoded by the subsequent branches, our asymetric hash table design enables continual learning of dynamic scenes and high scalability in streaming scenarios simultaneously.
+Since the spatial hash tables occupy the most parameters in the model, and most chunks are encoded by the subsequent branches, our asymmetric hash table design enables continual learning of dynamic scenes and high scalability in streaming scenarios simultaneously.
 The total bandwidth consumption of our method is $\mathcal{O}(N_{chunk}\cdot 2^{P_{2}}LF+2^{P_{1}}LF)$, while the previous method [48] requires offline training and consumes $\mathcal{O}(\mathbf{K}\cdot 2^{P_{1}}LF)$ bandwidth, where $\mathbf{K}$ ranging from $0.5$ to $16$ is their bandwidth budget controlling the bitrates in streaming scenarios.
@@
 Following the approach of HexPlane [7], we excluded the unsynchronized scene coffee martini and conducted experiments on the remaining scenes, using the first 300 frames of the 1200-frame video flame salmon for evaluation.
-Meetroom dataset [27] is another dataset for dynamic view synthesis that contains 3 scenes captured by 13 synchronized cameras at $1280\times 720$ resolution.
-The scenes are 300 frames long and captured at 30 FPS.
+The Meetroom dataset [27] is another dataset for dynamic NVS that recorded by 13 synchronized cameras with $1280\times 720$ resolution at 30 FPS. Each of the 3 scenes contains 300 frames.
+Although existing datasets have been widely used for dynamic NVS, their limited lengths (300 frames) prevent the evaluation of the scalability of the methods on long video sequences.
 To further validate the effectiveness of our method on long video sequences, we introduce a new dataset comprising 5 dynamic scenes at a resolution of $3840\times 2160$ as shown in Fig.
@@
 Our dataset features more challenging dynamics, including significant non-rigid human motions and transient effects like screen slides, compared to the DyNeRF dataset.
-To comprehensively assess the scalability of our method, we provide 5 scenes with a duration exceeding 200 seconds (6000 frames), which is over 10 times longer than the previous DyNeRF dataset.
-Our dataset is collected to evaluate the scalability of our method and to support further research in the community.
+To comprehensively assess the scalability of our method, we provide 5 scenes with a duration of 6000 frames.
+We hope this dataset will serve as a valuable benchmark for future research in dynamic novel view synthesis, especially in long video sequences.
 We have carefully reviewed the videos to ensure they do not contain any controversial or offensive content before using them for research purposes.
@@
 3.2 Implementation details
-Our implementation is based on HexPlane [7], Instant-NGP [31], and NGP-pl [25].
+We adopt the hash-encoding and tiny MLP implementation of tiny-cuda-nn [33] by Instant-NGP [34], and customize the CUDA-based training and rendering pipelines for the dynamic scenes based on [25].
+The data preparation process is based on HexPlane [7] and LLFF [31].
+Please see the supplementary material for more detailed explanations and analysis.
 All the metrics are measured on a single RTX 3090 GPU if not otherwise specified.
-Similar to the previous work [7, 28], we select the view in the center with index 00 for evaluation on both datasets.
+Similar to the previous work [7, 28], we select the view in the center with index 00 for evaluation on all datasets.
 Following most of the state-of-the-art methods [28, 27, 17, 50, 48], we train and evaluate the model at resolution $1352\times 1014$ on the DyNeRF dataset.
@@
 The model parameters are optimized by Adam [23], with hyper parameters $\epsilon=10^{-15},\beta=(0.9,0.96)$ for stability.
-The learning rate is set as 0.001 and scheduled by cosine function.
+The learning rate is set as 0.001 and scheduled by a cosine function.
 We set the hash table sizes as $2^{P_{1}}=2^{19}$ in the first branch and $2^{P_{2}}=2^{14}$ in the subsequent branches.
@@
 We report the model sizes with the occupancy grid under bfloat16 precision and parameters in float32 precision.
-The episode length $T_{episode}$ is set as $30$ only for videos longer than 300 frames in our own dataset.
+The episode length $T_{episode}$ is set as $30$ only for videos longer than 300 frames in our dataset.
 We set the dimensions of the latent feature $\mathbf{H}$ in Eq.
@@
 We compare our method with both online and offline baselines.
-As demonstrated in Table 1, our CD-NGP is 32×32$\times$32 × faster than the MLP-based continual representation INV [48], $\times4$ smaller than the voxel-based continual representation StreamRF [27] and significantly outperforms the two additional baselines in the continual learning setting.
+Our method shows significantly better efficiency and quality than the other online methods.
+As demonstrated in Table 1, our CD-NGP is 32×32$\times$32 × faster than the MLP-based continual representation INV [53], $\times4$ smaller than the voxel-based continual representation StreamRF [27] and significantly outperforms these baselines in quality metrics.
 The undesirable results from HexPlane-PR and HexPlane-GR show the limitations of the regularization-based and replay-based methods in dynamic scenes: unlike the static scenes, the objects are not stationary across time and the models with limited complexity are not capable of reconstructing unlimited transient effects.
-Benefiting from the parameter isolation like StreamRF [27] and INV [48], our method achieves fast convergence and compact model size simultaneously.
+Benefiting from parameter isolation like StreamRF [27] and INV [53], our method achieves fast convergence and compact model size simultaneously.
 It leverages the hash tables to ensure fast convergence and reuses the features from differently configured spatial hash tables to ensure high scalability.
-Compared with prevailing offline methods [28, 34, 7, 17, 45, 55], our method only requires caching $T_{chunk}=10$ frames into the buffer by continual learning and thus reduces memory occupation by $85\%$.
-Our method also reduces $40\%$ less model size (to $0.4$MB/frame) by leveraging the feature fusion of the base and the auxiliary branches and delivers comparable quality to the offline methods.
+Compared with most of the prevailing offline methods [28, 37, 7, 17, 50], our method only requires caching $T_{chunk}=10$ frames into the buffer by continual learning and thus reduces memory occupation significantly.
+Although the offline baseline 4DGS [61] consumes lower memory by using a lazy load strategy with a trade-off on training speed, their implementation consumes another 27 GB hard disk space on 300-frame DyNeRF dataset, and the space requirement also scales at a complexity of O($N_{frames}$) as the number of frames increases.
+On the contrary, our method leverages the continual learning strategy and enables directly decoding the frames from the videos without any additional storage consumption for training (scales at $O(1)$ complexity).
+It also balances training speed and model size by leveraging the feature fusion of the base and the auxiliary branches and delivers high quality.
 (a) HexPlane-offline
@@
 Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 4: The results of the proposed CD-NGP on DyNeRF dataset.
+Figure 4: The results of the proposed CD-NGP on the DyNeRF dataset.
 Figure 4 provides qualitative comparisons.
@@
 6(a) is the PSNR curve of the compared methods on the trimming scene.
-We use the released configuration of StreamRF [27] for the meetroom dataset to obtain the results.
+We use the default setup of StreamRF [27] for the Meetroom dataset to obtain the results.
 As shown in Table 2, HexPlane-GR and HexPlane-PR produce significantly lower quality than the other methods on the Meetroom dataset, we thus exclude them from Fig.
@@
 6(a) show the advantage of our method over both online and offline baselines.
+We note that the Meetroom dataset provides only 13 views (much fewer than 20+ views in the DyNeRF dataset), which leads to limited coverage for multi-view stereo.
+As a result, COLMAP [44, 43] fails to generate high-quality dense point clouds, making it difficult for 4DGS [61] to obtain reliable initialization.
+Consequently, 4DGS does not produce valid results on this dataset and is excluded from the comparison.
+On the contrary, our method does not require initial point clouds and is more robust.
 It outperforms the implicit MLP-based online method [53] regarding reconstruction quality, speed, and model size.
-Moreover, our method consumes only a fraction of storage than the explicit voxel-based online baseline [27] while maintaining comparable reconstruction quality as compared in Fig.
-5.
+It obtains slightly lower metrics than the explicit voxel-based online method [27], which is likely due to the robustness advantage of explicit voxel representations under sparse-view conditions.
 When compared with offline baselines, our method produces comparable quality to MixVoxels [50] and outperforms HexPlane [7] in terms of quality and speed by much less memory cost.
@@
 (d) the sear steak scene in the DyNeRF dataset.
-Solid and dotted lines denote symmetric and asymmetric setting, respectively.
+Please note solid and dotted lines denote symmetric and asymmetric settings, respectively.
+The CD-NGP/MERF-$P_{1}$-$P_{2}$ annotation denotes different hash table sizes in the base and auxiliary branches defined in Section 2.3.
 Table 2: Comparisons of the metrics on the Meetroom dataset.
@@
 The alternatives of spatial representations include CD-MERF and CD-Plane.
-The element-wise product for plane representation in [17] produces unsatisfactory results in the experiments, we thus remove the $P_{m}(x,y,z)$ in Eq.
+The element-wise product for plane representation in [17] produces unsatisfactory results in the experiments.
+We thus remove the $P_{m}(x,y,z)$ in Eq.
 12 and fuse the features $Q_{mij}(i,j)$ by concatenation in the implementation.
@@
 The metrics show the advantage of our separated encoding of spatial-temporal $(x,y,z,t)$ features explained in Eqs.
-3 and 6: representing the features of spatial coordinates $(x,y,z)$ and temporal coordinate $(t)$ separately avoids feature collisions and lead to better results.
+3 and 6: representing the features of spatial coordinates $(x,y,z)$ and temporal coordinate $(t)$ separately avoids feature collisions and leads to better results.
 The results in Table 3 also show the flexibility in the feature fusion methods.
@@
 3.3.4 Long sequence dataset and scalability analysis.
-As introduced in Section 2.3, our method shows superior scalability: the total size of the model is $\mathcal{O}(N_{chunk}\cdot 2^{P_{2}}LF+2^{P_{1}}LF)$, where $N_{chunk}$ is the number of chunks in the sequence.
-To prove the scalability of our method, we provide quantitative comparisons on our proposed long multi-view video dataset in Table 4 and Fig.
-6(b) to further show the superior scalability of our method.
-We split the sequence into 300-frame chunks for the compared methods since the offline baseline [7] is designed for 300-frame sequences in the DyNeRF dataset.
-We set the hash table size as $2^{12}$ in the auxiliary branches and use $T_{chunk}=5$ in our model.
-As shown in Table 4, our method reduces the memory and storage consumption by at least 4×4$\times$4 × with comparable quality to HexPlane [7].
+As detailed in Section 2.3, our method exhibits superior scalability.
+The overall model size is $\mathcal{O}(N_{chunk}\cdot 2^{P_{2}}LF+2^{P_{1}}LF)$, where $N_{chunk}$ denotes the number of chunks in the sequence.
+We provide quantitative comparisons on our proposed long multi-view video dataset in Table 4 and Fig.
+6(b).
+Following the default configurations on the DyNeRF dataset, the offline baselines HexPlane [7] and MixVoxels [50] are trained with a chunk length of $T_{chunk}=300$.
+Consistent with its DyNeRF configuration, 4DGS [61] adopts a lazy loading strategy, caching frames to disk in order to mitigate memory usage at the expense of training speed.
+In CD-NGP, we set the hash table size in the auxiliary branches to $2^{12}$ and adopt a compact chunk length of $T_{chunk}=5$.
+As shown in Table 4, our method reduces storage consumption by at least 4×4$\times$4 ×, delivers rendering quality comparable to HexPlane [7], and maintains low memory usage and fast training speed.
+Although 4DGS produces better results than our method, it consumes approximately 100 times more storage space.
 Compared with the results in Table 1, the performance gap in storage consumption is even larger in long video sequences.
 This is because the majority of the parameters in our method are used for the auxiliary branches, and the base branch only takes up a small fraction of the parameters.
-Consequently, applying smaller hash tables in the auxiliary branches and leverage the re-use of the features in the base branch leads to a significant reduction in the model size.
+Consequently, applying smaller hash tables in the auxiliary branches and leveraging the reuse of the features in the base branch lead to a significant reduction in the model size.
 The PSNR curves in Fig.
@@
 $(P_{1},P_{2})$ are the log2 values of the hash table lengths of the base branch and the auxiliary branches.
-The whole model contains hash encoders, MLP regressors, and the occupancy gridd, the sizes of the models are thus larger than the value deduced from the combinations of the hash table and MLP regressors.
-Although larger hash tables lead to less hash collisions, the default configuration $(19,14)$ in CD-NGP produces better results than the naive $(16,16)$ configuration even if the storage consumption is $40\%$ lower.
+The whole model contains hash encoders, MLP regressors, and the occupancy grid, the sizes of the models are thus larger than the value deduced from the combinations of the hash table and MLP regressors.
+Although larger hash tables lead to fewer hash collisions, the default configuration $(19,14)$ in CD-NGP produces better results than the naive $(16,16)$ configuration even if the storage consumption is $40.8\%$ lower.
 Moreover, the $(19,12)$ configuration of CD-NGP significantly outperforms the naive $(14,14)$ configuration at a similar storage cost.
@@
 The advantage comes from our re-use of the spatial features encoded by different branches: the base branch with a large hash table recovers more environment details, and the auxiliary branches could thus use fewer parameters to cope with recent changes only.
-The combination of the base encode and the auxiliary encoders lead to a more efficient and scalable model.
+The combination of the base encoder and the auxiliary encoders leads to a more efficient and scalable model.
 Table 4: Quantitative comparisons between our CD-NGP and baselines on the long video sequences in our dataset.
@@
 HexPlane $\checkmark$ 27.74 0.022 0.117 800MB 8 hours 74GB
+4DGS $\checkmark$ 28.85 0.019 0.078 19GB 24 hours 3GB
 HexPlane-PR $\checkmark$ 20.90 0.095 0.286 1272MB 8 hours 12GB
@@
 3.4.2 Temporal representations.
-Our continual learning pipeline is also compatible with a variety of representations for the time axis.
+Our continual learning pipeline is also compatible with a variety of representations for time.
 Currently, there are 2 other prevailing representations for the time axis alone: pure frequency encoding [39] and MLP with frequency-encoding [16].
-We replace the temporal hash encoding in our method with the ones mentioned above on DyNeRF dataset and show the results in Table 6.
+We replace the temporal hash encoding in our method with them on the DyNeRF dataset and show the results in Table 6.
 Our time latent code slightly outperforms others on DyNeRF dataset.
@@
 $(\sigma,\mathbf{c})=(\sigma_{i}+\sigma_{a},\mathbf{c}_{i}+\mathbf{c}_{a}).$ (14)
-Full composition in Table 7 denotes the naive composition of 2 fields from the base branch and the current branch respectively, i.e.
+Full composition in Table 7 denotes the naive composition of two fields from the base branch and the current branch respectively, i.e.
 $\sigma_{i}$ and $\mathbf{c}_{i}$ are the field from the base branch and $\sigma_{a}$ and $\mathbf{c}_{a}$ in Eq.
@@
 3.
-Static-dynamic in Table 7 denotes a naive composition of the field from a static branch by the base encoder and a dynamic branch by the auxiliary encoders similar to [45], i.e.
+Static-dynamic in Table 7 denotes a naive composition of fields from a static branch by the base encoder and a dynamic branch by the auxiliary encoders similar to [50], i.e.
 $\sigma_{i}$ and $\mathbf{c}_{i}$ in Eq.
@@
 9(c) and 9(d) suffer from salt and pepper noise because of the explicit composition, while our CD-NGP circumvents the problem by fusing features in the latent space and produces better results.
+Moreover, the feature combination in the latent space speeds up the inference by reducing the number of MLP queries.
+Please refer to the supplementary materials for more details.
 (a) Ground truth
@@
 5 30.39 0.02602 0.2075 162MB
+Table 10: Comparison between CD-NGP and MVSGS-FT on DyNeRF dataset.
+[…]
+CD-NGP 30.23 0.027 0.198
+MVSGS-FT 26.05 0.086 0.258
 3.4.4 Chunk size.
@@
 We select $T_{chunk}=10$ as the default configuration.
+3.5 Further Discussion
+CD-NGP leverages a highly scalable continual learning framework to tackle dynamic-scene NVS, with a principal benefit being a substantially reduced memory footprint (the feasible 14GB usage on the DyNeRF dataset).
+The multi-branch feature reusing design leads to fast convergence and high fidelity, and the encoder-decoder structure inherited from neural radiance fields (NeRF) [32] enables the model to represent the scenes compactly.
+CD-NGP supports varied kinds of encoders and field composition methods and exhibits high scalability in long video sequences.
+However, the volume rendering process is still required, which limits the real-time rendering capability of our method.
+Please see the supplementary materials for detailed latency analysis.
+Moreover, the current parameter isolation-based design does not support fusing features from different model branches, and thus the reconstruction quality of our method is lower than the state-of-the-art offline baselines.
+In addition, due to the hybrid hash encoder representations, our method suffers from more quality downgrades than the explicit baseline [27] in sparse-view scenarios.
+We acknowledge that several recent cost volume-based multi-view stereo (MVS) NVS techniques also support real-time processing with low memory requirements and thus hold promise for dynamic-scene applications.
+To further evaluate the unique value of continual learning in this context, we include the state-of-the-art method MVSGS [30] as a representative baseline.
+MVSGS [30] leverages a pre-trained model to directly render novel views, as well as generating Gaussian point clouds to support per-scene fine-tuning for best quality.
+We refer to the direct-inference results as MVSGS-D and the fine-tuning results as MVSGS-FT.
+We note that the cost volume module of MVSGS does not support the standard $1352\times 1014$ resolution in the DyNeRF dataset, we thus use the default setup (including resolution) for MVSGS-D.
+Although processing the frames at $\sim 1$ FPS, MVSGS-D produces limited reconstruction quality on the DyNeRF dataset.
+We further compare MVSGS-FT and our CD-NGP in Table 10.
+We train the Gaussians for 5k iterations for each frame in the DyNeRF dataset.
+The training takes around 1.5 minutes for each frame, i.e.
+$\sim 8$ hours for the whole 300-frame sequence.
+The checkpoints consume around 60MB per frame (18 GB in total), which is significantly larger than the most recent explicit offline baselines for dynamic scenes.
+By comparing the quality metrics in Table 10 and the resource consumption, it is clear that our CD-NGP is more efficient: ours provides high-quality results, but consumes much less model size (113MB in total) and training time (75 minutes) by leveraging the re-use of the spatial features in the first branch.
+Hence, the significance of continual learning in dynamic-scene NVS is further validated.
+Please see the supplementary materials for details.
 4 Conclusion and Limitation
 We propose CD-NGP, a continual learning framework for multi-view dynamic scenes based on parameter isolation, along with a self-collected dataset for novel view synthesis under multi-view long-view scenarios.
-Our proposed method splits the input multi-view videos into different chunks, trains different model branches, and performs reconstruction accordingly.
+Our approach partitions the input multi-view videos into different chunks and uses individual model branches to reconstruct the scene accordingly.
 Each branch encodes spatial and temporal features separately using the multi-resolution hash encoding structure.
-It leverages the redundancy of the dynamic scenes and reuses the spatial features in the first branch to achieve fast convergence, high fidelity, and superior scalability.
-On the prevailing DyNeRF dataset, it consumes $40\%$ less bandwidth and $85\%$ less memory than recent explicit offline baselines, achieves comparable quality, and maintains the speed advantage over pure-MLP-based offline methods.
+CD-NGP leverages the redundancy of the dynamic scenes and reuses the spatial features in the first branch to achieve fast convergence, high fidelity, and superior scalability.
+On the prevailing DyNeRF dataset, it consumes much less memory, achieves comparable quality with a smaller model size than most recent explicit offline baselines, and maintains the speed advantage over traditional pure MLP-based offline methods.
 Compared with online baselines, it delivers state-of-the-art quality and achieves a new balance among reconstruction quality, training speed, bandwidth occupation, and memory cost.
@@
 CD-NGP 24.43 0.037 0.193 26.01 0.039 0.191
-Table 12: Detailed results of the proposed CD-NGP and other hash-based methods on our long video dataset.
+Table 13: Detailed results of the proposed CD-NGP and other methods on our long video dataset.
 Model Demo Discussion Kungfu
@@
 HexPlane 28.92 0.016 0.095 27.17 0.024 0.136 25.14 0.032 0.136
+4DGS 28.38 0.022 0.092 29.92 0.016 0.064 25.53 0.029 0.099
 HexPlane-PR 21.10 0.090 0.268 20.98 0.094 0.315 18.85 0.112 0.309
@@
 HexPlane 27.88 0.018 0.099 29.59 0.018 0.120 27.74 0.022 0.117
+4DGS 29.38 0.014 0.067 31.02 0.015 0.067 28.85 0.019 0.078
 HexPlane-PR 21.85 0.087 0.287 21.78 0.092 0.280 20.91 0.095 0.292
@@
 CD-NGP 26.70 0.031 0.212 28.97 0.024 0.169 27.13 0.039 0.201
+Appendix B Access to the long video dataset.
+We provide a public link to access our self-collected long video dataset. It can be accessed via the following BaiduNetdisk link:
+Appendix C Discussion on rendering latency.
+The proposed method CD-NGP fuses spatial and temporal features in a multi-branch NeRF model, which is efficient in terms of training and memory consumption.
+Moreover, the feature combination in the latent space avoids querying multiple fields for rendering, and thus achieves better rendering latency.
+However, the volume rendering process adopted from NeRF-variants is inherently computationally intensive, which limits the real-time rendering capability.
+In this section, we provide a detailed analysis of the rendering latency of our method on the DyNeRF dataset.
+Taking the cut roasted beef scene in the DyNeRF dataset (at $1352\times 1014$ resolution) as an example, we provide a detailed analysis of the rendering latency in Table 14.
+The rendering process of CD-NGP is divided into three parts: ray marching, NeRF model, and pixel-level volume rendering as follows:
+  • Ray marching: The ray marching process is the same as that in the implementation [25] of Instant-NGP [34], which is a standard process in hash-based NeRF.
+This process generates a split (batch) of sampling coordinates from the camera to the scene along each ray, and each image requires multiple passes to render.
+Typically, each pass generates a split of up to 9M samples on the 1M rays (1M pixels), and rendering the entire image requires around 200M samples.
+  • NeRF model: As elaborated in the manuscript, our CD-NGP is a multi-branch NeRF model that incorporates hash-based encoders and MLP-based decoders.
+As shown in Table 14, both the spatial and temporal hash encoders are efficient and lightweight, requiring only 7 ms and 5 ms, respectively.
+The primary computational cost arises from the two MLP decoders (55 ms in total), which is common and inevitable in NeRF-based methods.
+Although CD-NGP avoids multiple passes of MLP decoders by fusing features in the latent space, one fused decoding pass of MLP decoders is still required for each sampled point.
+This fused pass takes approximately 33 ms for the XYZT-fusion MLP and 22 ms for the RGB MLP, resulting in a total NeRF model latency of around 76 ms per pass.
+  • Volume rendering: This process loads the optical attributes $\sigma,\textbf{c}$ of the sampled points and computes the pixel color C by numerical integral (accumulating the optical attributes along the ray).
+The rendering process is similar to that of Instant-NGP [34] and other NeRF-based methods [25].
+If the accumulated opacity of a ray (pixel) reaches $10^{-4}$, the sampling and rendering are terminated.
+Thus, the whole rendering time is smaller than the product of the number of passes and the time of each pass.
+Table 14: Rendering time breakdown and point scale of each component.
+Module Time Scale
+Ray marching 3ms $\sim$1M pixels
+NeRF model
+    Spatial representation 7ms $\sim$9M samples
+    Time representation 5ms $\sim$9M samples
+    XYZT-fusion MLP 33ms $\sim$9M samples
+    RGB MLP 22ms $\sim$9M samples
+    Total (NeRF model) 76ms $\sim$9M samples
+Rendering a split 2ms $\sim$1M pixels
+Rendering the whole image 500ms $\sim$200M samples
+As illustrated, each query of the hash table-based NeRF model takes less than 100ms, but the pixel-based volume rendering process requires multiple passes to render the whole image.
+This limitation is a common issue in all NeRF-based methods and is orthogonal to the contribution of the proposed feature fusion method and continual learning pipeline.
+Appendix D Comparison with real-time processing method.
+We compare the per-scene results of our method and MVSGS-FT on the DyNeRF dataset in Table 15.
+Our method produces comparable quality with MVSGS-FT in the first 4 scenes and is more robust in the last sear steak scene.
+The method MVSGS-D does not support the standard $1352\times 1014$ resolution on the DyNeRF dataset and produces limited quality as shown in Fig.
+11, we thus exclude it from Table 15.
+Refer to caption Refer to caption
+(a) (b)
+Figure 10: Output images of MVSGS-FT on the DyNeRF dataset. (a) iteration 3,000, (b) iteration 5,000.
+Figure 11: Output images of MVSGS-D on the DyNeRF dataset. The left side is the ground truth, and the right side is the rendering output.
+Table 15: Detailed results of the proposed CD-NGP and real-time processing method MVSGS.
+Model Flame Salmon Cook Spinach Cut Roasted Beef
+[…]
+CD-NGP 26.37 0.046 0.289 30.76 0.024 0.179 31.03 0.024 0.180
+MVSGS-FT 24.75 0.054 0.190 30.57 0.023 0.123 29.68 0.025 0.131
+Model Flame Steak Sear Steak Average
+[…]
+CD-NGP 31.27 0.021 0.174 31.71 0.019 0.170 30.23 0.027 0.198
+MVSGS-FT 30.02 0.027 0.126 15.26 0.301 0.722 26.05 0.086 0.258
+We provide the rendered images of MVSGS-FT for better visualization in Fig.
+10.
+We note that there are two approaches for fine-tuning the MVSGS: the first is to use the output of the MVS network to re-initialize Gaussians for each frame, and the second is to use the Gaussians of the previous frame for initialization.
+However, we find that the second approach leads to unstable training and produces unsatisfactory results (much lower PSNRs).
+Therefore, we only report the metrics of the first approach and refer to it as MVSGS-FT.
+Appendix E Implementation Details about the spatial and temporal hash encoding
+E.1 Implementation
+In the implementation of tiny-cuda-nn, the authors have provided interfaces for hash-encoders with different input dimensions, such as 1D, 2D, and 3D.
+However, the 1D hash encoding is not enabled by default.
+In our implementation, we enable the interfaces for one-dimension hash encodings of tiny-cuda-nn [33] to support the temporal encoding in our continual learning framework.
+Specifically, we use the 1D hash encoding for the time dimension and the 3D hash encoding for the spatial dimensions.
+E.2 Scalability on size.
+We measure the hash encoder size and the inference latency of different configurations of the encodings in Table 16.
+The latency is measured on a single NVIDIA RTX 3090 GPU.
+Since all hash tables are stored in float32 precision, each element of the table consumes $F$ parameters, and the hash encoder contains $L$ hash tables for different encoding resolutions.
+The upper bound of the total hash encoder size is $4\times L\times F\times 2^{P}$ bytes.
+Since some spatial hash table resolutions are coarse (e.g., $16\times 16\times 16=2^{12}<2^{14}<2^{P_{1}=19}$), and the tables are not fully filled, the actual hash encoder size is smaller than the upper bound.
+Clearly, the hash encoder size is primarily influenced by the length of the hash tables (i.e., $2^{P}$).
+By controlling the length of the hash tables and reusing features from the base spatial hash table, we can achieve a good trade-off between hash encoder size and rendering quality.
+E.3 Scalability on inference speed.
+Table 16 additionally presents the inference speed across different configurations.
+We render the first frame of the cut roasted beef scene from the DyNeRF dataset, which has a resolution of $1352\times 1014$.
+The latency increases approximately linearly with both the number of layers $L$ and the feature vector length $F$, with $L$ exhibiting a greater impact.
+Increasing the hash table length $P$ also affects inference speed, primarily due to reduced memory locality for larger table sizes.
+Table 16: Inference speed and model size under different configurations.
+L: levels, F: feature dim, P: log2 values of the hash table length.
+Base and auxiliary denote the hash encoders of the base spatial representation and auxiliary spatial representation, respectively.
+Temporal denotes the hash encoder of the time dimension.
+Type L F P Size (MB) Samples Latency(ms)
+base 12 2 19 37.83 5483712 3.33
+auxiliary 12 2 14 1.41 5483712 3.16
+temporal 2 20 7 0.006 5483712 2.89
+base 16 2 19 51.03 5483712 4.48
+auxiliary 16 2 14 1.91 5483712 4.27
+temporal 2 20 7 0.006 5483712 2.96
+base 24 2 19 77.53 5483712 6.82
+auxiliary 24 2 14 2.86 5483712 6.51
+temporal 2 20 7 0.006 5483712 2.89
+base 12 4 19 75.65 5483712 5.33
+auxiliary 12 4 14 2.82 5483712 5.08
+temporal 2 20 7 0.006 5483712 3.01
+base 12 2 21 140.83 5483712 3.64
+auxiliary 12 2 14 1.41 5483712 3.16
+temporal 2 20 7 0.006 5483712 3.19
+base 12 2 20 73.82 5483712 3.45
+auxiliary 12 2 14 1.41 5483712 3.16
+temporal 2 20 7 0.006 5483712 3.23
+base 12 2 19 37.83 5483712 3.31
+auxiliary 12 2 16 5.26 5483712 3.24
+temporal 2 20 7 0.006 5483712 2.87
+base 12 2 19 37.83 5483712 3.30
+auxiliary 12 2 14 1.41 5483712 3.14
+temporal 1 20 7 0.003 5483712 1.39
+base 12 2 19 37.83 5483712 3.29
+auxiliary 12 2 14 1.41 5483712 3.15
+temporal 1 40 7 0.006 5483712 2.87